### v1

# Pesin theory for transcendental maps and applications

Abstract
In this paper, we develop Pesin theory for the boundary map of some Fatou components of transcendental functions, under certain hyptothesis on the singular values and the Lyapunov exponent. That is, we prove that generic inverse branches for such maps are well-defined and conformal. In particular, we study in depth the Lyapunov exponents with respect to harmonic measure, providing results which are of independent interest.
As an application of our results, we describe in detail generic inverse branches for centered inner functions, and we prove density of periodic boundary points for a large class of Fatou components. Our proofs use techniques from measure theory, ergodic theory, conformal analysis, and inner functions, as well as estimates on harmonic measure.
1 Introduction
In the setting of smooth dynamical systems, hyperbolic dynamical systems play a distinguished role, since they are the easiest to study and exhibit the simplest possible behaviour. Indeed, hyperbolic dynamics are characterized by the presence of expanding and contracting directions for the derivative at every point, which provides strong local, semilocal or even global information about the dynamics. However, the assumption of hyperbolicity is quite restrictive.
A weaker (and hence, more general) form of hyperbolicity, known as non-uniform hyperbolicity, was initially developed by Yasha Pesin in his seminal work [Pes76, Pes77]. Since then, Pesin’s approach to hyperbolicity, also known as Pesin theory, has been extended, generalized and refined in numerous articles and research books (see e.g. [Pol93], [KH95, Supplement], [BP23]). Although results apply to both discrete and continuous dynamical systems, in this paper we focus on the discrete ones.
Roughly speaking, Pesin studied originally $\mathcal{C}^{1}$-diffeomorphisms on compact smooth Riemannian manifolds. Under the assumption that such a map is measure-preserving and ergodic, and no Lyapunov exponent vanishes except on a set of zero measure, the forward and backwards contraction or expansion around almost every point is controlled asymptotically by the Lyapunov exponents.
One of the natural generalizations of Pesin theory is to the setting of iteration of rational maps in the Riemann sphere. That is, let $f\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be holomorphic, and consider the discrete dynamical system generated by $f$. The phase space $\widehat{\mathbb{C}}$ is commonly split into two totally invariant sets: the Fatou set $\mathcal{F}(f)$, where the family of iterates is normal, and hence the dynamics are in some sense stable; and its complement, the Julia set $\mathcal{J}(f)$. Although the Fatou set is well-understood, the dynamics in the Julia set are more intricate and worthy of study. For general background in rational iteration we refer to [CG93, Mil06].
In contrast with the setting of $\mathcal{C}^{1}$-diffeomorphisms considered by Pesin, now the iterated function is no longer bijective, which is overcome by assuming a higher degree of regularity on the function. Indeed, a rational map fails to be bijective only at the finitely many points where $f^{\prime}(z)=0$, known as critical points, and hence inverse branches fail to be defined in their images, known as critical values.
In this setting, a rational map is said to be hyperbolic if all orbits of critical values are compactly contained in the Fatou set, which already implies that all inverse branches around points in $\mathcal{J}(f)$ are well-defined and uniformly contracting (see e.g. [CG93, Sect. V.2], [Mil06, Sect. 19]). Hence, following Pesin’s approach for diffeomorphisms, it is natural to ask whether, for a general map (not necessarily hyperbolic), generic inverse branches are well-defined and contracting. Note that one should make precise the notion of generic inverse branches, by defining the abstract space of backward orbits for points in $\mathcal{J}(f)$ and endow it with a measure (using Rokhlin’s natural extension, see Sect. 2.5).
One can prove that, under the assumption of existence of an ergodic invariant probability with positive Lyapunov exponent, for almost every backbard orbit $\left\{x_{n}\right\}_{n}$ there exists a disk around the initial point $x_{0}$, such that the corresponding inverse branches of $f^{n}$ are well-defined and contracting in this disk (see [PU10, Sect. 11.2], [KU23, Chap. 9.3], [URM23, Sect. 28.3]). The proof relies strongly on the fact that $\widehat{\mathbb{C}}$ is compact (and hence, $\mathcal{J}(f)$ is also compact), the finiteness of critical values, and the existence of an ergodic invariant probability, with positive Lyapunov exponent.
We note that the existence of ergodic invariant probabilities supported on the Julia set of rational maps has been historically a topic of wide interest, in connection with the measure of maximal entropy. For polynomials, the existence of such a measure was already proved by Brolin [Bro65], whereas for rational maps it was done by Freire, Lopes and Mañé [FLM83], and Lyubich [Lyu83], independently. Such a measure of maximal entropy is known to be an ergodic invariant probability, and hence it can be used as an initial cornerstone to develop Pesin theory. Moreover, Lyapunov exponents with respect to any ergodic invariant probability supported on $\mathcal{J}(f)$ have been studied in depth [Prz85, Prz93].
The goal of this paper is to extend these well-known results for rational maps to the transcendental setting, that is, for maps $f\colon\mathbb{C}\to\widehat{\mathbb{C}}$ (transcendental) meromorphic. Although it is important to note that, under the presence of poles, some orbits get truncated, one can define the Fatou and Julia set for $f$ in a similar way than for rational maps, so that the Fatou set $\mathcal{F}(f)$ is the set where the dynamics are stable, and the Julia set $\mathcal{J}(f)$ is the locus of chaos (for precise definitions and properties, see Sect. 3.1 and references therein). The question we want to address is the same: whether generic inverse branches are well-defined and contracting around points in $\mathcal{J}(f)$.
Note that the cornerstones from which the rational Pesin theory is built (namely, compact phase space, finitely many critical values, and existence of ergodic invariant probabilities) no longer hold in general. Indeed, first, the phase space is now $\mathbb{C}$, which is no longer compact, and nor is the Julia set. In fact, this lack of compactness causes difficulties even for the extension of the notion of hyperbolicity from the rational setting [RGS17].
Additionally, critical values are not the only values where inverse branches fail to be defined. Indeed, one shall consider the set of singular values (i.e. critical and asymptotic values, and accumulation thereof), and it may be uncountable. Note also that the functions we iterate have always infinite degree.
Finally, the existence of invariant measures on the Julia set is much more delicate and remains somewhat unexplored, as well as Lyapunov exponents (which depend on the existence of the previous measures). Indeed, although the existence of invariant ergodic probabilities supported on the Julia set has been proved for certain families (such as the hyperbolic exponential family [UZ03], and also in [IR22], for a wider family of hyperbolic maps), in other cases it is known that they do not exist [DS08]. Hence, in contrast with rational maps, the existence of an ergodic invariant measure supported in the Julia set is unknown in the general setting.
In order to overcome some of these difficulties, we restrict ourselves to some forward invariant subsets of the Julia set which are of special interest: the boundaries of invariant (or periodic) connected components of the Fatou set (known as Fatou components). If we let $U$ be an invariant Fatou component for $f$, then its boundary $\partial U$ is forward invariant under $f$, and in the seminal work of Doering and Mañé [DM91], invariant ergodic measures for $f\colon\partial U\to\partial U$ supported on $\partial U$ are given, following the approach initiated by Przytycki to study the boundary map of attracting basins of rational maps [Prz85].
Taking advantage of these invariant measures, under some mild assumptions on the geometric distribution of singular values and the order of growth of the function, we are able to overcome the difficulties arising from the lack of compactness, the infinite degree and the presence of infinitely many singular values. Our techniques include refined estimates on harmonic measure and the construction of an appropriate conformal metric. In this manner, we can develop Pesin theory in the boundary of some transcendental Fatou components in a quite successful way, which is presented next.
Statement of results
Let $f\colon\mathbb{C}\to\widehat{\mathbb{C}}$ be a transcendental meromorphic function, i.e. so that $\infty$ is an essential singularity for $f$, and let $U$ be an invariant Fatou component for $f$. It is well-known that such an invariant Fatou component is either an attracting basin, a parabolic basin, a rotation domain or a Baker domain (see Sect. 3.1). In the sequel, we denote by $\partial U$ the boundary of $U$ in $\mathbb{C}$. All the derivatives are understood to be with respect to the spherical metric on $\widehat{\mathbb{C}}$, and hence $\left|f^{\prime}\right|$ is bounded on compact subsets of the plane.
The transversal assumption throughout the paper is that singular values in ${\partial}U$ satisfy some geometric condition, to which we refer as singular values being thin in ${\partial}U$. Since the definition for general maps is rather technical, we restrict ourselves to meromorphic maps and inner functions in the introduction, and refer to Definition 3.8 for details.
Roughly speaking, singular values being thin in ${\partial}U$ means that we do allow singular values to accumulate on ${\partial}U$ as long as they do it fast enough, or if they accumulate at $\infty$, if they stay at a positive distance from $\partial U$. In particular, we do allow infinitely many singular values in ${\partial}U$. Next we state the precise definition, taking into account that a subset $K\subset X$ is $\delta$-separated if for all $x,y\in K$, $x\neq y$, we have $\textrm{dist}(x,y)\geq\delta$.
Definition.
(Thin singular values for meromorphic functions) Let $f\colon\mathbb{C}\to\widehat{\mathbb{C}}$ be a transcendental meromorphic function, and let $U$ be a Fatou component for $f$. We say that singular values are thin in ${\partial}U$ if there exist $R,\delta>0$, $\mu\in(0,1)$ and $d\in\mathbb{N}$ such that
  1. (a)
    for every $n\in\mathbb{N}$, there are at most $n^{d}$ singular values in
    $\left\{z\in{\mathbb{C}}\colon\left|z\right|<R,\ \textrm{dist}(z,{\partial}U)<\mu^{n}\right\}$
    which are $\mu^{n}$-separated;
  2. (b)
    if $v\in\mathbb{C}$ is a singular value with $\left|z\right|>R$, then
    $\textrm{dist}(z,\partial U)>\delta.$
Intuitively, the first condition says that, if there are accumulation points of singular values (as a set) in $\partial U$, then these points are accumulated exponentially fast. The second condition states that singular values of large modulus lie at a positive (Euclidean) distance of $\partial U$. Moreover, if singular values are thin in $\partial U$, they have zero harmonic measure (Rmk. 3.11). Note also that, if there are only finitely many singular values in $\partial U$, then singular values are trivially thin in $\partial U$, but our definition allow singular values in $\partial U$ to be infinite. For more details, see Section 3.2.
Attracting basins are the natural candidates to perform Pesin theory on its boundary, since the harmonic measure $\omega_{U}$ (with basepoint the fixed point $p\in U$) is invariant under $f$ and ergodic (Thm. 3.3). Hence, our main result is the following, which claims that generic inverse branches are well-defined in the boundary of an attracting basin, under the following hypotheses.
Theorem A.
(Pesin theory for attracting basins of transcendental maps) Let $f\colon\mathbb{C}\to\widehat{\mathbb{C}}$ be a meromorphic function, and let $U$ be a simply connected attracting basin for $f$, with attracting fixed point $p\in U$. Let $\omega_{U}$ be the harmonic measure in $\partial U$ with base point $p$. Assume $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$ with $\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}>0$, and singular values are thin in ${\partial}U$.
Then, for every countable collection of measurable sets $\left\{A_{k}\right\}_{k}\subset\partial U$ with $\omega_{U}(A_{k})>0$, and for $\omega_{U}$-almost every $x_{0}\in\partial U$, there exists a backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ and $r>0$ such that
  1. (a)
    there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A_{k}$;
  2. (b)
    the inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},r)$;
  3. (c)
    diam $\textrm{\em diam }F_{n}(D(x_{0},r))\to 0$, as $n\to\infty$.
Note that, in particular, for $\omega_{U}$-almost every $x_{0}\in\partial U$ there exists a backward orbit $\left\{x_{n}\right\}_{n}$, with well-defined inverse branches $\left\{F_{n}\right\}_{n}$, which is dense in $\partial U$.
If we consider parabolic basins or Baker domains, the situation is even more unfavorable, since no harmonic measure on $\partial U$ is $f$-invariant. However, restricting ourselves to entire functions and using the first return map, we develope a similar result for parabolic basins and recurrent Baker domains (i.e. those such that $f|_{\partial U}$ is recurrent, see Thm. 8.3). In these cases no harmonic measure is $f$-invariant, but there exists a $\sigma$-finite measure which is absolutely continuous with respect to it, invariant under $f$, recurrent and ergodic (Thm. 3.3), and hence exploitable by means of the first return map. As far as we are aware of, this result is new even for polynomials.
Theorem B.
(Pesin theory for entire functions) Let $f\colon\mathbb{C}\to\mathbb{C}$ be an entire function, and let $U$ be an attracting or parabolic basin, or recurrent Baker domain. Let $\omega_{U}$ be a harmonic measure on $\partial U$, such that $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$ with $\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}>0$. Assume critical values of $f$ are finite.
Then, for every countable collection of measurable sets $\left\{A_{k}\right\}_{k}\subset\partial U$ with $\omega_{U}(A_{k})>0$, and for $\omega_{U}$-almost every $x_{0}\in\partial U$ there exists a backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ and $r>0$ such that
  1. (a)
    there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A_{k}$;
  2. (b)
    the inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},r_{0})$;
  3. (c)
    for every subsequence $\left\{x_{n_{j}}\right\}_{j}$ with $x_{n_{j}}\in D(x_{0},r)$, diam $\textrm{\em diam }F_{n_{j}}(D(x_{0},r))\to 0$, as $j\to\infty$.
Applications
Next we present two applications of the theorems above: developing Pesin theory for centered inner functions, and finding periodic points in the boundary of Fatou components of transcendental maps.
Let $\mathbb{D}$ denote the unit disk, and $\partial\mathbb{D}$ the unit circle, and let $\lambda$ be the normalized Lebesgue measure in $\partial\mathbb{D}$. An inner function is, by definition, a holomorphic self-map of the unit disk, $g\colon\mathbb{D}\to\mathbb{D}$, which preserves the unit circle $\lambda$-almost everywhere in the sense of radial limits. If, in addition, we have that $g(0)=0$, we say that the inner function is centered.
In general, inner functions may present a highly discontinuous behaviour in $\partial\mathbb{D}$. Indeed, a point $\xi\in\partial\mathbb{D}$ is called a singularity of $g$ if $g$ cannot be continued analitically to any neighbourhood of $\xi$. Denote the set of singularities of $g$ by $E(g)$.
It is well-known that radial extension of centered inner functions preserve the Lebesgue measure and are ergodic (see e.g. [DM91, Thm. A, B]). For these reasons, centered inner functions have been widely studied as measure-theoretical dynamical systems [Aar78, Pom81, DM91, Cra91, Cra92, Aar97, IU23, IU24].
An important subset of centered inner functions are the ones with finite entropy, or equivalently, when $\log\left|g^{\prime}\right|\in L^{1}(\partial\mathbb{D})$ [Cra91]. Such a property translates to a greater control on the dynamics, from different points of view (see e.g [Cra91, Cra92, IU23, IU24]). In particular, centered inner functions with finite entropy are natural candidates to apply the theory developed above. Moreover, due to its rigidity and symmetries, we will deduce some additional properties. Since, in general, inner functions present a highly discontinuous behaviour in $\partial\mathbb{D}$, it is noteworthy the great control we achieve, only by assuming that singular values are thin, which for inner functions is defined as follows.
Definition.
(Thin singular values for inner functions) Let $f\colon\mathbb{D}\to{\mathbb{D}}$ be an inner function. We say that singular values are thin in ${\partial}\mathbb{D}$ if there exists $\varepsilon>0$, $\mu\in(0,1)$, $d\in\mathbb{N}$ and $v_{1},\dots,v_{k}\in{\partial}\mathbb{D}$ such that
  1. (a)
    for every $n\in\mathbb{N}$, there are at most $n^{d}$ singular values in
    $\left\{z\in\mathbb{D}\colon\left|z\right|>1-\varepsilon\right\}\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$
    which are $\mu^{n}$-separated;
  2. (b)
    for $i=1,\dots,k$, there exists a horodisk $H_{i}$ of center $v_{i}\in\partial\mathbb{D}$ and radius $\delta_{i}>0$ such that all singular values in $D(v_{i},\varepsilon)\cap\mathbb{D}$ lie in the horodisk $H_{i}$.
Let us denote the radial segment at $\xi$ of length $\rho>0$ by
$R_{\rho}(\xi)\coloneqq\left\{r\xi\colon r\in(1-\rho,1)\right\}.$
and the Stolz angle at $\xi$ of length $\rho>0$ and opening $\alpha\in(0,\frac{\pi}{2})$ by
$\Delta_{\alpha,\rho}(\xi)=\left\{z\in\mathbb{D}\colon\left|\textrm{Arg }\xi-\textrm{Arg }(\xi-z)\right|<\alpha,\left|z\right|>1-\rho\right\}.$
Then, even though inner functions may have more than one essential singularity, we are able to do the same construction as in Theorem A (which only depend on singular values being thin, not in the number of singularities). We deduce the following.
Corollary C.
(Pesin theory for centered inner functions) Let $g\colon\mathbb{D}\to\mathbb{D}$ be a centered inner function, such that $\log\left|g^{\prime}\right|\in L^{1}(\partial\mathbb{D})$. Fix $\alpha\in(0,\pi/2)$. Assume singular values are thin in $\partial\mathbb{D}$. Then, for every countable collection of measurable sets $\left\{A_{k}\right\}_{k}\subset\partial\mathbb{D}$ with $\lambda(A_{k})>0$, and for $\lambda$-almost every $\xi_{0}\in\partial\mathbb{D}$ there exists a backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ and $\rho_{0}>0$ such that
  1. (a)
    there exists a sequence $n_{k}\to\infty$ such that $\xi_{n_{k}}\in A_{k}$;
  2. (b)
    the inverse branch $G_{n}$ of $g^{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho_{0})$;
  3. (c)
    for all $\rho\in(0,\rho_{0})$,
    $G_{n}(R_{\rho}(\xi_{0}))\subset\Delta_{\alpha,\rho}(\xi_{n}).$
In particular, the set of singularities $E(g)$ has zero $\lambda$-measure.
Another application of the previous construction of inverse branches in the boundary of a Fatou component is to find density of periodic boundary points, as in the seminal paper of F. Przytycki and A. Zdunik for rational maps [PZ94]. To do so, we need a stronger assumption on the orbits of singular values in $U$. Recall that, given a simply connected domain $U$, we say that $C\subset U$ is a crosscut if $C$ is a Jordan arc such that $\overline{C}=C\cup\left\{a,b\right\}$, with $a,b\in\partial U$, $a\neq b$. Any of the two connected components of $U\smallsetminus C$ is a crosscut neighbourhood. We define the postsingular set of $f$ in $U$ as
$P_{U}(f)\coloneqq\overline{\bigcup\limits_{s\in SV(f)\cap U}\bigcup\limits_{n\geq 0}f^{n}(s)}.$
Corollary D.
(Periodic boundary points are dense) Under the hypotheses of Theorem A or Theorem B, assume, in addition, there exists a crosscut neighbourhood $N_{C}$ with $N_{C}\cap P_{U}(f)=\emptyset$. Then, accessible periodic points are dense in $\partial U$.
Note that, although we need some control on the postsingular set in $U$, we do not put any restriction on the postsingular set outside $U$. Observe that the assumptions of Corollary D are always satisfied if singular values in $U$ are compactly contained in $U$, even if $f|_{U}$ has infinite degree.
Lyapunov exponents of transcendental maps
Finally, we note that one essential hypothesis in our results is that $\log\left|f^{\prime}\right|$ is integrable with respect to the harmonic measure $\omega_{U}$, and hence the Lyapunov exponent
$\chi_{\omega_{U}}(f)=\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}$
is well-defined. We also require that $\chi_{\omega_{U}}$ is positive. These facts are well-known for simply connected basins of attraction of rational maps [Prz85, Prz93], but unexplored for transcendental maps. In this paper we give several conditions, concerning the order of growth of the function and the shape of the Fatou component, which implies that the Lyapunov exponent is well-defined and non-negative.
One of the main challenges that appear when considering transcendental maps is that $\left|f^{\prime}\right|$ may not be bounded in $\partial U$, even when taking the derivative with respect to the spherical metric. Indeed, $\left|f^{\prime}\right|$ is not bounded around the essential singularity, and the growth can be arbitrarily fast. Thus, we introduce the following concept, which relates the growth of the function with the shape of the Fatou component.
Definition.
(Order of growth in a sector for meromorphic functions) Let $f\colon\mathbb{C}\to\widehat{\mathbb{C}}$ be a transcendental meromorphic function, and let $U\subset\mathbb{C}$ be an invariant Fatou component for $f$. We say that $U$ is asymptotically contained in a sector of angle $\alpha\in\left(0,1\right)$ with order of growth $\beta>0$ if there exists $R_{0}>0$, $\xi\in\partial\mathbb{D}$ and $\alpha\in(0,1)$, such that, if
$S_{R}=S_{R,\alpha}\coloneqq\left\{z\in\mathbb{C}\colon\left|z\right|>R,\ \left|\textrm{Arg }\xi-\textrm{Arg }(1/\overline{z})\right|<\pi\alpha\right\}$
then,
  1. (a)
    $U\cap\left\{z\in\mathbb{C}\colon\left|z\right|>R_{0}\right\}\subset S_{R_{0}}$;
  2. (b)
    $f$ has order of growth $\beta>0$ in $S_{R_{0}}$, i.e. there exists $A,B>0$ such that, for all $R>R_{0}$ and $z\in S_{R_{0}}\smallsetminus S_{R}$,
    $A\cdot e^{B\cdot\left|z\right|^{-\beta}}\leq\left|f^{\prime}(z)\right|\leq A\cdot e^{B\cdot\left|z\right|^{\beta}}.$
Under this asusmption on the growth, we are able to prove the following.
Proposition E.
($\log\left|f^{\prime}\right|$ is $\omega_{U}$-integrable) Let $f\colon\mathbb{C}\to\widehat{\mathbb{C}}$ be a meromorphic function, and let $U$ be an invariant Fatou component for $f$. Let $\omega_{U}$ be a harmonic measure on $\partial U$. Assume $U$ is asymptotically contained in a sector of angle $\alpha\in\left(0,1\right)$, with order of growth $\beta\in(0,\frac{1}{2\alpha})$. Then, $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$.
Proposition F.
(Non-negative Lyapunov exponents) Let $f\colon\mathbb{C}\to\widehat{\mathbb{C}}$ be a meromorphic function, and let $U$ be a simply connected attracting basin. Let $\omega_{U}$ be the harmonic measure in $\partial U$ with base point $p$. Assume
  1. (a)
    $U$ is asymptotically contained in a sector of angle $\alpha\in\left(0,1\right)$, with order of growth $\beta\in(0,\frac{1}{2\alpha})$;
  2. (b)
    $SV(f)$ are thin in $\partial U$.
Then,
$\chi_{\omega_{U}}=\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}\geq 0.$
Remark.
The statements of Theorems A, B and C are a simplified version of the ones we prove inside the paper (respectively, Thms. 5.1, 6.4 and 7.14). These stronger statements are formulated in terms of the Rohklin’s natural extension of the corresponding dynamical systems. Since this construction is not common in transcendental dynamics (although it is standard in ergodic theory), we chose to present our results in this simplified (and weaker) form. For convenience of the reader, all the needed results about Rohklin’s natural extension can be found in Section 2.5. Nevertheless, for simplicity, the proof of Corollary D is done without using Rohklin’s natural extension.
Finally, although the results are stated here for meromorphic functions, we shall work in the more general class $\mathbb{K}$ of functions with countably many singularities; in particular, this allows us to consider periodic attracting basins of meromorphic maps, not only invariant ones. The technicalities that arise when working in class $\mathbb{K}$ are explained in Section 3.1. Note also that the definition of thin singular values has to be adapted to deal with several singularities (see Def. 3.8), as well as the definiton of order of growth in a sector (see Def. 3.13).
Notation. We denote by $\widehat{\mathbb{C}}$, $\mathbb{C}$ and $\mathbb{D}$, the Riemann sphere, complex plane, and unit disk, respectively. The normalized Lebesgue measure in the unit circle $\partial\mathbb{D}$ is denoted by $\lambda$. We designate by $D(x,r)$ the Euclidean disk of center $x$ and radius $r>0$. We denote by $\textrm{dist}(z,w)$ the Euclidean distance between two points in the complex plane.
The boundary (in $\widehat{\mathbb{C}}$) of a domain $U\subset\widehat{\mathbb{C}}$ is denoted by ^$\widehat{\partial}U$. The notation ${\partial}U$ is reserved to denote the boundary in the domain of definition of the function we iterate, that is, if we fix $f\in\mathbb{K}$, we denote by $\Omega(f)$ its domain of definition, and by ${\partial}U$ the boundary of $U$ in $\Omega(f)$ (more details are given in Sect. 3.1). The class of harmonic measures in $U$ is denoted by $\omega_{U}$, and by $\omega_{U}(p,\cdot)$ when we need to specify the basepoint (see Sect. 2.6).
Finally, $f^{-1}(z)$ denotes all the preimages under $f$ of the point $z$ (setwise). When we refer to the inverse branch, we write $F_{n,z,w}$ meaning that $F_{n,z,w}$ is an inverse branch of $f^{n}$ sending $z$ to $w$. When it is clear from the context, we just write $F_{n}$ to lighten the notation.
Structure of the paper. Section 2 is devoted to gather all the preliminary results used throughout the paper, including distortion estimates on conformal maps, measure theory and abstract ergodic theory. Due to its relevance in the paper, we describe in detail the construction of Rokhlin’s natural extension for $\sigma$-finite measures (Sect. 2.5) and estimates on harmonic measure (Sect. 2.6). Section 3 is devoted to comment on the setting and the hypotheses we work with, namely, functions of class $\mathbb{K}$, thin singular values and growth in sectors. Section 4 deals with Lyapunov exponents for boundaries of Fatou components of transcendental maps, proving Propositions E and F. Sections 5 and 6 are devoted to the development of Pesin theory, proving Theorems A and B, respectively. Sections 7, 8 and 9 are dedicated to the applications, proving Corollaries C and D. Further questions are discussed in Section 10.
We note that, although it is customary to introduce the inner function associated to a simply connected Fatou component as the main tool for the study of its boundary dynamics, we postpone this until Section 8, when dealing with the associated inner function to a Fatou component. This way, the reader can go through the exposition on Pesin theory (Thms. A and B), and on Lyapunov exponents (Prop. E and F), with no previous knowledge on inner functions, only assuming the ergodic properties of the map $f|_{\partial U}$ (which of course come from the ergodic properties of the associated inner function in $\partial\mathbb{D}$). This simplifies considerably the exposition. To prove that periodic boundary points are dense (Corol. D), we do need the associated inner function, and all the machinery developed before. Therefore this proof is adjourned until the end of the paper (Sect. 9).
Acknowledgments. First of all, I am indebted to my supervisor, Núria Fagella. I also want to thank Lasse Rempe, for asking me the question which motivates this work. Besides, I am indebted with Anna Zdunik, for all her explanations and her encouragement for starting this project, as well as with Oleg Ivrii, for interesting discussions and his valuable insights. I also want to thank Phil Rippon for his help and his kindness for sharing with me some of his knowledge in harmonic measure. I am also thankful to Jana Rodríguez-Hertz and the KTH in Stockholm, for a master class in Pesin theory, and Nikolai Prochorov and the Séminaire Rauzy in Marseille, for interesting discussions on the topic.
2 Preliminaries
In this section we gather the tools we use throughout the article, including distortion estimates for univalent maps, measure theory, and abstract ergodic theory. We put special emphasis on Rokhlin’s natural extension, which will be an essential tool throughout the paper. Although all the results in this section seem to be well-known, we include the proof of those for which we could not find a written reference.
We start with the following easy remark on the composition of inverse branches, which will use recurrently throughout the paper, without noticing explicitly.
Remark.
(Composition of inverse branches) Let $\left\{x_{n}\right\}_{n}$ be an infinite backward orbit, i.e. $f(x_{n+1})=x_{n}$ for $n\geq 0$. Assume there exists $r>0$ such that all inverse branches $F_{n,x_{0},x_{n}}$ are well-defined and conformal in $D(x_{0},r)$, for $n\geq 0$. We claim that this is equivalent to say that the inverse branches $F_{1,x_{n},x_{n+1}}$ are well-defined and conformal in $F_{n,x_{0},x_{n}}(D(x_{0},r))$, for $n\geq 0$.
It is clear that if the inverse branches $F_{1}$ are well-defined along the backward orbit, then by composing them we get $F_{n}$. The converse is more subtle; assume $F_{n,x_{0},x_{n}}$ and $F_{n,x_{0},x_{n+1}}$ are well-defined in $D(x_{0},r)$, and let us see that $F_{1,x_{n},x_{n+1}}$ is well-defined in $F_{n,x_{0},x_{n}}(D(x_{0},r))$. Indeed, we have the following commutative diagram.
[…]
Then, the map
$\varphi\colon F_{n,x_{0},x_{n}}(D(x_{0},r))\longrightarrow F_{n+1,x_{0},x_{n+1}}(D(x_{0},r)),$
defined as $F_{n+1,x_{0},x_{n+1}}\circ f^{n}$, is well-defined and conformal in $F_{n,x_{0},x_{n}}(D(x_{0},r))$, $\varphi(x_{n})=x_{n+1}$, and it is a local inverse branch of $f$. Hence, $\varphi=F_{1,x_{n},x_{n+1}}$ in $F_{n,x_{0},x_{n}}(D(x_{0},r))$, as desired.
2.1 Distortion estimates for univalent maps
We need the following results concerning the distortion for univalent maps.
Theorem 2.1.
(Koebe’s distortion estimates, [BFJK20, p. 639]) Let $z\in\mathbb{C}$, $r>0$, and let $\varphi\colon D(z,r)\to\mathbb{C}$ be a univalent map. Then,
$D\left(\varphi(z),\frac{1}{4}\cdot\left|\varphi^{\prime}(z)\right|\cdot r\right)\subset\varphi(D(z,r)).$
Moreover, for all $\lambda\in(0,1)$ and $z\in\overline{D(x,\lambda r)}$, it holds
$\left|\varphi^{\prime}(x)\right|\cdot\dfrac{1-\lambda}{(1+\lambda)^{3}}\leq\left|\varphi^{\prime}(z)\right|\leq{\left|\varphi^{\prime}(x)\right|}\cdot\dfrac{1+\lambda}{(1-\lambda)^{3}},$
$\varphi(D(x,\lambda r))\subset D\left(\varphi(x),r\cdot{\left|\varphi^{\prime}(x)\right|}\cdot\dfrac{1+\lambda}{(1-\lambda)^{3}}\right).$
2.2 Conformal metrics
Let $V\subset{\mathbb{C}}$ be an open connected set. A conformal metric on $V$ is a Riemannian metric of the form $d\rho(z)=\rho(z)\left|dz\right|$ for a positive continuous function $\rho$ on $V$, where $\left|dz\right|$ denotes the Euclidean metric in $\mathbb{C}$.
Given a piecewise $\mathcal{C}^{1}$-curve $\gamma\subset V$, we define its length with respect to $\rho$ as
$\textrm{length}_{\rho}(\gamma)\coloneqq\int_{\gamma}\rho(z)\left|dz\right|.$
The distance between two points $z,w\in V$ with respect to the metric $d\rho$, denoted by $\textrm{dist}_{\rho}(z,w)$, is defined as the infimum of lengths (with respect to $\rho$) of piecewise $\mathcal{C}^{1}$-curves $\gamma\subset V$ joining $z$ and $w$.
The diameter of a set $A\subset V$ with respect to $d\rho$ is defined as
$\textrm{diam}_{\rho}\coloneqq\sup\left\{\textrm{dist}_{\rho}(z,w)\colon z,w\in A\right\},$
while $D_{\rho}(x,r)$ denotes the disk of center $x\in V$ and radius $r>0$, with respect to $d\rho$.
2.3 Measure theory
We gather here the concepts of measure theory that are needed throughout the paper, including elementary properties of measurable transformations, the definitions of Lebesgue space and Lebesgue density point, and the first Borel-Cantelli Lemma.
In the sequel, we let $(X,\mathcal{A},\mu)$ be a measure space. If $\mu(X)=1$, we say that $(X,\mathcal{A},\mu)$ is a probability space. Note that any measure space with finite measure (i.e. $\mu(X)<\infty$) can be turned into a probability space by reescaling the measure. We say that the measure $\mu$ is $\sigma$-finite if there exists a countable collection of measurable sets $\left\{A_{n}\right\}_{n}\subset\mathcal{A}$ such that $X=\bigcup_{n}A_{n}$ and $\mu(A_{n})<\infty$.
Recall that, given $(X_{1},\mathcal{A}_{1})$ and $(X_{2},\mathcal{A}_{2})$ measurable spaces, we say that $T\colon X_{1}\to X_{2}$ is measurable if $T^{-1}(A)\in\mathcal{A}_{1}$, for every $A\in\mathcal{A}_{2}$.
Definition 2.2.
(Properties of measurable maps) Let $(X_{1},\mathcal{A}_{1},\mu_{1})$ and $(X_{2},\mathcal{A}_{2},\mu_{2})$ be measure spaces, and let $T\colon X_{1}\to X_{2}$ be measurable. Then, $T$ is:
  • non-singular, if, for every $A\in\mathcal{A}_{2}$, it holds $\mu_{1}(T^{-1}(A))=0$ if and only if $\mu_{2}(A)=0$;
  • measure-preserving, if, for every $A\in\mathcal{A}_{2}$, it holds $\mu_{1}(T^{-1}(A))=\mu_{2}(A)$.
Definition 2.3.
(Lebesgue space) A measure space $(X,\mathcal{A},\mu)$ is called a Lebesgue space if it is isomorphic to the interval $\left[0,\mu(X)\right)$ (equipped with classical Lebesgue measure) together with countably many atoms. If $\mu(X)=1$, we say it is a Lebesgue probability space.
Mostly, we will use the measure space $(\partial\mathbb{D},\mathcal{B}(\partial\mathbb{D}),\lambda)$, where $\mathcal{B}(\partial\mathbb{D})$ denotes the Borel $\sigma$-algebra of $\partial\mathbb{D}$, and $\lambda$, its normalized Lebesgue measure. It is clear that $(\partial\mathbb{D},\mathcal{B}(\partial\mathbb{D}),\lambda)$ is a Lebesgue space. Throughout the article we will omit the dependence in the $\sigma$-algebra whenever it is clear from the context. Note that the next theorems are stated for $(\partial\mathbb{D},\mathcal{B}(\partial\mathbb{D}),\lambda)$ but hold for any compact Euclidean space endowed with the (normalized) Lebesgue measure.
Definition 2.4.
(Lebesgue density) Given a Borel set $A\in\mathcal{B}(\partial\mathbb{D})$, the Lebesgue density of $A$ at $\xi\in\partial\mathbb{D}$ is defined as
$d_{\xi}(A)\coloneqq\lim\limits_{\rho\to 0}\dfrac{\lambda(A\cap D(\xi,\rho))}{\lambda(D(\xi,\rho))}.$
A point $\xi\in\partial\mathbb{D}$ is called a Lebesgue density point for $A$ if $d_{\xi}(A)=1$.
Proposition 2.5.
(Almost every point is a Lebesgue density point, [Rud87, p. 138]) Given a Borel set $A\in\mathcal{B}(\partial\mathbb{D})$, with $\lambda(A)>0$, then $\lambda$-almost every point in $A$ is a Lebesgue density point for $A$.
Lemma 2.6.
(First Borel-Cantelli lemma, [Bog07, 1.12.89]) Let $(X,\mathcal{A},\mu)$ be a probability space, let $\left\{A_{n}\right\}_{n}\subset\mathcal{A}$, and let
$B\coloneqq\left\{x\in X\colon x\in A_{n}\textrm{ for infinitely many }n\textrm{'s}\right\}=\bigcap\limits_{k=1}^{\infty}\bigcup_{n=k}^{\infty}A_{n}.$
Then, if $\sum_{n=1}^{\infty}\mu(A_{n})<\infty$, it holds $\mu(B)=0$.
2.4 Abstract Ergodic Theory
We recall some basic notions used in abstract ergodic theory (for more details, see e.g. [PU10, Haw21, URM22]).
Definition 2.7.
(Ergodic properties of measurable maps) Let $(X,\mathcal{A},\mu)$ be a measure space, and let $T\colon X\to X$ be measurable. Then,
  • $\mu$ is $T$-invariant if $T$ is measure-preserving;
  • $T$ is recurrent with respect to $\mu$, if for every $A\in\mathcal{A}$ and $\mu$-almost every $x\in A$, there exists a sequence $n_{k}\to\infty$ such that $T^{n_{k}}(x)\in A$;
  • $T$ is ergodic with respect to $\mu$, if $T$ is non-singular and for every $A\in\mathcal{A}$ with $T^{-1}(A)=A$, it holds $\mu(A)=0$ or $\mu(X\smallsetminus A)=0$.
In the sequel, if it is clear with which measure are we working with, we shall omit the dependence on the measure. We note that, if $T$ is invertible and ergodic, then $T^{-1}$ is also ergodic.
Theorem 2.8.
(Poincaré Recurrence Theorem, [Haw21, Thm. 2.12]) Let $(X,\mathcal{A},\mu)$ be a probability space, and let $T\colon X\to X$ be a measurable transformation. Assume $T$ is $\mu$-preserving. Then, $T$ is recurrent with respect to $\mu$.
Theorem 2.9.
(Almost every orbit is dense, [Aar97, Prop. 1.2.2]) Let $(X,\mathcal{A},\mu)$ be a measure space, and let $T\colon X\to X$ be non-singular. Then, the following are equivalent.
  1. (a)
    $T$ is ergodic and recurrent with respect to $\mu$.
  2. (b)
    For every $A\in\mathcal{A}$ with $\mu(A)>0$ and $\mu$-almost every $x\in X$, there exists a sequence $n_{k}\to\infty$ such that $T^{n_{k}}(x)\in A$.
Theorem 2.10.
(Birkhoff Ergodic Theorem, [KH95, Sect. 4.1]) Let $(X,\mathcal{A},\mu)$ be a probability space together with a measure-preserving transformation $T\colon X\to X$, and let $\varphi\in L^{1}(\mu)$. Then,
$\lim\limits_{n}\frac{1}{n}\sum_{k=0}^{n-1}\varphi(T^{k}(x))$
exists for $\mu$-almost every $x\in X$. If $T$ is an automorphism, the equality
$\lim\limits_{n}\frac{1}{n}\sum_{k=0}^{n-1}\varphi(T^{k}(x))=\lim\limits_{n}\frac{1}{n}\sum_{k=0}^{n-1}\varphi(T^{-k}(x))$
holds $\mu$-almost everywhere.
Finally, if $T$ is ergodic with respect to $\mu$, then for $\mu$-almost every $x\in X$ it holds
[…]
2.5 Rokhlin’s natural extension
A useful technique in the study of non-invertible measure-preserving tranformations is the so-called Rokhlin’s natural extension, which allows us to construct a measure-preserving automorphism in an abstract measure space, mantaining its ergodic properties. However, this technique is often developed for Lebesgue spaces with invariant probabilities (see e.g [PU10, Sect. 1.7] , [URM22, Sect. 8.5]). Since we will work with both probabilities and $\sigma$-finite measures, we sketch how we can develop the theory in this more general case.
Therefore, let $(X,\mathcal{A},\mu)$ be a measure space, together with a measure-preserving transformation $T\colon X\to X$. We assume that the measure $\mu$ is either finite (and in this case we assume, without loss of generality, that it is a probability), or a $\sigma$-finite measure.
Consider the space of backward orbits for $T$
$\widetilde{X}=\left\{\left\{x_{n}\right\}_{n}\subset X\colon\hskip 4.26773ptx_{0}\in X,\hskip 4.26773ptT(x_{n+1})=x_{n},\hskip 4.26773ptn\geq 0\right\},$
and define, in a natural way, the following maps. On the one hand, for $k\geq 0$, let $\pi_{k}\colon\widetilde{X}\to X$ be the projection on the $k$-th coordinate of $\left\{x_{n}\right\}_{n}$, that is $\pi_{k}(\left\{x_{n}\right\}_{n})=x_{k}$. On the other hand, we define Rokhlin’s natural extension of $T$ as $\widetilde{T}\colon\widetilde{X}\to\widetilde{X}$, with
$\widetilde{T}(\left\{x_{n}\right\}_{n})=\widetilde{T}(x_{0}x_{1}x_{2}\dots)=f(x_{0})x_{0}x_{1}\dots$
It is clear that $\widetilde{T}$ is invertible and $\widetilde{T}^{-1}$ is the shift-map, i.e.
$\widetilde{T}^{-1}(\left\{x_{n}\right\}_{n})=\widetilde{T}^{-1}(x_{0}x_{1}x_{2}\dots)=x_{1}x_{2}x_{3}\dots=\left\{x_{n+1}\right\}_{n}.$
Moreover, for each $k\geq 0$, the following diagram commutes.
[…]
Note that, up to here, the construction is purely symbolic and measures have not come out yet. In fact, the next step in the construction is to endow the space $\widetilde{X}$ with an appropriate $\sigma$-algebra $\widetilde{\mathcal{A}}$ and a measure $\widetilde{\mu}$, which makes the previous projections $\pi_{k}$ and the map $\widetilde{T}$ measure-preserving. To do so, we will need the following more general result.
Theorem 2.11.
(Kolmogorov Consistency Theorem, [Par67, Thm. V.3.2]) Let $(X_{n},\mathcal{A}_{n},\mu_{n})$ be Lebesgue probability spaces, and let $T_{n}\colon X_{n+1}\to X_{n}$ be measure-preserving. Let
$\widetilde{X}=\left\{\left\{x_{n}\right\}_{n}\colon\hskip 4.26773ptx_{n}\in X_{n},\hskip 4.26773ptT_{n}(x_{n+1})=x_{n},\hskip 4.26773ptn\geq 0\right\}.$
and let $\pi_{k}\colon\widetilde{X}\to X_{k}$ be the projection on the $k$-th coordinate. Then, there exists a $\sigma$-algebra $\widetilde{A}$ and a probability $\widetilde{\mu}$ in $\widetilde{X}$ such that $(\widetilde{X},\widetilde{A},\widetilde{\mu})$ is a Lebesgue probability space and, for each $k\geq 0$,
$\widetilde{\mu}(\pi^{-1}_{k}(A))=\mu_{k}(A),\hskip 14.22636ptA\in\mathcal{A}_{k}.$
Notice that the theorem above holds whenever $(X_{n},\mathcal{A}_{n},\mu_{n})$ are Lebesgue measure spaces with finite measure. The $\sigma$-algebra $\widetilde{A}$ can be taken to be the smallest which makes each projection $\pi_{k}\colon\widetilde{X}\to X_{k}$ measurable [Par67, Thm. V.2.5]. Note that $T_{k}\circ\pi_{k+1}=\pi_{k}$. Observe that now $\widetilde{X}$ stands for the space of backward orbits under the sequence of maps $\left\{T_{n}\right\}_{n}$. Hence, one has to think of $\widetilde{X}$ as the infinite product of the spaces $\left\{X_{n}\right\}_{n}$, since the spaces in $\left\{X_{n}\right\}_{n}$ are a priori different, and hence there is no endomorphism $\widetilde{T}\colon\widetilde{X}\to\widetilde{X}$ in general. However, we will use these extensions $(\widetilde{X},\widetilde{A},\widetilde{\mu})$ of some appropriate spaces as building blocks for Rokhlin’s natural extension for transformations with $\sigma$-finite invariant measures.
Theorem 2.12.
(Rokhlin’s natural extension for $\sigma$-finite invariant measures) let $(X,\mathcal{A},\mu)$ be a Lebesgue space, and let $T\colon X\to X$ be a measure-preserving transformation. Assume $\mu$ is a $\sigma$-finite measure, and consider Rokhlin’s natural extension $\widetilde{T}\colon\widetilde{X}\to\widetilde{X}$. Then, there exists a $\sigma$-algebra $\widetilde{\mathcal{A}}$ and a $\sigma$-finite measure $\widetilde{\mu}$ such that the maps $\pi_{k}$ and $\widetilde{T}$ are measure-preserving.
Proof.
In the case of $(X,\mathcal{A},\mu)$ being a Lebesgue probability space, the statement follows from applying Kolmogorov Consistency Theorem 2.11 with $X_{n}=X$, for all $n\geq 0$, as indicated in [URM22, Thm. 8.4.2].
Otherwise, let $\left\{X^{j}_{0}\right\}_{j}$ be a partition of $X$ such that $\mu(X_{0}^{j})$ is finite, for each $j\geq 0$. Without loss of generality, we assume $\mu(X_{0}^{j})=1$, for each $j\geq 0$, to simplify the computations. Then, for all $n\geq 0$, $\left\{X_{n}^{j}\coloneqq T^{-n}(X_{0}^{j})\right\}_{j}$ is also a partition of $X$ such that $\mu(X_{n}^{j})=1$, for each $j\geq 0$, since $T$ is measure-preserving and preimages of disjoint sets are disjoint.
If we write $\mathcal{A}_{n}^{j}$ and $\mu_{n}^{j}$ for the restrictions of $\mathcal{A}$ and $\mu$ to $X_{n}^{j}$, we have that, for each $j\geq 0$, $(X_{n}^{j},\mathcal{A}_{n}^{j},\mu_{n}^{j})$ is a Lebesgue probability space, and $T\colon X^{j}_{n+1}\to X^{j}_{n}$ is measure-preserving. Hence, by Theorem 2.11, there exists a Lebesgue probability space $(\widetilde{X^{j}},\widetilde{\mathcal{A}^{j}},\widetilde{\mu^{j}})$ such that
$\widetilde{X^{j}}=\left\{\left\{x_{n}\right\}_{n}\colon\hskip 4.26773ptx_{n}\in X^{j}_{n},\hskip 4.26773ptT(x_{n+1})=x_{n},\hskip 4.26773ptn\geq 0\right\}.$
and the projections $\pi^{j}_{k}\colon\widetilde{X^{j}}\to X_{k}^{j}$ are measure-preserving. It is clear that the space of backward orbits
$\widetilde{X}=\left\{\left\{x_{n}\right\}_{n}\colon\hskip 4.26773ptx_{n}\in X,\hskip 4.26773ptT_{n}(x_{n+1})=x_{n},\hskip 4.26773ptn\geq 0\right\}$
is the disjoint union of the $\widetilde{X^{j}}$, $j\geq 0$. We consider $\mathcal{A}$ to be the $\sigma$-algebra generated by $\left\{\widetilde{\mathcal{A}^{j}}\right\}_{j}$, and the measure $\widetilde{\mu}$ on $(\widetilde{X},\widetilde{A})$ unambiguously determined by the $\widetilde{\mu^{j}}$’s. It is clear that the maps $\pi_{k}$ are measure-preserving, for all $k\geq 0$.
In other words, once we fix an initial condition $x_{0}\in X^{j}_{0}$, for some $j\geq 0$, then all backward orbits starting at $x_{0}$ can be followed using $(X_{n}^{j},\mathcal{A}_{n}^{j},\mu_{n}^{j})$. Indeed, if we have $\left\{x_{n}\right\}_{n}\in\widetilde{X}$ with $x_{0}\in X^{j}_{0}$, and we label the partition so that $x_{n}\in X^{j_{n}}_{0}$, then for the natural extension $\widetilde{T}\colon\widetilde{X}\to\widetilde{X}$, we have
$\widetilde{T}\colon\widetilde{X^{j_{n+1}}}\longrightarrow\widetilde{X^{j_{n}}}.$
Therefore, we have that the $X_{0}^{j_{n}}=X^{j}_{n}$, and the projection morphisms $\pi^{j_{n}}_{0}\colon\widetilde{X^{j_{n}}}\to X_{0}^{j_{n}}$ and $\pi^{j}_{n}\colon\widetilde{X^{j_{0}}}\to X^{j_{0}}_{n}$ coincide. Thus, we have the following measure-preserving commutative diagram.
[…]
It is left to see that $\widetilde{T}$ is measure-preserving. To do so, note that
$\left\{(\pi_{n}^{j})^{-1}(A\cap X^{j}_{n})\colon A\in\mathcal{A}\right\}_{n,j}$
generates the $\sigma$-algebra $\widetilde{A}$, hence it is enough to prove invariance for such sets. Thus, without loss of generality, let $A\subset X^{j}_{n}$, and then
$\widetilde{\mu}\circ\widetilde{T}^{-1}((\pi_{n}^{j})^{-1}(A))$ $=\widetilde{\mu}\circ(\pi_{n}^{j}\circ\widetilde{T})^{-1}(A)=\widetilde{\mu}\circ(T\circ\pi_{0}^{j_{n}})^{-1}(A)=\widetilde{\mu}\circ(\pi_{0}^{j_{n}})^{-1}\circ T^{-1}(A)$
$=\mu(T^{-1}(A))=\mu(A)=\widetilde{\mu}((\pi_{n}^{j})^{-1}(A)),$
as desired. $\square$
It follows from the previous theorem that $\widetilde{\mu}$ is a probability if and only if so is $\mu$. Natural extensions share many ergodic properties with the original map, as shown in the following proposition for probability spaces.
Proposition 2.13.
(Ergodic properties of Rokhlin’s natural extension) Let $(X,\mathcal{A},\mu)$ be a Lebesgue probability space, endowed with a measure-preserving transformation $T\colon X\to X$, and consider its Rokhlin’s natural extension $\widetilde{T}$ acting in $(\widetilde{X},\widetilde{\mathcal{A}},\widetilde{\mu})$, given by Theorem 2.12. Then, the following holds.
  1. (a)
    $\widetilde{T}$ is recurrent with respect to $\widetilde{\mu}$.
  2. (b)
    $\widetilde{T}$ is ergodic with respect to $\widetilde{\mu}$ if and only if ${T}$ is ergodic with respect to ${\mu}$.
Proof.
  1. (a)
    Since $\mu$ is assumed to be a probability, $\widetilde{\mu}$ is also a probability, and the recurrence of $\widetilde{T}$ follows from Poincaré Recurrence Theorem 2.8.
  2. (b)
    It is proven in [URM22, Thm. 8.4.3].
$\square$
Under the assumption of ergodicity and recurrence, we can prove that every subset of positive measure in the phase space is visited by almost every backward orbit.
Corollary 2.14.
(Almost every backward orbit is dense) Let $(X,\mathcal{A},\mu)$ be a Lebesgue space, endowed with a measure-preserving transformation $T\colon X\to X$, and consider its Rokhlin’s natural extension $\widetilde{T}$ acting in $(\widetilde{X},\widetilde{\mathcal{A}},\widetilde{\mu})$, given by Theorem 2.12. Assume $\widetilde{T}$ is ergodic and recurrent with respect to $\widetilde{\mu}$, and $A\subset X$ is a measurable set with $\mu(A)>0$. Then, for $\widetilde{\mu}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{X}$, there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A$.
Proof.
Since $\widetilde{T}$ is ergodic and recurrent with respect to $\widetilde{\mu}$, by Theorem 2.9, for every $\widetilde{A}\in\widetilde{\mathcal{A}}$ with $\widetilde{\mu}(\widetilde{A})>0$ and $\widetilde{\mu}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{X}$, there exists a sequence $n_{k}\to\infty$ such that $\widetilde{T}^{-n_{k}}(\left\{x_{n}\right\}_{n})\in\widetilde{A}$. Taking $\widetilde{A}$ to be $\pi_{0}^{-1}(A)$, we have that $\widetilde{\mu}(\widetilde{A})>0$, so for $\widetilde{\mu}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{X}$, there exists a sequence $n_{k}\to\infty$ with
$\widetilde{T}^{-n_{k}}(\left\{x_{n}\right\}_{n\geq 0})=\left\{x_{n}\right\}_{n\geq n_{k}}\in\pi_{0}^{-1}(A).$
Hence, $x_{n_{k}}\in A$, as desired. $\square$
2.6 Radial extensions and harmonic measure
Let $U\subset\widehat{\mathbb{C}}$ be a hyperbolic simply connected domain (i.e. $U$ omits at least three points), and let $\varphi\colon\mathbb{D}\to U$ be a Riemann map. We are concerned with the extension of $\varphi$ to the unit circle $\partial\mathbb{D}$ given in terms of radial limits. More in general, we say that a function $h\colon\mathbb{D}\to\widehat{\mathbb{C}}$ has radial limit at $\xi$ if the limit
$h^{*}(\xi)\coloneqq\lim\limits_{t\to 1^{-}}h(t\xi)$
exists, and we say that the map $h^{*}\colon\partial\mathbb{D}\to\widehat{\mathbb{C}}$ is the radial extension of $h$. The following results imply that radial extensions of continuous maps are measurable, and, for Riemann maps, radial extensions are well-defined $\lambda$-almost everywhere.
Theorem 2.15.
(Radial extensions are measurable, [Pom92, Prop. 6.5]) Let $h\colon\mathbb{D}\to\widehat{\mathbb{C}}$ be continuous. Then, the points $\xi\in\partial\mathbb{D}$ where the radial limit $h^{*}$ exists form a Borel set, and if $B\subset\widehat{\mathbb{C}}$ is a Borel set, then
$(h^{*})^{-1}(B)\coloneqq\left\{\xi\in\partial\mathbb{D}\colon h^{*}(\xi)\in B\right\}\subset\partial\mathbb{D}$
is also a Borel set.
Theorem 2.16.
(Existence of radial limits; Fatou, Riesz and Riesz, [Mil06, Thm. 17.4]) Let $\varphi\colon\mathbb{D}\to U$ be a Riemann map. Then, for $\lambda$-almost every point $\xi\in\partial\mathbb{D}$, the radial limit $\varphi^{*}(\xi)$ exists. Moreover, if we fix $\xi\in\partial\mathbb{D}$ for which $\varphi^{*}(\xi)$ exists, then $\varphi^{*}(\xi)\neq\varphi^{*}(\zeta)$, for $\lambda$-almost every point $\zeta\in\partial\mathbb{D}$.
Definition of harmonic measure
Given a simply connected domain $U$, we use the radial extension of its Riemann map
$\varphi^{*}\colon\partial\mathbb{D}\to\widehat{\partial}U$
to define a measure in ^$\widehat{\partial}U$, the harmonic measure, in terms of the push-forward of the normalized Lebesgue measure on the unit circle $\partial\mathbb{D}$.
Definition 2.17.
(Harmonic measure) Let $U\subset\widehat{\mathbb{C}}$ be a hyperbolic simply connected domain, $z\in U$, and let $\varphi\colon\mathbb{D}\to U$ be a Riemann map, such that $\varphi(0)=z\in U$. Let $(\partial\mathbb{D},\mathcal{B},\lambda)$ be the measure space on $\partial\mathbb{D}$ defined by $\mathcal{B}$, the Borel $\sigma$-algebra of $\partial\mathbb{D}$, and $\lambda$, its normalized Lebesgue measure. Consider the measurable space $(\widehat{\mathbb{C}},\mathcal{B}(\widehat{\mathbb{C}}))$, where $\mathcal{B}(\widehat{\mathbb{C}})$ is the Borel $\sigma$-algebra of $\widehat{\mathbb{C}}$. Then, given $B\in\mathcal{B}(\widehat{\mathbb{C}})$, the harmonic measure at $z$ relative to $U$ of the set $B$ is defined as
$\omega_{U}(z,B)\coloneqq\lambda((\varphi^{*})^{-1}(B)).$
Note that the harmonic measure $\omega_{U}(z,\cdot)$ is well-defined. Indeed, by Theorem 2.15, the set
$(\varphi^{*})^{-1}(B)=\left\{\xi\in\partial\mathbb{D}\colon\varphi^{*}(\xi)\in B\right\}$
is a Borel set of $\partial\mathbb{D}$, and hence measurable. We also note that the definition of $\omega_{U}(z,\cdot)$ is independent of the choice of $\varphi$, provided it satisfies $\varphi(0)=z$, since $\lambda$ is invariant under rotations.
We refer to [GM05, Pom92] for equivalent definitions and further properties of the harmonic measure. The rest of the section is devoted to collect the properties that we will use throughout the paper.
Properties of harmonic measure
We start with the following simple facts.
Lemma 2.18.
(Sets of zero and full harmonic measure) Let $U\subset\widehat{\mathbb{C}}$ be a hyperbolic simply connected domain, and $B\in\mathcal{B}(\widehat{\mathbb{C}})$. If there exists $z_{0}\in U$ such that $\omega_{U}(z_{0},B)=0$ (resp. $\omega_{U}(z_{0},B)=1$), then $\omega_{U}(z,B)=0$ (resp. $\omega_{U}(z,B)=1$) for all $z\in U$. In this case, we say that the set $B$ has zero (resp. full) harmonic measure relative to $U$, and we write $\omega_{U}(B)=0$ (resp. $\omega_{U}(B)=1$).
Note that, in particular, properties such as ergodicity and recurrence, which only depend on zero and full measure sets, are well-defined for harmonic measure without specifying the basepoint.
Lemma 2.19.
(Comparison of harmonic measures, [Con95, Prop. 21.1.13]) Let $V,\ U\subset\widehat{\mathbb{C}}$ be hyperbolic simply connected domains, with $V\subset U$. Then, for all $z\in V$ and any measurable set $B\subset\widehat{\partial}V\cap\widehat{\partial}U$,
$\omega_{V}(z,B)\leq\omega_{U}(z,B).$
We are interested in the support of $\omega_{U}$. Recall that
supp $\textrm{supp }\omega_{U}\coloneqq\left\{x\in\widehat{\mathbb{C}}\colon\textrm{for all }r>0,\textrm{ }\omega_{U}(D(x,r))>0\right\}.$
Note that it only depends on the sets of zero measure, hence it is well-defined without specifying the basepoint of the harmonic measure.
Lemma 2.20.
(Support of harmonic measure) Let $U\subset\widehat{\mathbb{C}}$ be a hyperbolic simply connected domain. Then,
supp $\textrm{\em supp }\omega_{U}=\widehat{\partial}U.$
Therefore, for all $x\in\widehat{\partial}U$ and $r>0$, $\omega_{U}(D(x,r))>0$.
Proof.
It follows easily from considering an equivalent definition of harmonic measure in terms of solutions to the Dirichlet problem, see e.g. [Con95, Chap. 21]. $\square$
We will use several times the fact that harmonic measure is invariant under Möbius transformations, which follows from the uniqueness of the Riemann map. Although it seems to be a folklore result, we did not find an explicit reference for it, so we include its proof.
Lemma 2.21.
(Harmonic measure and Möbius transformations) Let $U\subset\widehat{\mathbb{C}}$ be a hyperbolic simply connected domain, $z\in U$, and consider the harmonic measure $\omega_{U}(z,\cdot)$. Then, if $M$ is a Möbius transformation, $M(U)$ is a hyperbolic simply connected domain, and, for all $B\in\mathcal{B}(\widehat{\mathbb{C}})$ it holds
$\omega_{U}(z,B)=\omega_{M(U)}(M(z),M(B)).$
Proof.
Since Möbius transformations are conformal maps of the Riemann sphere, it is clear that $M(U)$ is a hyperbolic simply connected domain.
Now, let $\varphi_{U}\colon\mathbb{D}\to U$ be a Riemann map such that $\varphi(0)=z\in U$. Then, $\varphi_{M(U)}\colon\mathbb{D}\to M(U)$ defined as $\varphi_{M(U)}=M\circ\varphi_{U}$ is a Riemann map such that $\varphi(0)=M(z)\in M(U)$. Since harmonic measure does not depend on the Riemann map we use to define it, provided it sends 0 to the basepoint $M(z)$, we have
$\omega_{M(U)}(M(z),M(B))=\lambda((\varphi_{M(U)}^{*})^{-1}(M(B)))=\lambda((\varphi_{U}^{*})^{-1}(B))=\omega_{U}(z,B),$
as desired. $\square$
Finally, we will need the following facts concerning integrability.
Lemma 2.22.
(Logarithm is integrable, [Con95, Prop. 21.1.18]) For every $a\in\mathbb{C}$, $\log\left|z-a\right|\in L^{1}(\omega_{U})$.
Lemma 2.23.
(Shrinking targets, [PU10, Lemma 11.2.1]) Let $\mu$ be a finite Borel measure on $\mathbb{C}$, and let $a\in\mathbb{C}$ be such that $\log\left|z-a\right|\in L^{1}(\mu)$. Then, for every $C>0$ and $10<t<$ 1,
$\sum_{n\geq 0}\mu(D(a,C\cdot t^{n}))<\infty.$
Estimates on the harmonic measure of disks
Throughout the article, we will need different estimates on the harmonic measure of disks centered in a boundary point. For some applications, it is enough to have the summability condition given by Lemma 2.23. However, sometimes it will be necessary a stronger control on the measure in terms of the radius of the disk. More precisely, for $x\in\partial U$, we are interested in quantifying the decay of
$\omega_{U}(z_{0},D(x,r))$
when $r\to 0$, as a function of $r$.
We note that the choice of the basepoint $z_{0}$ is irrelevant. Indeed, it follows from Harnack’s inequality (see e.g. [Con95, Thm. 21.1.7]) that for any $z_{0},z_{1}\in U$ there exists $\rho\coloneqq\rho(z_{0},z_{1})>0$ such that
$\rho^{-1}\cdot\omega_{U}(z_{1},\cdot)\leq\omega_{U}(z_{0},\cdot)\leq\rho\cdot\omega_{U}(z_{1},\cdot).$
Thus, the decay of $\omega_{U}(\cdot,D(x,r))$ when $r\to 0$ is comparable.
We start by recalling the following estimate on harmonic measure of disks for some particular simply connected domains, which follows from Beurling’s Projection Theorem [GM05, Thm. 9.2].
Theorem 2.24.
(Harmonic measure of disks I, [GM05, p. 281]) Let $U\subset\widehat{\mathbb{C}}$ be a simply connected domain, such that $\infty\in U$ and $\textrm{\em diam}(\partial U)=2$. Then, for all $x\in\partial U$ and $r>0$, it holds
$\omega_{U}(\infty,D(x,r))\leq\sqrt{r}.$
Next we prove that similar estimates hold for any arbitrary simply connected domain.
Lemma 2.25.
(Harmonic measure of disks II) Let $U$ be a simply connected domain, and let $z_{0}\in U$. Then, there exists $C\coloneqq C(U,z_{0})$ and $r_{0}\coloneqq r_{0}(U,z_{0})$ such that, for all $r\in(0,r_{0})$ and $x\in\widehat{\mathbb{C}}$,
$\omega_{U}(z_{0},D(x,r))\leq C\cdot\sqrt{r}.$
Without the explicit independence of $C$ on the point $x$, this statement can be found in [GM05, Corol. 9.3]. [^1] We are indebted to Phil Rippon for providing us the idea of this quantitative version of the estimate.
Proof of Lemma 2.25.
We start by proving the following claim, which asserts that similar bounds of Theorem 2.24 holds for any $x\in\widehat{\mathbb{C}}$, not only for $x\in\partial U$.
Claim.
Let $U\subset\widehat{\mathbb{C}}$ be a simply connected domain, such that $\infty\in U$ and $\textrm{\em diam}(\partial U)=2$. Then, for all $x\in\widehat{\mathbb{C}}$ and $r>0$, it holds
$\omega_{U}(\infty,D(x,r))\leq\sqrt{2r}.$
Proof.
Let $x\in\widehat{\mathbb{C}}$ and $r>0$. If $D(x,r)\cap\partial U=\emptyset$, the statement is clear. Otherwise, $D(x,r)\cap\partial U\neq\emptyset$. Then, there exists $y\in\partial U$, such that $\left|x-y\right|<r$. Hence $D(x,r)\subset D(y,2r)$, and
$\omega_{U}(\infty,D(x,r))\leq\omega_{U}(\infty,D(y,2r))\leq\sqrt{2r},$
as desired. $\square$
We shall consider a Möbius transformation $M$ sending $z_{0}$ to $\infty$ and such that $\textrm{diam}(M(\partial U))=2$, and then apply the previous claim. We construct the Möbius transformation $M$ as follows. First, if $z_{0}\neq\infty$, let
$M_{1}\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}},\quad M_{1}(z)=\frac{1}{z-z_{0}}.$
Otherwise, let $M_{1}=\textrm{id}_{\widehat{\mathbb{C}}}$. In either case, $M_{1}(\partial U)$ is a compact subset of $\mathbb{C}$, so there exists $x_{1},x_{2}\in\partial U$ such that
$\left|x_{1}-x_{2}\right|=\textrm{diam}(M_{1}(\partial U))\eqqcolon R.$
Then, let
$M_{1}\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}},\quad M_{2}(z)=\frac{2}{R}\left(z-\frac{x_{1}+x_{2}}{2}\right),$
and $M\coloneqq M_{2}\circ M_{1}$. Notice that $M(z_{0})=\infty$ and $\textrm{diam}(M(\partial U))=2$, as desired.
Now let $R_{0}\coloneqq\textrm{dist}(z_{0},\partial U)>0$, and $r_{0}\coloneqq R_{0}/2$. We note that both $M$ and $r_{0}$ depend only on the domain $U$ and the point $z_{0}$. Consider any $r\in(0,r_{0})$ and $x\in\partial U$. Then, if we write $M(D(x,r))=D(y,\rho)$ (recall that the image of an Euclidean disk under a Möbius transformation is still an Euclidean disk), we have
$\omega_{U}(z_{0},D(x,r))=\omega_{M(U)}(M(z_{0}),M(D(x,r)))=\omega_{M(U)}(\infty,D(y,\rho))\leq\sqrt{2\rho},$
by the previous claim. Hence, it is left to prove that there exists a constant $C>0$ such that $\rho\leq Cr$. Indeed,
$\rho\leq\max_{z\in\overline{D(x,r)}}\left|M^{\prime}(z)\right|\cdot r.$
But
$\left|M^{\prime}(z)\right|=\left|M_{2}^{\prime}(M_{1}(z))\right|\cdot\left|M^{\prime}_{1}(z)\right|\leq\frac{2}{R}\max\left\{1,\frac{1}{\left|z-z_{0}\right|^{2}}\right\}\leq\frac{2}{R}\max\left\{1,\frac{1}{(R_{0}-r)^{2}}\right\},$
since $z\in D(x,r)$, and hence $\left|z_{0}-z\right|\geq R_{0}-r$. Since $r<R_{0}/2$, we can take
$C\coloneqq\frac{8}{R\cdot R_{0}^{2}},$
which only depends on $R$ and $R_{0}$ (and hence on $z_{0}$ and $U$), as desired. This ends the proof of the lemma. $\square$
Estimates on the harmonic measure of sectors
Finally, we ask ourselves if the estimates obtained in the previous lemma can be improved under some additional assumptions on the shape of the domain. Note that the previous estimates work for any simply connected domain, and that they are sharp when the domain considered is a slit domain. Indeed, up to applying a Möbius transformation, we can assume our slit domain is $U=\mathbb{C}\smallsetminus\left(-\infty,0\right]$, and we want to estimate the harmonic measure of disks centered at 0.
[…]
Figure 2.1: The map $z\mapsto\sqrt{z}$ is well-defined and conformal in the slit domain $\mathbb{C}\smallsetminus\left(-\infty,0\right]$, and allows us to estimate the harmonic measure of disks around 0 (the turning point).
To do so, the natural procedure would be to consider a a conformal map from $\mathbb{D}$ to $U$ (which extends continuously to the boundaries), and estimate the Lebesgue measure of the push-forward of disks $D(0,r)$ under this map. To make computations easier, let us fix the basepoint 1, and consider instead a conformal map from the right half-plane $\mathbb{H}$ to $U$, fixing 1. The fact that the Möbius transformation $M\colon\mathbb{D}\to\mathbb{H}$ is bi-Lipschitz around the considered disks justifies that we can work with $\mathbb{H}$. Such map is the square root $\sqrt{\cdot}\colon\mathbb{C}\smallsetminus\left(-\infty,0\right]\to\mathbb{H}$. This way, one can show that the estimate in Lemma 2.25 is sharp (see Fig. 2.1).
Hence, for $\xi\in\partial\mathbb{D}$ and $x\in\partial U$, consider the set
$S_{\alpha,r}(x,\xi)=\left\{z\in{\mathbb{C}}\colon\left|z-x\right|<r,\left|\textrm{Arg }\xi-\textrm{Arg }(z-x)\right|<\pi\alpha\right\},$
which is a sector of angle $2\pi\alpha$, $\alpha\in(0,1)$, and side-length $r>0$, with vertex at $x\in{\partial}U$ (see Fig. 2.2). In the case that $x=\infty$, we define the sector with vertex at infinity in a natural way: as the image of a sector at a finite point under a Möbius transformation, that is
$S_{\alpha,R}(\infty,\xi)\coloneqq\left\{z\in\mathbb{C}\colon\left|z\right|>R,\ \left|\textrm{Arg }\xi-\textrm{Arg }(1/\overline{z})\right|<\pi\alpha\right\}.$
However, for the sake of simplicity, since Möbius transformations preserve harmonic measure, in the sequel we will assume, without loss of generality, that the vertex of a sector is always a finite point in the plane.
$z\mapsto z-x$$x$$r$$r$$\pi\alpha$$\xi$$S_{\alpha,r}(x,\xi)$$U$
Figure 2.2: A visual representation of the definition of the sector $S_{\alpha,r}(x,\xi)$.
Next, we assume that $U$ is contained in some sector around $x\in\partial U$ and we obtain improved estimates of harmonic measure for disks centered at $x$.
Lemma 2.26.
(Harmonic measure of sectors) Let $U\subset\mathbb{C}$ be a simply connected domain, and let $z_{0}\in U$, $x\in{\partial}U$. Assume there exists $r_{0}>0$, $\alpha\in\left(0,1\right)$ and $\xi\in\partial\mathbb{D}$, such that
$D(x,r_{0})\cap U\subset S_{\alpha,r_{0}}(x,\xi).$
Then, there exists $C>0$ and $r_{1}\in(0,r_{0})$ such that, for all $r\in(0,r_{1})$,
$\omega_{U}(z_{0},D(x,r))\leq C\cdot r^{\frac{1}{2\alpha}}.$
Proof.
Without loss of generality we assume $z_{0}\notin S_{\alpha,r_{0}}(x,\xi)$, and let $r\in(0,r_{0})$. First observe that, if $V$ denotes the connected component of $U\smallsetminus D(x,r)$ that contains $z_{0}$, we have that
$\omega_{U}(z_{0},D(x,r))=\omega_{U}(z_{0},D(x,r)\cap\partial U)\leq\omega_{V}(z_{0},\partial D(x,r)\smallsetminus\partial U)\leq\omega_{V}(z_{0},\partial D(x,r)),$
where in the first inequality we applied the Comparison Lemma 2.19 (note that we apply it to the complements, and hence the inequality is reversed), and the second follows from the inclusion of the measured sets.
Next we observe that, without loss of generality, we can assume that
$V\subset S_{\alpha}(x,\xi)=\left\{z\in{\mathbb{C}}\colon\left|\textrm{Arg }\xi-\textrm{Arg }(z-x)\right|<\pi\alpha\right\}.$
Indeed, since we want to estimate the harmonic measure of disks $D(x,r)$ centered at $x\in\partial U$ (which is a local property of the boundary around the point $x$), and $D(x,r_{0})\cap U\subset S_{\alpha,r_{0}}(x,\xi)$, for $r>0$ small enough, we can disregard the part of $\partial U$ outside $D(x,r_{0})$.
Therefore, up to composing by appropriate Möbius transformations, it is left to see that, if
$S=\left\{z\in\mathbb{C}\colon\left|\textrm{Arg }z\right|<\pi\alpha\right\},$
then, for some constant $C>0$, we have
$\omega_{S\smallsetminus D(0,r)}(1,\partial D(0,r))\leq C\cdot r^{\frac{1}{2\alpha}}.$
However, since the length of a circumference of radius $r$ is $2\pi r$ (i.e. proporcional to the radius), it is enough to see that $\omega_{S}(1,D(0,r))$ decays to 0 like $r^{\frac{1}{2\alpha}}$, when $r\to 0$ (see Fig. 2.3). But, since $M(z)=z^{\frac{1}{2\alpha}}$ is a conformal map from $S$ to the right half-plane fixing 1, this follows straighforward (see again Fig. 2.3)
$S\smallsetminus D(0,r)$$S$$\mathbb{H}$$r$$r$$r^{\theta}$$z\mapsto z^{\theta}$$\theta=1/2\alpha$
Figure 2.3: A visual scheme to approximate harmonic measure of sectors.
$\square$
3 On the setting and the hypothesis
In this section we shall discuss the setting we are working with (that is, iteration of functions of class $\mathbb{K}$, and the dynamics on the boundary of Fatou components), and the hypotheses we establish (specifically, thin singular values and order of growth in sectors). Note that the definitions of thin singular values and order of growth in sectors are stated here in a broader sense that in the introduction in order to deal with, not only with transcendental meromorphic functions, but also with functions in class $\mathbb{K}$.
3.1 Fatou components of functions of class $\mathbb{K}$
As mentioned in the introduction, consider $f\in\mathbb{K}$, i.e.
$f\colon\widehat{\mathbb{C}}\smallsetminus E(f)\to\widehat{\mathbb{C}},$
where $\Omega(f)\coloneqq\widehat{\mathbb{C}}\smallsetminus E(f)$ is the largest set where $f$ is meromorphic and $E(f)$ is the set of singularities of $f$, which is assumed to be closed and countable. Note that $\Omega(f)$ is open.
Notation. Once a function $f\in\mathbb{K}$ is fixed, we denote $\Omega(f)$ and $E(f)$ simply by $\Omega$ and $E$, respectively. Given a domain $U\subset\Omega$, we denote by $\partial U$ the boundary of $U$ in $\Omega$, and we keep the notation ^$\widehat{\partial}U$ for the boundary with respect to $\widehat{\mathbb{C}}$.
The dynamics of functions in class $\mathbb{K}$ was studied in [Bol96, Bol97, Bol99, BDH01, BDH04, Dom10, DMdOS22]. The Fatou set $\mathcal{F}(f)$ is defined as the largest open set in which $\left\{f^{n}\right\}_{n}$ is well-defined and normal, and the Julia set $\mathcal{J}(f)$, as its complement in $\widehat{\mathbb{C}}$. The standard theory of Fatou and Julia for rational or entire functions extends successfully to this more general setting. We shall need the following properties.
Theorem 3.1.
(Properties of Fatou and Julia sets, [BDH01, Thm. A]) Let $f\in\mathbb{K}$. Then,
  1. (a)
    $\mathcal{F}(f)$ is completely invariant in the sense that $z\in\mathcal{F}(f)$ if and only if $f(z)\in\mathcal{F}(f)$;
  2. (b)
    for every positive integer $k$, $f^{k}\in\mathbb{K}$, $\mathcal{F}(f^{k})=\mathcal{F}(f)$ and $\mathcal{J}(f^{k})=\mathcal{J}(f)$;
  3. (c)
    $\mathcal{J}(f)$ is perfect;
  4. (d)
    repelling periodic points are dense in $\mathcal{J}(f)$.
By (a), Fatou components (i.e. connected components of $\mathcal{F}(f)$) are mapped among themselves, and hence classified into periodic, preperiodic or wandering. By (b), the study of periodic Fatou components reduces to the invariant ones, i.e. those for which $f(U)\subset U$.
Theorem 3.2.
(Connectivity of Fatou components, [Bol99]) Let $f\in\mathbb{K}$, and let $U$ be a periodic Fatou component of $f$. Then, the connectivity of $U$ is 1, 2, or $\infty$.
In this paper, we focus on simply connected periodic Fatou components, which we assume to be invariant. Those Fatou components are classified into attracting basins, parabolic basins, Siegel disks, and Baker domains [BDH01, Thm. C]. A Baker domain is, by definition, a periodic Fatou component $U$ of period $k\geq 1$ for which there exists $z_{0}\in\widehat{\partial}U$ such that $f^{nk}(z)\to z_{0}$, for all $z\in U$ as $n\to\infty$, but $f^{k}$ is not meromorphic at $z_{0}$. In such case, $z_{0}$ is accessible from $U$ [BDH01, 658]. Baker domains are classified according to its internal dynamics in doubly parabolic, hyperbolic and simply parabolic (see e.g. [FH06], also Sect. 8). For some doubly parabolic Baker domains, $f|_{\partial U}$ is recurrent with respect to harmonic measure, and we call them recurrent Baker domains.
Recall that, since ergodicity and recurrence depend only on sets of zero and full measure (see Def. 2.7), these properties are well-defined for harmonic measure without specifying the basepoint.
To prove Theorems A and B we shall need the following ergodic results about Fatou components, which collect essentially the work of Doering and Mañé [DM91], with some refinements in [BFJK19] and [Jov24]. The curious reader can find more details in Sections 7 and 8 (Theorems 7.6 and 8.3, and Corollary 8.4), which are interesting on its own, and needed to prove density of periodic boundary points (Corollary D).
Theorem 3.3.
(Ergodic properties of $f|_{\partial U}$) Let $f\in\mathbb{K}$, and let $U$ be an attracting or parabolic basin, or a recurrent Baker domain. Then, $f|_{\partial U}$ is ergodic and recurrent with respect to the harmonic measure $\omega_{U}$. Moreover,
  1. (i)
    if $U$ is an attracting basin with fixed point $p\in U$, the harmonic measure $\omega_{U}(p,\cdot)$ is invariant under $f$;
  2. (ii)
    if $U$ is a parabolic basin or a recurrent Baker domain, with convergence point $p\in\widehat{\partial}U$, let $\varphi\colon\mathbb{H}\to U$ be a Riemann map with $\varphi^{*}(\infty)=p$, and consider the Lebesgue measure $\lambda_{\mathbb{R}}$ in $\mathbb{R}$. Then, the push-forward
    $\mu\coloneqq(\varphi^{*})_{*}\lambda_{\mathbb{R}},$
    is invariant under $f$. Moreover, $\mu$ and $\omega_{U}$ are mutually absolutely continuous.
Regular and singular values for holomorphic maps
Throughout the paper, we will make an extensive use of the concepts of regular and singular values, in the sense of [Ive14, BE95], for both functions of class $\mathbb{K}$ and inner functions. We review the basic definitions.
We consider the following class of meromorphic functions, denoted by $\mathbb{M}$, consisting of functions
$f\colon\widehat{\mathbb{C}}\smallsetminus E(f)\longrightarrow\widehat{\mathbb{C}},$
where ^$\widehat{\mathbb{C}}\smallsetminus E(f)$ is the largest set where $f$ is meromorphic, and, for all $z\in E(f)$, the cluster set $Cl(f,z)$ of $f$ at $z$ is $\widehat{\mathbb{C}}$. Note that functions in class $\mathbb{K}$ are in $\mathbb{M}$, as well as inner functions (see e.g. [BD99, BDH01], and also the discussion in [Jov24, Sect. 2.6]).
In this general setting, regular and singular values, and critical and asymptotic values, are defined as follows. Note that appropriate charts have to be used when dealing with $\infty$.
Definition 3.4.
(Regular and singular values) Given $v\in\widehat{\mathbb{C}}$, we say that $v$ is a regular value for $f$ if there exists $r\coloneqq r(v)>0$ such that all branches $F_{1}$ of $f^{-1}$ are well-defined (and hence, conformal) in $D(v,r)$. Otherwise we say that $v$ is a singular value for $f$.
The set of singular values of $f$ is denoted by $SV(f)$. Note that $SV(f)$ is closed by definition, and it is the smallest set for which
$f\colon\widehat{\mathbb{C}}\smallsetminus(E(f)\cup f^{-1}(SV(f)))\longrightarrow\widehat{\mathbb{C}}\smallsetminus SV(f)$
is a covering map.
Definition 3.5.
(Critical and asymptotic values) Given $v\in\widehat{\mathbb{C}}$, we say that $v$ is a critical value if there exists $z\in\Omega$ such that $f^{\prime}(z)=0$ and $f(z)=v$. We say that $z$ is a critical point.
We say that $v$ is an asymptotic value if there exists a curve $\gamma\colon\left[0,1\right)\to\Omega$ such that $\gamma(t)\to\partial\Omega$ and $f(\gamma(t))\to v$, as $t\to 1$.
The set of critical values of $f$ is denoted by $CV(f)$, while $AV(f)$ stands for the set of asymptotic values. It holds that singular values are precisely the critical values, the asymptotic values, and the accumulation thereof.
Lemma 3.6.
(Characterization of singular values, [Jov24, Lemma 2.21]) Let $f\in\mathbb{M}$. Then,
$SV(f)=\overline{CV(f)\cup AV(f)}.$
Finally, recall that we define the postsingular set of $f$ as
$P(f)\coloneqq\overline{\bigcup\limits_{s\in SV(f)}\bigcup\limits_{n\geq 0}f^{n}(s)}.$
3.2 Thin singular values in $\partial U$
In this section we discuss the hypothesis on the singular values used in Theorem A and Proposition F. The general idea is to allow infinitely many singular values in ^$\widehat{\partial}U$, as long as they cluster together in a controlled way (geometrically).
Without loss of generality, by conjugating the map $f$ by a Möbius transformation, we shall assume $\infty\in U$, hence ^$\widehat{\partial}U$ is a compact subset of the plane and we can work with the Euclidean distance. Alternatively, we could have worked with the spherical metric in $\widehat{\mathbb{C}}$, but we choose to use the Euclidean one to simplify the notation and the computations. Recall that we denote the Euclidean distance by dist.
Definition 3.7.
(Separated sets) Let $X$ be a compact subset of $\mathbb{C}$, endowed with the Euclidean distance. We say that a subset $K\subset X$ is $\delta$-separated if for all $x,y\in K$, $x\neq y$, we have $\textrm{dist}(x,y)\geq\delta$.
Definition 3.8.
(Thin singular values) Let $f\in\mathbb{K}$, and let $U$ be a Fatou component. We say that singular values are thin in ^$\widehat{\partial}U$ if there exists $\varepsilon>0$, $\mu\in(0,1)$, $d\in\mathbb{N}$ and $v_{1},\dots,v_{k}\in\widehat{\partial}U$ such that
  1. (a)
    for every $n\in\mathbb{N}$, there are at most $n^{d}$ singular values in
    $\left\{z\in\widehat{\mathbb{C}}\colon\textrm{dist}(z,\widehat{\partial}U)<\mu^{n}\right\}\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$
    which are $\mu^{n}$-separated;
  2. (b)
    for $i=1,\dots,k$, there exists $\delta_{i}$ such that, if $M_{i}$ is a Möbius transformation with
    $M_{i}(D(v_{i},\varepsilon))=\widehat{\mathbb{C}}\smallsetminus\overline{D(0,\varepsilon)}\quad\textrm{and}\quad M_{i}(v_{i})=\infty,$
    then
    $\textrm{dist}(M_{i}(\widehat{\partial}U\cap D(v_{i},\varepsilon)),M_{i}(SV(f)\cap D(v_{i},\varepsilon)))>\delta_{i}.$
We say that $\left\{v_{1},\dots,v_{k}\right\}$ is the set of punctures, and we put
$\mathcal{S}\coloneqq SV(f)\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon).$
$U$$M_{i}(U)$$M_{i}$$\varepsilon$$v_{i}$
Figure 3.1: Schematic representation of the notion of thin singular values in the boundary of Fatou components. Indeed, there are a finite number of punctures $v_{1},\dots,v_{k}$ (red), such that there are at most $n^{d}$ singular values in $\left\{z\in\widehat{\mathbb{C}}\colon\textrm{dist}(z,\widehat{\partial}U)<\mu^{n}\right\}\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$ which are $\mu^{n}$-separated (pink region). Moreover, in the disk $D(v_{i},\varepsilon)$ (grey) singular values lie at a bounded distance from the boundary when applying the Möbius transformation $M_{i}$ (right).
Intuitively, outside from the disks $D(v_{i},\varepsilon)$, $i=1,\dots,k$, singular values cluster to their accumulation points exponentially fast, and in the disks $D(v_{i},\varepsilon)$ they remain at a bounded distance of the boundary.
Observe that we allow infinitely many singular values in ^$\widehat{\partial}U$, as long as their distribution satisfies the metric conditions specified above. Without loss of generality, we shall make the following assumptions.
  • Choosing $\delta\coloneqq\min(\delta_{1},\dots,\delta_{k})>0$, we can assume $\delta$ is uniform for all $i=1,\dots,k$.
  • Choosing $\varepsilon$ small enough, we can assume that the disks $D(v_{1},\varepsilon),\dots,D(v_{k},\varepsilon)$ are pairwise disjoint and the Euclidean distance between any such two disk is greater than $\varepsilon$. In fact, if there exists $\varepsilon_{0}>0$ for which the conditions in Definition 3.8 are satisfied, then there are fulfilled for any $\varepsilon\in(0,\varepsilon_{0})$.
    Indeed, assume condition (a) is satisfied for some $\varepsilon_{0}>0$, and let $\varepsilon\in(0,\varepsilon_{0})$. Then, there exists $n_{0}\in\mathbb{N}$ large enough so that, if
    $A(v_{i},\varepsilon,\varepsilon_{0})\coloneqq\left\{z\in{\mathbb{C}}\colon\varepsilon<\left|z-v_{i}\right|<\varepsilon_{0}\right\},$
    then there are no singular values in
    $\left\{z\in\mathbb{C}\colon\textrm{dist}(z,\widehat{\partial}U)<\mu^{n_{0}}\right\}\cap\bigcup_{i=1}^{k}A(v_{i},\varepsilon,\varepsilon_{0}).$
    (Note that singular values in any compact subset of $D(v_{i},\varepsilon_{0})\smallsetminus\left\{v_{i}\right\}$ lie at a positive Euclidean distance of $\partial U$).
    Hence, for $n\geq n_{0}$, there are at most $n^{d}$ singular values in
    $\left\{z\in\widehat{\mathbb{C}}\colon\textrm{dist}(z,\widehat{\partial}U)<\mu^{n_{0}}\right\}\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$
    which are $\mu^{n}$-separated.
    It is trivial to see that, if condition (b) holds for some $\varepsilon_{0}$, then it holds for any $\varepsilon$ smaller.
  • We shall choose a particular form for the Möbius transformations $M_{i}$. Imposing that $M_{i}(v_{i}+\varepsilon)=\varepsilon$, it follows that
    $M_{i}(z)=\dfrac{\varepsilon^{2}}{z-v_{i}},$
    with derivative
    $M^{\prime}_{i}(z)=-\dfrac{\varepsilon^{2}}{(z-v_{i})^{2}}\quad\textrm{and}\quad\left|M^{\prime}_{i}(z)\right|=\dfrac{\varepsilon^{2}}{\left|z-v_{i}\right|^{2}}.$
  • Finally, note that the notion of singular values being thin in ^$\widehat{\partial}U$ is invariant under conjugating $f$ by a Möbius transformation $M$. Indeed, $M^{\prime}$ is uniformly bounded around ^$\widehat{\partial}U$ and hence distances are distorded in a controlled way when applying $M$.
In this situation, neither the Euclidean metric nor the spherical are appropriate to work with, basically because they do not give a differentiated role to the points $\left\{v_{1},\dots,v_{k}\right\}$. Hence, we shall define the following metric, which we call the $\rho$-metric and for which the points $\left\{v_{1},\dots,v_{k}\right\}$ are punctures: they are at infite distance of any other point.
Definition 3.9.
(The $\rho$-metric) Let $U$ be a Fatou component, and assume singular values are thin in ^$\widehat{\partial}U$. Then, we define the $\rho$-metric in as
$d\rho(z)=\rho(z)\left|dz\right|,$
where
$\rho\colon\widehat{\mathbb{C}}\smallsetminus\left\{v_{1},\dots,v_{k}\right\}$ $\longrightarrow\left[0,+\infty\right)$
$z$$\longmapsto\left\{\begin{array}[]{lcc}&\dfrac{\varepsilon^{2}}{\left|z-v_{i}\right|^{2}}&\textrm{if }z\in D(v_{i},\varepsilon),\\ \\ &1&\textrm{otherwise.}\end{array}\right.$
Observe that $\rho$ is a continuous function, and hence defines a conformal metric on ^$\widehat{\mathbb{C}}\smallsetminus\left\{v_{1},\dots,v_{k}\right\}$. Hence, one has to interpret this new metric $\rho$ as the gluing of the Euclidean metric in places which are not problematic (that is, in ^$\widehat{\mathbb{C}}\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$), and the “Euclidean metric around infinity” in the neighbourhoods of the problematic points $\left\{v_{1},\dots,v_{k}\right\}$. Indeed, for $i=1,\dots,k$ and $z,w\in D(v_{i},\varepsilon)$,
$\textrm{dist}_{\rho}(z,w)=\textrm{dist}(M_{i}(z),M_{i}(w)).$
Note that the $\rho$-metric is not defined at the punctures $\left\{v_{1},\dots,v_{k}\right\}$. We devote the rest of the section to prove several properties of the $\rho$-metric.
First note that the punctures are singular values, since they are accumulated by other singular values. However, $f$ has, in general, more singular values than the punctures $\left\{v_{1},\dots,v_{k}\right\}$. Nevertheless, these singular values either lie uniformly far (with respect to the $\rho$-metric) from the boundary of $U$, or they satisfy the condition of being $\mu^{n}$-separated. In the later case, there is a great control on harmonic measure around such points. This is the content of the following lemma.
Lemma 3.10.
($\rho$-distribution of singular values) Let $f\in\mathbb{K}$, and let $U$ be a Fatou component for $f$. Assume that singular values are thin in ^$\widehat{\partial}U$, and consider the $\rho$-metric defined above. Then, there exists $\eta>0$ such that there are no singular values in
$\bigcup_{i=1}^{k}D(v_{i},\varepsilon)\cap\left\{z\in\widehat{\mathbb{C}}\colon\textrm{\em dist}_{\rho}(z,\widehat{\partial}U)<\eta\right\}.$
Moreover, if
$\mathcal{S}\coloneqq SV(f)\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon),$
then, for any $C>0$ and $t\in(0,1)$,
$\sum_{n\geq 0}\omega_{U}\left(\bigcup\limits_{s\in\mathcal{S}}D(s,C\cdot t^{n})\right)<\infty.$
Remark 3.11.
(Singular values have zero $\omega_{U}$-measure) Note that Lemma 3.10 already implies that $SV(f)\cap\widehat{\partial}U$ have zero harmonic measure. Indeed,
$SV(f)\cap\widehat{\partial}U\subset\mathcal{S}\cup\left\{v_{1},\dots,v_{k}\right\},$
and the convergence of the series $\sum_{n\geq 0}\omega_{U}(\bigcup\limits_{s\in\mathcal{S}}D(s,C\cdot t^{n}))$ ensures that $SV(f)\cap\widehat{\partial}U$ have zero harmonic measure.
Proof of Lemma 3.10.
First we shall see that there exists $\eta\in(0,\varepsilon)$ such that there are no singular values in
$\left\{z\in\widehat{\mathbb{C}}\colon\textrm{dist}_{\rho}(z,\widehat{\partial}U)<\eta\right\},$
for $z\in\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$. By assumption, there exists $\delta>0$ such that for all $x\in\widehat{\partial}U\cap D(v_{i},\varepsilon)$ and $s\in SV(f)\cap D(v_{i},\varepsilon)$, we have
$\textrm{dist}_{\rho}(x,v)=\textrm{dist}(M_{i}(x),M_{i}(v))>\delta.$
Choosing $\eta\coloneqq\min(\delta,\varepsilon)$ we proved the first statement of the Lemma.
It is left to prove the second claim. First note that, without loss of generality, we can assume $C=1$. Indeed, for any $s\in(t,1)$, there exists $n_{0}\in\mathbb{N}$ such that $Ct^{n}\leq s^{n}$, for all $n\geq n_{0}$.
Assume first that $t\leq\mu$. Then, for all $n\in\mathbb{N}$, there exist $s^{n}_{1},\dots,s^{n}_{n^{d}}\in\mathcal{S}$ such that
$D_{n}\coloneqq\bigcup\limits_{s\in\mathcal{S}}D(s,t^{n})\subset\bigcup\limits_{s\in\mathcal{S}}D(s,\mu^{n})\subset\bigcup\limits_{i=1}^{n^{d}}D(s^{n}_{i},2\mu^{n}).$
Then, applying the estimate of harmonic measure given in Lemma 2.25, there exists a constant $C>0$ such that
$\sum_{n\geq 0}\omega_{U}(D_{n})\leq\sum_{n\geq 0}\omega_{U}\left(\bigcup\limits_{i=1}^{n^{d}}D(s^{n}_{i},2\mu^{n})\right)\leq\sum_{n\geq 0}n^{d}\cdot C\cdot\mu^{n/2}=C\cdot\sum_{n\geq 0}n^{d}\cdot\mu^{n/2}.$
To see that the last series converges, apply the ratio test: if $a_{n}=n^{d}\cdot\mu^{n/2}$
$\lim\limits_{n}\dfrac{a_{n+1}}{a_{n}}=\lim\limits_{n}\dfrac{(n+1)^{d}\cdot\mu^{(n+1)/2}}{n^{d}\cdot\mu^{n/2}}=\mu^{1/2}<1,$
as desired.
Now assume $t>\mu$. Since there are at most $n^{d}$ singular values in $\mathcal{S}$ which are $\mu^{n}$-separated, there exist $s^{n}_{1},\dots,s^{n}_{n^{d}}\in\mathcal{S}$ such that
$D_{n}\coloneqq\bigcup\limits_{s\in\mathcal{S}}D(s,t^{n})\subset\bigcup\limits_{i=1}^{n^{d}}D(s^{n}_{i},2t^{n}).$
Then the proof ends applying the same arguments as before. $\square$
Finally, we give the following estimate, which provides a quantitative way of comparing the $\rho$-metric with the Euclidean one.
Lemma 3.12.
(Euclidean versus $\rho$-metric) Let $x\in D(v_{i},\varepsilon)$. If $v_{i}\notin D(x,2r)$, then
$D_{\rho}(x,r)\subset D(x,r)\subset D_{\rho}(x,16\cdot r\cdot\rho(x)).$
Proof.
The first inclusion is clear since $\rho(z)\geq 1$, for all $z\in\mathbb{C}$. The second inclusion comes from applying Koebe’s Distortion Theorem 2.1. Indeed, for $z\in D(x,r)$,
$\textrm{dist}_{\rho}(z,x)=\min\left\{\textrm{dist}(z,x),\textrm{dist}(M_{i}(z),M_{i}(x))\right\}.$
Note that $M_{i}$ is univalent in $D(x,2r)$, so for $z\in D(x,r)$,
$\textrm{dist}(M_{i}(z),M_{i}(x))\leq\frac{2\cdot\left|M_{i}^{\prime}(x)\right|}{(1-1/2)^{3}}\cdot\textrm{dist}(z,x)=16\cdot\rho(x)\cdot r.$
Since $\rho(x)\geq 1$, it follows that
$\textrm{dist}_{\rho}(z,x)=\min\left\{\textrm{dist}(z,x),\textrm{dist}(M_{i}(z),M_{i}(x))\right\}\leq 16\cdot\rho(x)\cdot r,$
as desired. $\square$
3.3 Order of growth in sectors
Since we are working in the general context of functions of class $\mathbb{K}$, we shall define the following notion of order of growth, which allow derivatives to blow up exponentially fast in a finite numbers of points (singularities) on the boundary of the Fatou component.
To simplify the notation, in the sequel we shall assume $\infty\in U$, hence ^$\widehat{\partial}U$ is a compact subset of the plane, and that none of the singularities is placed at $\infty$. This can be achieved by conjugating $f$ by an appropriate Möbius transformation. Alternatively, one can work with the spherical metric, but the notation is more involved.
Definition 3.13.
(Order of growth in sectors) Let $f\in\mathbb{K}$, and let $U\subset\widehat{\mathbb{C}}$ be an invariant Fatou component for $f$. We say that $U$ is asymptotically contained in a sector of angle $\alpha\in\left(0,1\right)$ with order of growth $\beta>0$ if there exists $r_{0}>0$, $s_{1},\dots,s_{k}\in\widehat{\partial}U\smallsetminus\left\{\infty\right\}$ and $\xi_{1},\dots,\xi_{k}\in\partial\mathbb{D}$, such that, if
$S_{\alpha,r}=\bigcup\limits_{i=1}^{k}\left\{z\in{\mathbb{C}}\colon\left|z-s_{i}\right|<r,\left|\textrm{Arg }\xi_{i}-\textrm{Arg }(z-s_{i})\right|<\pi\alpha\right\}$
satisfies
  1. (a)
    $\overline{U}\cap\bigcup\limits_{i=1}^{k}D(s_{i},r_{0})\subset S_{\alpha,r_{0}}$;
  2. (b)
    $f$ has order of growth $\beta>0$ in $S_{\alpha,r_{0}}$, i.e. there exist $A,B>0$ such that, for all $r<r_{0}$ and $z\in S_{\alpha,r_{0}}\smallsetminus S_{\alpha,r}$,
    $A\cdot e^{B\cdot r^{\beta}}\leq\left|f^{\prime}(z)\right|\leq A\cdot e^{B\cdot r^{-\beta}}.$
Geometrically, each of the sets
$S_{i}=\left\{z\in{\mathbb{C}}\colon\left|z-s_{i}\right|<r,\left|\textrm{Arg }\xi_{i}-\textrm{Arg }(z-s_{i})\right|<\alpha\right\}$
is a sector of angle $\alpha\in(0,\pi)$, and side-length $r>0$, with vertex at $s_{i}\in\widehat{\partial}U$ (as considered in Sect. 2.6).
Note that this notion of order of growth in sectors is invariant under conjugating $f$ by a Möbius transformation $M$, as long as $M(U)\subset\mathbb{C}$ and $M(s_{i})\neq\infty$, $i=1,\dots,k$. Indeed, $M^{\prime}$ is uniformly bounded around ^$\widehat{\partial}U$ and hence distances are distorded in a controlled way when applying $M$.
4 Lyapunov exponents for transcendental maps
We define Lyapunov exponents as follows.
Definition 4.1.
(Lyapunov exponent) Let $f\in\mathbb{K}$, and let $X\subset\widehat{\mathbb{C}}$. Let $\mu$ be a measure supported on $X$, and assume $\log\left|f^{\prime}\right|\in L^{1}(\mu)$. Then,
$\chi_{\mu}(f)\coloneqq\int_{\partial U}\log\left|f^{\prime}\right|d\mu$
is called the Lyapunov exponent of the map $f$ (with respect to the measure $\mu$).
When $f$ is clear from the context, we shall write simply $\chi_{\mu}$ to lighten the notation. In the case when $\mu$ is an $f$-invariant ergodic probability measure, Birkhoff Ergodic Theorem 2.10 yields
$\chi_{\mu}(f)=\lim\limits_{n}\frac{1}{n}\sum_{k=0}^{n-1}\log\left|f^{\prime}(f^{k}(x))\right|,$
for $\mu$-almost every $x\in X$ [^2] We note that it is customary to define the Lyapunov exponent $\chi_{\mu}(f)$ as the previous limit, and then prove that it is equal, by Birkhoff Ergodic Theorem, to $\int_{\partial U}\log\left|f^{\prime}\right|d\mu$. In our setting, as we will see, it is more convenient to proceed the way we chosen, following the approach of [Prz85, p. 162], [PU10, Def. 9.1.2].
We are interested in the particular case where $X$ is the boundary of an invariant Fatou component $U$. If $U$ is an attracting basin, one shall consider the harmonic measure $\omega_{U}$ with basepoint the attracting fixed point. By Theorem 8.3 and Corollary 8.4, such a measure is a $f$-invariant ergodic probability, and by the previous argument, $\chi_{\omega_{U}}$ controls the asymptotic expansion of $f$ along $\omega_{U}$-almost every orbit in ^$\widehat{\partial}U$.
Note that we need to assume $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$. It is well-known that this holds for rational maps [Prz85]. In Section 4.1, we prove integrability of $\log\left|f^{\prime}\right|$ with respect to $\omega_{U}$, under some assumptions on the shape of the Fatou component and the growth of the function. We note that, given an explicit example, it may be possible to get better estimates than the ones we give in general, using an explicit computation of the derivatives or better approximations for the harmonic measure. In Section 4.2 we give conditions under which Lyapunov exponents are non-negative. Again, this is well-known for rational maps [Prz93], but unexplored in the transcendental case. Finally, Section 4.3 is devoted to extend some of the results to parabolic basins and Baker domains.
4.1 Integrability of $\log\left|f^{\prime}\right|$. Proposition E
Next we examine the integrability of $\log\left|f^{\prime}\right|$ with respect to harmonic measure. First observe that, by Harnack’s inequality, if $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U}(p,\cdot))$, then $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U}(q,\cdot))$, for all $q\in U$. Hence, in this case, we write simply $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$.
To begin with, we prove that the integrability of $\log\left|f^{\prime}\right|$ and the Lyapunov exponent is invariant under conjugating $f$ by Möbius transformations. This is the content of the following lemma, which follows from [Prz85, p. 165]. We include its proof for completeness.
Lemma 4.2.
(Lyapunov exponent invariant under Möbius transformations) Let $f\in\mathbb{K}$, and let $U$ be an invariant Fatou component for $f$. Let $M\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be a Möbius transformation, and let $g\in\mathbb{K}$ be defined as $g\coloneqq M\circ f\circ M^{-1}$. Then, $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$ if and only if $\log\left|g^{\prime}\right|\in L^{1}(\omega_{M(U)})$. Moreover, if $\omega_{U}$ is $f$-invariant, then $\omega_{M(U)}$ is $g$-invariant, and
$\chi_{\omega_{U}}(f)=\chi_{\omega_{M(U)}}(g).$
To proof Lemma 4.2, we need the following version of Löwner’s lemma. Although it is not exactly the statement in [DM91, Corol. 1.5(b)], [BEF+23, Thms. 2.6, 2.7], it can be deduced straighforward from them, taking into account that the associated self-map of $\mathbb{D}$ is an inner function (see Sect. 8).
Lemma 4.3.
(Löwner’s lemma for Fatou components) Let $f\in\mathbb{K}$, and let $U$ be an invariant Fatou component for $f$. Let $z\in U$, and consider the harmonic measure $\omega_{U}(z,\cdot)$. Then, for all $B\in\mathcal{B}(\widehat{\mathbb{C}})$ it holds
$\omega_{U}(z,f^{-1}(B))=\omega_{U}(f(z),B).$
Proof of Lemma 4.2.
By the previous observation and the bijectivity of Möbius transformations, it is enough to prove that, for a given $p\in U$ and $q=M(p)\in M(U)$, if $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U}(p,\cdot))$, then $\log\left|g^{\prime}\right|\in L^{1}(\omega_{U}(q,\cdot))$.
Any Möbius transformation $M\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$, with derivative $M^{\prime}$, can be writen explicitly as
$M(z)=\dfrac{az-b}{cz-d},\hskip 14.22636ptM^{\prime}(z)=\dfrac{bc-ad}{(cz-d)^{2}}.$
Note that, for any such a Möbius transformation, $\log\left|M^{\prime}\right|\in L^{1}(\omega_{U})$. Indeed,
[…]
$\leq\left|\log\left|bc-ad\right|\right|+2\int_{\partial U}\left|\log\left|cx-d\right|\right|d\omega_{U}(p,x).$
The last integral is finite due to Lemma 2.22.
Next, on the one hand, let $\varphi\colon\mathbb{D}\to U$ be a Riemann map, such that $\varphi(0)=p$. Then, $M\circ\varphi\colon\mathbb{D}\to M(U)$ is a Riemann map for $M(U)$ with $M\circ\varphi(0)=M(p)=q$, and
[…]
On the other hand, since $g=M\circ f\circ M^{-1}$, it follows that
$g^{\prime}(M(z))=M^{\prime}(f(z))\cdot f^{\prime}(z)\cdot M^{-1}(M(z))=M^{\prime}(f(z))\cdot f^{\prime}(z)\cdot\frac{1}{M^{\prime}(z)}.$
Thus,
$\int_{\partial\mathbb{D}}\left|\log\left|g^{\prime}\circ\varphi^{*}\right|\right|d\lambda\leq\int_{\partial\mathbb{D}}\left|\log\left|M^{\prime}(f(\varphi^{*}))\right|\right|d\lambda+\int_{\partial\mathbb{D}}\left|\log\left|f^{\prime}(\varphi^{*})\right|\right|d\lambda+\int_{\partial\mathbb{D}}\left|\log\left|M^{\prime}(\varphi^{*}))\right|\right|d\lambda.$
Note that $\log\left|M^{\prime}\circ\varphi^{*}\right|,\ \log\left|f^{\prime}\circ\varphi^{*}\right|\in L^{1}(\lambda)$. Moreover, applying Löwner’s lemma 4.3, it follows that $\log\left|M^{\prime}\circ f\circ\varphi^{*}\right|\in L^{1}(\lambda)$. Indeed, Löwner’s lemma implies that
[…]
Therefore, $\log\left|g^{\prime}\circ\varphi^{*}\right|\in L^{1}(\lambda)$, as desired. The other implication follows by symmetry.
For the second statement, notice that, for any Borel set $A\subset\widehat{\mathbb{C}}$,
$\omega_{M(U)}(q,g^{-1}(A))$ $=\omega_{U}(p,M^{-1}(g^{-1}(A)))=\omega_{U}(p,f^{-1}(M^{-1}(A)))$
$=\omega_{U}(p,M^{-1}(A))=\omega_{M(U)}(q,A),$
and hence $\omega_{M(U)}$ is $g$-invariant. Then, proceeding as before, write
$\int_{\partial\mathbb{D}}\log\left|g^{\prime}(M(\varphi^{*}))\right|d\lambda=\int_{\partial\mathbb{D}}\log\left|M^{\prime}(f(\varphi^{*}))\right|d\lambda+\int_{\partial\mathbb{D}}\log\left|f^{\prime}(\varphi^{*})\right|d\lambda-\int_{\partial\mathbb{D}}\log\left|M^{\prime}(\varphi^{*}))\right|d\lambda.$
Since $\omega_{U}$ is assumed to be $f$-invariant and $\log\left|M^{\prime}\circ\varphi^{*}\right|\in L^{1}(\lambda)$, it follows that
$\int_{\partial\mathbb{D}}\left|\log\left|M^{\prime}(f(\varphi^{*}))\right|\right|d\lambda=\int_{\partial\mathbb{D}}\left|\log\left|M^{\prime}(\varphi^{*}))\right|\right|d\lambda,$
implying that
[…]
as desired. $\square$
Next we check that, if $U$ is asymptotically contained in a sector of angle $\alpha\in(0,1)$, with order of growth $\beta\in(0,1/2\alpha)$, then $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$.
Observe that, for rational maps, $\left|f^{\prime}\right|$ is bounded, so $\chi_{\mu}(f)$ is well-defined (although a priori may be equal to $\infty$- ∞). By a careful study of $f$ around critical points, it is established that it is never the case, and in fact Lyapunov exponents are always non-negative ([Prz93], see also [URM23, Sect. 28.1]). In the case of transcendental maps, $\left|f^{\prime}\right|$ may not be bounded, and this is why we need the assumption on the growth.
Proposition 4.4.
($\log\left|f^{\prime}\right|$ is $\omega_{U}$-integrable) Let $f\in\mathbb{K}$, and let $U$ be an invariant Fatou component for $f$. Assume $U$ is asymptotically contained in a sector of angle $\alpha\in(0,1)$, with order of growth $\beta\in(0,\frac{1}{2\alpha})$. Then, $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$.
Note that Proposition 4.4 implies Proposition E.
Proof.
By conjugating by a Möbius transformation if needed, we can assume $\infty\notin U$. Recall that the integrability of $\log\left|f^{\prime}\right|$ is preserved by conjugating by Möbius transformation (Lemma 4.2), as well as the hypothesis of $U$ being asymptotically contained in a sector of angle $\alpha$, with order of growth $\beta$.
First note that $\log\left|f^{\prime}\right|$ is integrable with respect to harmonic measure when restricted to compact subsets of the domain ^$\widehat{\mathbb{C}}\smallsetminus\left\{s_{1},\dots,s_{k}\right\}$. Indeed, the only difficulty is to see that $\log\left|f^{\prime}\right|$ is integrable around critical points. It is easy to check this by considering the Taylor expansion of $f$ around the critical point, and using that $\log\left|z-a\right|$ is integrable with respect to $\omega_{U}$ for all $a\in\mathbb{C}$ (Lemma 2.22).
It is left to check integrability near the singularities, and here is where we use the estimates on the growth. Let us use the notation
$\log^{+}\left|f^{\prime}(z)\right|\coloneqq\max(0,\log\left|f^{\prime}(z)\right|),\hskip 19.91684pt\log^{-}\left|f^{\prime}(z)\right|\coloneqq-\min(0,\log\left|f^{\prime}(z)\right|),$
so that
$\left|\log\left|f^{\prime}\right|\right|=\log^{+}\left|f^{\prime}\right|+\log^{-}\left|f^{\prime}\right|.$
Since $\log^{+}\left|f^{\prime}\right|$ and $\log^{-}\left|f^{\prime}\right|$ satisfy analogous estimates on their growth, we will check only that $\log^{+}\left|f^{\prime}\right|\in L^{1}(\omega_{U})$.
Since we are assuming that $U$ is asymptotically contained in a sector of angle $\alpha\in(0,1)$, with order of growth $\beta\in(0,\frac{1}{2\alpha})$, we have that
$\log^{+}\left|f^{\prime}(z)\right|\leq\log A+Br^{-\beta},$
for all $z\notin S_{\alpha,r}$. Note that
$\max_{z\notin D(s_{i},1/n)}\log^{+}\left|f^{\prime}\right|\leq\log A+Bn^{\beta}.$
Hence, the estimate $n^{\beta}$ holds except in the disk $D(s_{i},\frac{1}{n})$. Doing the same for $n+1$, we have that the estimate $(n+1)^{\beta}$ holds except in the disk $D(s_{i},\frac{1}{n+1})$. Therefore, for all $n$, the estimate $n^{\beta}$ is the best possible in the annulus
$A_{i,n}\coloneqq\left\{z\in\mathbb{C}\colon 1/(n+1)<\left|z\right|<1/n\right\},$
for which
$\omega_{U}(A_{i,n})\leq\omega_{U}(D(s_{i},1/n))-\omega_{U}(D(s_{i},1/n+1))\leq\dfrac{1}{n^{1/2\alpha}}-\omega_{U}(D(s_{i},1/n+1)),$
applying the estimates for the harmonic measure on sectors (Lemma 2.26). Although we do not have lower bounds for $\omega_{U}(D(s_{i},1/n+1))$, we do know that the upper estimate $n^{\beta}$ holds for all points in $D(s_{i},\frac{1}{n})$ except for those in $D(s_{i},1/n+1)$, that is, except in a set of harmonic measure at most $\frac{1}{(n+1)^{1/2\alpha}}$. Hence, we shall use the estimate $n^{\beta}$ in a set of harmonic measure at most
$\dfrac{1}{n^{1/2\alpha}}-\dfrac{1}{(n+1)^{1/2\alpha}}.$
Thus, writing $\gamma=\frac{1}{2\alpha}$, we can bound the previous sum by
$\sum_{n=0}^{\infty}\dfrac{1}{n^{-\beta}}\cdot\left(\dfrac{1}{n^{\gamma}}-\dfrac{1}{(n+1)^{\gamma}}\right)\sim\sum_{n=0}^{\infty}\dfrac{1}{n^{-\beta}}\cdot\dfrac{(n+1)^{\gamma}-n^{\gamma}}{n^{\gamma}}\sim\sum_{n=0}^{\infty}\dfrac{1}{n^{-\beta}}\cdot\dfrac{n^{\gamma-1}}{n^{2\gamma}}\sim\sum_{n=0}^{\infty}\dfrac{1}{n^{-\beta+\gamma+1}},$
where we used the symbol $\sim$ to denote that the sums are comparable. It is clear that the last sum converges whenever $-\beta+\gamma+1>1$, or equivalently, when $\gamma>\beta$. Since $\gamma=\frac{1}{2\alpha}$, the hypothesis $\beta\in(0,\frac{1}{2\alpha})$ guarantees the convergence of the sum, and hence of the integral, as desired. $\square$
4.2 Non-negative Lyapunov exponents. Proposition F
Next, we give conditions under which $\chi_{\omega_{U}}$ is non-negative. Our result is inspired in [KU23, Lemma 9.1.2, Corol. 9.1.3], but we remark that we do not assume that $f$ extends holomorphically (in fact, not even continuously) to a neighbourhood of ^$\widehat{\partial}U$.
Proposition 4.5.
(Lyapunov exponents are non-negative) Let $f\in\mathbb{K}$, and let $U$ be an invariant Fatou component for $f$, such that $\omega_{U}$ is $f$-invariant. Assume
  1. (a)
    $U$ is asymptotically contained in a sector of angle $\alpha\in\left(0,1\right)$, with order of growth $\beta\in(0,\frac{1}{2\alpha})$;
  2. (b)
    $SV(f)$ are thin in $\partial U$.
Then,
$\chi_{\omega_{U}}=\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}\geq 0.$
Note that Proposition 4.5 implies Proposition F.
Proof.
By Theorem 3.3, if $\omega_{U}$ is $f$-invariant, then $U$ is either an attracting basin or a Siegel disk, and $\omega_{U}$ is precisely the harmonic measure with basepoint the fixed point $p\in U$. In particular, $f|_{\partial U}$ is ergodic with respect to $\omega_{U}$.
  1. 1.
    Asymptotic contraction of $f^{n}|_{\partial U}$, $\omega_{U}$-almost everywhere.
    By Proposition 4.4, the integral
    $\chi_{\omega_{U}}=\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}$
    is well-defined. Since $f|_{\partial U}$ is ergodic, Birkhoff Ergodic Theorem 2.10, for $\omega_{U}$-almost every $x\in\partial U$,
    $\lim\limits_{n}\frac{1}{n}\log\left|(f^{n})^{\prime}(x)\right|=\chi_{\omega_{U}}(f).$
    We want to see that $\chi_{\omega_{U}}(f)\geq 0$. We shall assume, on the contrary, that $\chi_{\omega_{U}}(f)<0$, and seek for a contradiction.
    Since $\chi_{\omega_{U}}(f)<0$ and, for $\omega_{U}$-almost every $x\in\partial U$,
    $\lim\limits_{n}\dfrac{\left|(f^{n})^{\prime}(x)\right|}{(e^{\chi_{\omega_{U}}})^{n}}=1,$
    it follows that there exists $M\in(e^{\frac{\chi_{\omega_{U}}}{4}},1)$ and $n_{0}\coloneqq n_{0}(x)\in\mathbb{N}$ such that, for all $n\geq n_{0}$,
    $\left|(f^{n})^{\prime}(x)\right|^{\frac{1}{4}}\leq M^{n}<1.$
    We fix $x\in\partial U$ satisfying the previous property, and we denote by $\left\{x_{n}\right\}_{n}$ its forward orbit.
  2. 2.
    Shrinking domains where $f^{n}|_{\partial U}$ is univalent, around $\omega_{U}$-almost everywhere.
    Let $M\in(0,1)$ be the constant fixed in the previous step.
    Lemma 4.5.1.
    For $\omega_{U}$-almost every $x\in\partial U$ and $\lambda\in(M,1)$, there exists $n_{1}\coloneqq n_{1}(x)\geq n_{0}(x)$ such that $f^{n}|_{D(x_{n},\lambda^{n})}$ is univalent, for all $n\geq n_{1}$.
    In particular, since $\lambda>M$, $f^{n}|_{D(x_{n},M^{n})}$ is univalent, for all $n\geq n_{1}$.
    Proof.
    As in Definition 3.13, let
    $S_{\alpha,r}=\bigcup\limits_{i=1}^{k}\left\{z\in{\mathbb{C}}\colon\left|z-x_{i}\right|<r,\left|\textrm{Arg }\xi_{i}-\textrm{Arg }(\xi_{i}-z)\right|<\alpha\right\}.$
    Since $\beta<\frac{1}{2\alpha}$, we can choose $\gamma\in(\beta,\frac{1}{2\alpha})$. Then, applying the estimates of Lemma 2.26, we have
    $\omega_{U}\left(S_{\alpha,n^{-\frac{1}{\gamma}}}\right)\leq C\cdot n^{-\frac{1}{\gamma\cdot 2\alpha}},$
    and therefore
    $\sum_{n\geq 1}\omega_{U}\left(S_{\alpha,n^{-\frac{1}{\gamma}}}\right)\leq\sum_{n\geq 1}C\cdot n^{-\frac{1}{\gamma\cdot 2\alpha}}<+\infty.$
    By the assumption on the growth, for all $z\notin S_{\alpha,n^{-\frac{1}{\gamma}}}$ and $n\in\mathbb{N}$ large enough, we have
    $\left|f^{\prime}(z)\right|\leq C\cdot e^{n^{\frac{\beta}{\gamma}}},$
    for some constant $C>0$. We claim that there exists $C^{\prime}>0$ such that, for $z\notin S_{\alpha,n^{-\frac{1}{\gamma}}}$ and $n\in\mathbb{N}$ large enough,
    $\left|f^{\prime}(z)\right|\leq C^{\prime}\cdot\lambda^{-n/4},$
    where $\lambda\in(M,1)$ is the constant given in the statement of the lemma. Indeed, since $\beta/\gamma<1$, and $\log\lambda<0$, we have
    $l=\lim\limits_{n}\frac{e^{n^{\beta/\gamma}}}{\lambda^{-n/4}}=\lim\limits_{n}e^{n^{\beta/\gamma}}\cdot{\lambda^{n/4}};$
    $\log l=\lim\limits_{n}\log{e^{n^{\beta/\gamma}}}+{\log\lambda^{n/4}}=\lim\limits_{n}{n^{\beta/\gamma}}+\frac{n}{4}\log\lambda=-\infty;$
    hence $l=0$, and this already proves the existence of the constant $C^{\prime}$.
    Now, using the previous computations and the assumption that singular values are thin in $\partial U$, the first Borel-Cantelli Lemma 2.6 yields that, for $\omega_{U}$-almost every $x\in\partial U$ and $n$ large enough (depending on $x$), it holds
    1. (1.1)
      $x_{n+1}\notin\bigcup\limits_{s\in SV(f)}D(s,\lambda^{(n+1)/2})$,
    2. (1.2)
      $x_{n}\notin S_{\alpha,n^{-\frac{1}{\gamma}}}$.
    By 2(1.1), all inverse branches of $f$ are well-defined in $D(x_{n+1},\lambda^{(n+1)/2})$ and are univalent. Denote by $F$ the inverse branch of $f$ defined in $D(x_{n+1},\lambda^{(n+1)/2})$ such that $F(x_{n+1})=x_{n}$. By Koebe’s distortion estimates 2.1, we have
    $F(D(x_{n+1},{\lambda^{(n+1)/2}}))\supset D(x_{n},R),$
    where
    $R=\frac{1}{4}\cdot\left|F^{\prime}(x_{n+1})\right|\cdot{\lambda^{(n+1)/2}}=\frac{\lambda^{(n+1)/2}}{4}\cdot\frac{1}{\left|f^{\prime}(x_{n})\right|}\geq\frac{\lambda^{(n+1)/2}}{4}\cdot\lambda^{n/4}=K\cdot\lambda^{\frac{3n}{4}},$
    for some constant $K>0$. Since
    $\lim\limits_{n\to+\infty}\frac{\lambda^{n}}{\lambda^{\frac{3n}{4}}}=\lim\limits_{n\to+\infty}\lambda^{\frac{n}{4}}=0,$
    it follows that there exists $n_{1}\coloneqq n_{1}(x)$ large enough so that, for $n\geq n_{1}$,
    $F(D(x_{n+1},{\lambda^{(n+1)/2}}))\supset D(x_{n},\lambda^{n}).$
    Hence, $f^{n}|_{D(x_{n},\lambda^{n})}$ is univalent, for all $n\geq n_{1}$. $\square$
    Hence, we fix a point $x\in\partial U$ such that its forward orbit $\left\{x_{n}\right\}_{n}$ satisfies the following conditions, with $M\in(0,1)$ and $n_{1}\coloneqq n_{1}(x)$ as above:
    1. (2.1)
      $\left|(f^{n})^{\prime}(x)\right|^{\frac{1}{4}}\leq M^{n}<1$, for all $n\geq n_{1}$;
    2. (2.2)
      $f^{n}|_{D(x_{n},M^{n})}$ is univalent, for all $n\geq n_{1}$;
    3. (2.3)
      $x_{n}\notin SV(f)$, for all $n\in\mathbb{N}$.
    Note that the third condition follows from the fact that singular values have zero harmonic measure (Rmk. 3.11).
  3. 3.
    Quantitative contraction of $f^{n}|_{D(x,\lambda^{n})}$, for $n$ large enough.
    Let
[…]
    Observe that, since $\sum b_{n}\leq\sum M^{n}<\infty$, the infinite product in $P$ is convergent. Moreover, we can choose $r\coloneqq r(x)>0$ small enough so that, if
    $D_{n_{1}}\coloneqq D(x,r\cdot\prod_{m\geq 1}^{n_{1}}(1-b_{m})),$
    then $12r<$ 1, $f^{n_{1}}|_{D_{n_{1}}}$ is univalent, and $f^{n_{1}}({D_{n_{1}}})\subset D(x_{n_{1}},M^{n_{1}})$.
    Claim 4.5.2.
    For $n\geq n_{1}$, let
    $D_{n}\coloneqq D(x,r\cdot\prod_{m\geq 1}^{n}(1-b_{m})).$
    Then $f^{n}|_{D_{n}}$ is univalent, and $f^{n}(D_{n})\subset D(x_{n},M^{n})$.
    Note that, by definition, $D_{n}\supset D_{n+1}\supset\dots\supset D(x,rP)\ni x$. It follows from the claim that, for all $n\in\mathbb{N}$, $f^{n}$ is univalent in $D(x,rP)$ and $f^{n}|_{D(x,rP)}\subset D(x_{n},M^{n})$.
    Proof.
    We prove the claim inductively: assume the claim is true for $n\geq n_{1}$, and let us see that it also holds for $n+1$.
    First note that, since $f^{n}(D_{n})\subset D(x_{n},M^{n})$ (by inductive assumption) and $f$ is univalent in $D(x_{n},M^{n})$ (by Lemma 4.5.1), it follows that $f^{n+1}|_{D_{n}}$ is univalent. In particular, since $D_{n+1}\subset D_{n}$, we have that $f^{n}|_{D_{n+1}}$ is univalent.
    Now we use Koebe’s distortion estimates (Thm. 2.1) to prove the bound on the size of $f^{n+1}(D_{n+1})$. Indeed, since $f^{n+1}|_{D_{n}}$ is univalent, we have $f^{n+1}(D_{n+1})\subset D(x_{n+1},R)$, where
    $R=r\cdot\prod_{m\geq 1}^{n}(1-b_{m}))\cdot\left|(f^{n+1})^{\prime}(x)\right|\cdot\dfrac{2}{b^{3}_{n}}\leq 2r\cdot\dfrac{\left|(f^{n+1})^{\prime}(x)\right|}{\left|(f^{n+1})^{\prime}(x)\right|^{\frac{3}{4}}}\leq\left|(f^{n+1})^{\prime}(x)\right|^{\frac{1}{4}}\leq M^{n+1},$
    as desired. $\square$
  4. 4.
    Contradiction with the blow-up property of the Julia set.
    Let $R>0$ be small enough so that $D(p,R)\subset U$, where $p$ is the fixed point of $f$ in $U$. Let $n_{2}\geq n_{1}$ be such that $M^{n_{2}}<\frac{R}{2}$ (recall that $M\in(0,1)$, so such $n_{2}$ exists).
    Then, $f^{n_{2}}(D(x,rP))$ is a neighbourhood of $x_{n_{2}}=f^{n_{2}}(x)\in\mathcal{J}(f)$. By the previous step,
    $\bigcup_{n\geq n_{2}}f^{n}(D(x,rP))\subset\bigcup_{n\geq n_{2}}D(x_{n},M^{n})\subset\bigcup_{n\geq n_{2}}D(x_{n},M^{n_{2}})\subset\widehat{\mathbb{C}}\smallsetminus D(p,R/2).$
    This is a contradiction of the blow-up property of the Julia set. Notice that the contradiction comes from assuming $\chi_{\omega_{U}}<0$. Therefore, $\chi_{\omega_{U}}\geq 0$, and this ends the proof of the Proposition.
$\square$
4.3 The Lyapunov exponent for parabolic basins and Baker domains
As discussed before, the boundary of parabolic basins and doubly parabolic Baker domains do not support invariant probabilities which are absolutely continuous with respect to the harmonic measure $\omega_{U}$. However, the measure
[…]
is invariant under the radial extension of the associated inner function $g$ (taken such that 1 is the Denjoy-Wolff point) and its push-forward
$\mu=(\varphi^{*})_{*}\lambda_{\mathbb{R}}$
is an infinite invariant measure supported in ^$\widehat{\partial}U$.
Hence, in the case of parabolic basins and doubly parabolic Baker domains, we shall consider the Lyapunov exponent of $f$ with respect to $\mu$:
$\chi_{\mu}(f)\coloneqq\int_{\partial U}\log\left|f^{\prime}\right|d\mu.$
Although a priori, the Birkhoff Ergodic Theorem cannot be applied to relate the previous quantity with the assymptotic expansion (or contraction) along almost every orbit (since the measure $\mu$ is infinite), a similar argument can be applied to the first return map to get a similar conclusion (see Lemma 6.2(1.3)).
Finally, we show that, for parabolic basins, if $\log\left|f^{\prime}\right|$ is integrable with respect to harmonic measure, then it is also integrable with respect to the invariant measure $\mu$. Note also that Lemma 4.2 and Proposition 4.4 give conditions for $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$ and do not assume that $\omega_{U}$ is invariant, so they are still true in this new setting.
Proposition 4.6.
(Parabolic Lyapunov exponents) Let $f\in\mathbb{K}$, and let $U$ be a parabolic basin with $f$-invariant measure $\mu$. If $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$, then $\log\left|f^{\prime}\right|\in L^{1}(\mu)$. If, in addition, $\chi_{\omega_{U}}>0$, then $\chi_{\mu}>0$.
Proof.
For the first statement note that, since $\omega_{U}$ and $\mu$ are comparable except in a neighbourhood of the parabolic fixed point $p\in\partial U$, it is enough to check that
$\int_{D(p,r)}\log\left|f^{\prime}\right|d\mu<\infty,$
for some $r>0$. We note that, in contrast with the situation considered in Proposition 4.4, $\log\left|f^{\prime}\right|$ achieves a (finite) maximum and minimum around $p$ (since $f^{\prime}(p)=1$), but now the difficulty comes from the fact that $\mu$ is an infinite measure.
Throughout this proof, we shall use an equivalent definition for $\mu$ which is more convenient to do computations. Indeed, one can think $\mu$ as the push-forward of the measure
[…]
on $\mathbb{R}$ under the radial extension of a Riemann map $\psi\colon\mathbb{H}\to U$, where $\mathbb{H}$ denotes the upper half-plane and $\psi^{*}(0)=p$. Hence, the measure of an interval $(a,b)\subset(0,+\infty)$ is $\frac{b-a}{ab}$.
Next we estimate the function and the harmonic measure of disks in order to bound the previous integral. On the one hand, around the parabolic fixed point (which we assume to be the origin), we have the following normal form,
$f(z)=z+az^{q}+\dots,$
with $a\in\mathbb{C}$ and $q\geq 2$ (see e.g. [Mil06, Sect. 10]). Therefore,
$\log\left|f^{\prime}(z)\right|\sim\log(1+qa\left|z\right|)^{q-1}.$
On the other hand, by Lemma 2.25, we have that
$\lambda((\varphi^{*})^{-1}(D(p,r)))=\omega_{U}(D(p,r))\leq C\cdot\sqrt{r}.$
Therefore, the measure $\mu$ is maximized when $(\varphi^{*})^{-1}(D(p,r))$ is a circular interval of length $C\cdot\sqrt{r}$ centered at 1. Equivalently, with respect to the upper half-plane, this corresponds to an interval of length $C\cdot\sqrt{r}$ centered at 0.
Therefore, estimating the integral as in Proposition 4.4, we have that the bound
$\log\left|1+qa\frac{1}{n^{q-1}}\right|$
is the best possible for a set of $\mu$-measure greater than $\sqrt{n+1}-\sqrt{n}$. Then, we have that
$\int_{D(p,r)}\log\left|f^{\prime}\right|d\mu\sim\sum_{n=n_{0}}^{\infty}\log\left|1+qa\frac{1}{n^{q-1}}\right|\cdot(\sqrt{n+1}-\sqrt{n})\sim\sum_{n=n_{0}}^{\infty}\frac{1}{n^{q-1}}\cdot\frac{1}{2n}<\infty,$
as desired.
For the second statement, applying the Leau-Fatou Flower Theorem (see e.g. [Mil06, Sect. 10]), we have that $\log\left|f^{\prime}\right|>0$ in $D(p,r)\cap\partial U$, for $r$ small enough. Since $\omega_{U}>0$ we can assume, without loss of generality, that $r$ satisfies
$\int_{\partial U\smallsetminus D(p,r)}\log\left|f^{\prime}\right|d\omega_{U}>0.$
Then, the statement follows directly. $\square$
Remark 4.7.
(Lyapunov exponent for Baker domains) Note that Proposition 4.6 is stated only for parabolic basins, and its proof used the normal form around a parabolic fixed point. For a Baker domain, there is no longer a normal form around the convergence point, since it is an essential singularity for $f$, and hence the argument cannot be applied in general. However, for some explicit Baker domains, similar estimates can be obtained and the argument may work ad hoc.
Indeed, consider for instance the Baker domain of the map $f(z)=z+e^{-z}$ (see [FJ23]). Since it is contained in a strip and $f$ has finite order, by Proposition 4.4, $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$. Hence, as before, to see that $\log\left|f^{\prime}\right|\in L^{1}(\mu)$, it is enough to check integrability in a neighbourhood of infinity. Note that $f^{\prime}(z)=1-e^{-z}$, so the estimates on $\left|f^{\prime}\right|$ are even better than in the parabolic case, and the same argument can be applied. Moreover, in this particular case, one can check that $\left|f^{\prime}\right|>1$ in a neighbourhood of $\partial U$ [FJ23, Prop. 3.6], so $\chi_{\mu}(f)>0$.
5 Pesin theory for attracting basins. Theorem A
In this section, we take on the main challenge of this paper: developing Pesin theory for a simply connected attracting basin $U$ of a function of class $\mathbb{K}$, or, in other words, proving that generic infinite inverse branches are well-defined in $\partial U$.
The easiest assumption one shall make to get that generic infinite inverse branches are well-defined in $\partial U$, is that there exists $x\in\partial U$ and $r>0$ so that $D(x,r)\cap P(f)=\emptyset$. Indeed, in such case, all iterated inverse branches are well-defined in $D(x,r)$. Moreover, since $f|_{\partial U}$ is ergodic and recurrent, and $D(x,r)$ has positive harmonic measure, it follows that the forward orbit of $\omega_{U}$-almost every $y\in\partial U$ eventually falls in $D(x,r)$, so all iterated inverse branches are well-defined around $y$.
Note that this idea works not only for attracting basins, but also for parabolic basins and recurrent Baker domains, since the boundary map has the same ergodic properties (Thm. 3.3). This is the way explored in [Jov24, Thm. 6.1].
The previous method has a main limitation: it does not work when $\partial U\subset P(f)$. Even in the case where $f$ is a polynomial, one can find examples for which $\partial U\subset P(f)$, or even $\mathcal{J}(f)\subset P(f)$. Our goal is precisely to show that, even in the case where $\partial U\subset P(f)$, if singular values are thin, generic inverse branches are well-defined.
Clearly, $D(x,r)\cap P(f)=\emptyset$ is equivalent to assuming that all inverse branches are well-defined in $D(x,r)$. Hence, in order to develop a general theory, one should work with each infinite backward orbit separatedly, and try to find a disk where the inverse branches corresponding to this backward orbit are well-defined, but other inverse branches may fail to be defined. Here is where Rohklin’s natural extension (Sect. 2.5) plays a crucial role.
Therefore, let $U$ be a simply connected attracting basin for a map $f\in\mathbb{K}$ with fixed point $p\in U$, and consider the measure-theoretical dynamical system given by $(\partial U,\omega_{U},f)$, where $\omega_{U}$ is the harmonic measure with basepoint $p$. Note that, through this section, $\omega_{U}$ stands for the harmonic measure with basepoint $p$, although we do not write it explicitly. Recall that $\omega_{U}$ is $f$-invariant, ergodic and recurrent (Thm. 3.3). Note also that we omit the dependence of the previous dynamical system on the $\sigma$-algebra $\mathcal{B}(\widehat{\mathbb{C}})$, in order to lighten the notation.
Now, consider the natural extension of $(\partial U,\omega_{U},f)$, denoted by $(\widetilde{\partial U},\widetilde{\omega_{U}},\widetilde{f})$, and given by the projecting morphisms $\left\{\pi_{U,n}\right\}_{n}$. We note that $(\partial U,\omega_{U},f)$ is indeed a Lebesgue probability space (in fact, it is isomorphic, in the measure-theoretical sense, to the unit interval), and hence Theorem 2.12 can be applied to guarantee the existence of Rokhlin’s natural extension. Thus, $\widetilde{\partial U}$ is the space of backward orbits $\left\{x_{n}\right\}_{n}\subset\partial U$, with $f(x_{n+1})=x_{n}$ for $n\geq 0$, and $\widetilde{f}\colon\widetilde{\partial U}\to\widetilde{\partial U}$ is the automorphism which makes the following diagram commute.
[…]
Since the natural extension inherits the ergodic properties of the original dynamical system, we have that $\widetilde{\omega_{U}}$ is an $\widetilde{f}$-invariant, ergodic and recurrent probability (Prop. 2.13). Moreover, for every measurable set $A\subset\partial U$ with $\mu(A)>0$ and $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{\partial U}$, there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A$ (Corol. 2.14).
We shall rephrase Theorem A in terms of Rokhlin’s natural extension as follows.
Theorem 5.1.
(Inverse branches are well-defined almost everywhere) Let $f\in\mathbb{K}$, and let $U$ be a simply connected attracting basin for $f$, with fixed point $p\in U$. Let $\omega_{U}$ be the harmonic measure in $\partial U$ with base point $p$. Assume:
  1. (a)
    $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$, and $\chi_{\omega_{U}}(f)>0$;
  2. (b)
    singular values are thin in ^$\widehat{\partial}U$.
Then, for $\widetilde{\omega_{U}}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\in\widetilde{\partial U}$, there exists $r\coloneqq r(\left\{x_{n}\right\}_{n})>0$ such that
  1. (i)
    the inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},r)$;
  2. (ii)
    for every $\chi\in(-\chi_{\omega_{U}},0)$, there exists a constant $C>0$ such that, for all $n\in\mathbb{N}$,
    $\left|F^{\prime}_{n}(x_{0})\right|<C\cdot e^{\chi\cdot n};$
  3. (iii)
    for every $r_{0}\in(0,r)$, there exists $m\in\mathbb{N}$ such that
    $F_{m}(D(x_{0},r))\subset D(x_{0},r_{0}).$
We show now how to deduce Theorem A from Theorem 5.1, and later we give the proof of it.
Proof of Theorem A.
Observe that the assumptions of Theorems A and 5.1 are the same. We have to see that the conclusions of Theorem A can be derived from the ones of Theorem 5.1. But this follows straightforward from Corollary 2.14. Indeed, since $\widetilde{f}$ is ergodic and recurrent with respect to $\widetilde{\omega_{U}}$, for any $A\subset X$ measurable set with $\omega_{U}(A)>0$, for $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{X}$, there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A$. It follows that, for every countable collection of measurable sets $\left\{A_{k}\right\}_{k}\subset\partial U$ with $\omega_{U}(A_{k})>0$, then for $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{X}$, there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A_{k}$. Using the measure-preserving morphism $\pi_{U,0}$ to project into the phase space, we get the desired conclusion. $\square$
Remark 5.2.
Before starting the proof let us note that we are assuming $f\in\mathbb{K}$ just because it is the largest class of functions in which Fatou components are defined. We do not use the fact that functions in class $\mathbb{K}$ have only countably many singularities, we only use that singular values are thin in $\partial U$.
The remaining of the section is devoted to prove Theorem 5.1.
5.1 Proof of Theorem 5.1
Recall that $\omega_{U}$ is a $f$-invariant ergodic probability in $\partial U$. We fix $M\in(e^{\frac{1}{4}\cdot\chi},1)$. Moreover, we are assuming that singular values are thin in ^$\widehat{\partial}U$ (Def. 3.8), write $\left\{v_{1},\dots,v_{k}\right\}$ for the punctures,
$\mathcal{S}\coloneqq SV(f)\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon),$
and consider the $\rho$-metric in ^$\widehat{\mathbb{C}}\smallsetminus\left\{v_{1},\dots,v_{k}\right\}$. Let $\eta>0$, $\mu\in(0,1)$, $d\in\mathbb{N}$ as in Lemma 3.10, i.e. such that there are no singular values in
$\bigcup_{i=1}^{k}D(v_{i},\varepsilon)\cap\left\{z\in\widehat{\mathbb{C}}\colon\textrm{dist}_{\rho}(z,\widehat{\partial}U)<\eta\right\},$
and there are at most $n^{d}$ singular values in
$\left\{z\in\widehat{\mathbb{C}}\colon\textrm{dist}(z,\widehat{\partial}U)<\mu^{n}\right\}\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$
which are $\mu^{n}$-separated.
Lemma 5.2.1.
(Almost every backward orbit does not come close to singular values) For $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{\partial U}$, it holds
  1. (1.1)
    $x_{0}\notin\bigcup\limits_{s\in SV(f)}\bigcup\limits_{n\geq 0}f^{n}(s)$,
  2. (1.2)
    $\lim\limits_{n}\dfrac{1}{n}\log\left|(f^{n})^{\prime}(x_{n})\right|=\chi_{\omega_{U}}(f)$,
  3. (1.3)
    if $D_{n}\coloneqq\bigcup\limits_{s\in\mathcal{S}}D(s,\eta\cdot M^{n})$, then $x_{n}\in D_{n}$ only for a finite number of $n$’s,
  4. (1.4)
    for $i=1,\dots,k$, $x_{n}\in D(v_{i},\frac{4}{\varepsilon}\cdot M^{n/2})$ only for a finite number of $n$’s.
Proof.
Since the finite intersection of sets of full measure has full measure, it is enough to show that each of the four conditions is satisfied in a set of full measure.
For condition (1.1), recall that $\omega_{U}(SV(f))=0$ (Rmk. 3.11), so
$\omega_{U}(\bigcup\limits_{s\in SV(f)}\bigcup\limits_{n\geq 0}f^{n}(s))=0,$
since $f$ is holomorphic, and hence absolutely continuous.
Hence, for $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}$, $x_{0}\notin\bigcup\limits_{s\in SV(f)}\bigcup\limits_{n\geq 0}f^{n}(s)$. Note that this is equivalent to say that, for $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}$, $x_{n}$ is not a singular value for $f$. In particular, the punctures $\left\{v_{1},\dots,v_{k}\right\}$ in which the $\rho$-metric is not defined are singular values for $f$. Hence, for the backward orbits $\left\{x_{n}\right\}_{n}$ we are considering the $\rho$-metric is well-defined around every point in the orbit.
Requirement (1.2) follows from Birkhoff Ergodic Theorem 2.10 applied to the map $\log\left|f^{\prime}\right|$, which is integrable by the assumption (a). Indeed, for $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{\partial U}$, it holds
$\chi_{\omega_{U}}(f)$ $=\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}=\lim\limits_{m}\frac{1}{m}\sum_{k=0}^{m-1}\log\left|f^{\prime}(f^{k}(x_{0}))\right|=$
$=\lim\limits_{m}\frac{1}{m}\sum_{k=0}^{n-1}\log\left|f^{\prime}(f^{k}(\pi_{U,0}(\left\{x_{n}\right\}_{n})))\right|=\lim\limits_{m}\frac{1}{m}\sum_{k=0}^{m-1}\log\left|f^{\prime}(\pi_{U,0}(\widetilde{f}^{k}(\left\{x_{n}\right\}_{n})))\right|,$
where in the last two equalities we used the properties of Rokhlin’s natural extension.
Now, $\widetilde{f}$ is a measure-preserving automorphism, and, since $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$, $\log\left|f^{\prime}\circ\pi_{U,0}\right|\in L^{1}(\widetilde{\omega_{U}})$. Then, Birkhoff Ergodic Theorem yields
$\lim\limits_{m}\frac{1}{m}\sum_{k=0}^{m-1}\log\left|f^{\prime}(\pi_{U,0}(\widetilde{f}^{k}(\left\{x_{n}\right\}_{n})))\right|=\lim\limits_{m}\frac{1}{m}\sum_{k=0}^{m-1}\log\left|f^{\prime}(\pi_{U,0}(\widetilde{f}^{-k}(\left\{x_{n}\right\}_{n})))\right|=$
$=$ $\lim\limits_{m}\frac{1}{m}\sum_{k=0}^{m-1}\log\left|f^{\prime}(x_{k})\right|=\lim\limits_{m}\frac{1}{m}\log(\left|f^{\prime}(x_{0})\right|\dots\left|f^{\prime}(x_{m})\right|)=\lim\limits_{m}\frac{1}{m}\log\left|(f^{m})^{\prime}(x_{m})\right|,$
where in the last equality we used the chain rule for the derivative with respect to the $\rho$-metric.
Putting everyting together, we get that for $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}$, it holds
$\lim\limits_{n}\dfrac{1}{n}\log\left|(f^{n})^{\prime}(x_{n})\right|=\chi_{\omega_{U}}(f)$
as desired.
Condition (1.3) follows from the first Borel-Cantelli Lemma 2.6. Since we are assuming that singular values are thin in ^$\widehat{\partial}U$ (hypothesis (b)), it follows from Lemma 3.10 that
$\sum_{n\geq 1}\widetilde{\omega_{U}}(\pi_{U,n}^{-1}(D_{n}))=\sum_{n\geq 1}\omega_{U}(D_{n})<\infty,$
Thus, by the Borel-Cantelli Lemma, for $\widetilde{\omega_{U}}$-almost every $\left\{x_{n}\right\}_{n}\in\widetilde{\partial U}$, $x_{n}\in D_{n}$ for only finitely many $n$’s, as desired.
Requirement (1.4) is again an application of the first Borel-Cantelli Lemma, using that, for all $i=1,\dots,n$,
$\sum_{n\geq 1}\widetilde{\omega_{U}}(\pi_{U,n}^{-1}(D(v_{i},\frac{4}{\varepsilon}\cdot M^{n/2})))=\sum_{n\geq 1}\omega_{U}(D(v_{i},\frac{4}{\varepsilon}\cdot M^{n/2}))<\infty,$
according to Lemma 2.23. This ends the proof of the Lemma. $\square$
Let us fix a backward orbit $\left\{x_{n}\right\}_{n}$ satifying the conditions of the previous lemma. By (1.3) and (1.4), there exists $n_{1}\in\mathbb{N}$ such that, for $n\geq n_{1}$,
[…]
Moreover, by (1.2), there exists $n_{2}\in\mathbb{N}$, $n_{2}\geq n_{1}$ such that, for $n\geq n_{2}$,
$\left|(f^{n})^{\prime}(x_{n})\right|^{-\frac{1}{4}}<M^{n}<1.$
Without loss of generality, we also assume $n_{2}$ is large enough so that, for $n\geq n_{2}$,
$M^{n/2}<2\cdot\dfrac{\eta}{\varepsilon}.$
Note that both $n_{1}$ and $n_{2}$ do not only depend on the starting point $x_{0}$, but on all the backward orbit $\left\{x_{n}\right\}_{n}$. Two different backward orbits starting at $x_{0}$ may require different $n_{1}$ or $n_{2}$.
We set the following notation.
$S\coloneqq\sum_{n\geq 1}\left|(f^{n})^{\prime}(x_{n})\right|^{-\frac{1}{4}}$
$b_{n}\coloneqq\frac{1}{2}\left|(f^{n+1})^{\prime}(x_{n+1})\right|^{-\frac{1}{4}}$
$P=\prod_{n\geq 1}(1-b_{n})^{-1}$
Note that the series in the definition of $S$ converges, since the $n$-th term is bounded by $M^{n}$, and $M\in(0,1)$. The convergence of that series gives also the convergence of the infinite product in $P$. Note that it is satisfied, for all $n\geq 1$,
$\prod_{m\geq n}(1-b_{m})^{-1}\geq\prod_{m\geq n+1}(1-b_{m})^{-1}.$
Choose $r\coloneqq r(\left\{x_{n}\right\}_{n})>0$ such that
  1. (2.1)
    $32rP<\eta$,
  2. (2.2)
    for $1\leq n\leq n_{2}$, the branch $F_{n,x_{0},x_{n}}$ of $f^{-n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},rP)$,
  3. (2.3)
    $\textrm{diam}\left(F_{n_{2}}(D(x_{0},r\prod\limits_{m\geq n_{2}}(1-b_{m})^{-1}))\right)\leq\eta\cdot M^{n_{2}}$,
  4. (2.4)
    $\textrm{diam}_{\rho}\left(F_{n_{2}}(D(x_{0},r\prod\limits_{m\geq n_{2}}(1-b_{m})^{-1}))\right)\leq\eta$.
For the sake of simplicity, we shall call the inverse branch considered in (2.2) just $F_{n}$. For any other inverse branch, we keep the notation $F_{n,z,w}$ (meaning that $F_{n}$ is an inverse branch of $f^{n}$ sending $z$ to $w$, as stated in the introduction).
We note that such a $r>0$ exists. Indeed, it is enough to check that the inverse branch $F_{n_{2}}$ sending $x_{0}$ to $x_{n_{2}}$ is well-defined in an open neighbourhood of $x_{0}$ (and then $r$ is chosen accordingly). Since the set of singular values is closed, for $n=0,\dots,n_{2}-1$, there exists an open neighbourhood around $x_{n}$ which not contain singular values. Hence, all inverse branches $F_{1,x_{n+1},x_{n}}$ of $f^{-1}$ are well-defined in such neighbourhoods. Since the finite intersection of open sets is open, it follows that $F_{n_{2}}$ is well-defined in an open neighbourhood of $x_{0}$. Then, choosing $r>0$ satisfying condition (2.1) is immediate, and this already implies the claim.
The following inverse branches are constructed by induction as follows.
Claim 5.2.2.
For every $n\geq n_{2}$, there exists an inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$, defined in $D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1})$, and such that
$\textrm{\em diam}\left(F_{n}(D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1}))\right)\leq\eta\cdot M^{n},$
$\textrm{\em diam}_{\rho}\left(F_{n}(D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1}))\right)\leq\eta.$
Note that proving the claim ends the proof of the theorem. Indeed, letting $n\to\infty$ we get that all inverse branches are well-defined in $D(x_{0},r)$, with $r>0$. The estimate on the derivative follows from the fact that, for $n\geq n_{2}$, $\left|(f^{n})^{\prime}(x_{n})\right|^{-\frac{1}{4}}<M^{n}<1$, with $M\in(e^{\frac{1}{4}\chi},1)$. Hence, there exists $C>0$ such that, for all $n\in\mathbb{N}$,
$\left|F_{n}^{\prime}(x_{0})\right|=\left|(f^{n})^{\prime}(x_{n})\right|^{-1}<e^{\chi\cdot n}.$
Proof of the claim.
Suppose the claim is true for $n\geq n_{2}$, and let us see that it also holds for $n+1$. First, using the inductive hypothesis and by the choice of $\eta$ and $n_{2}$, we have that there exists a neighbourhood $D$ of $x_{n}\in\partial U$ satisfying
$D_{\rho}(x_{n},\eta)\supset D,\,\,D(x_{n},\eta\cdot M^{n})\supset D,\,\textrm{and }D\cap SV(f)=\emptyset.$
(Note that singular values in $D_{\rho}(x_{n},\eta)$ are in $\mathcal{S}$, and these lie at a bounded Euclidean distance from $x_{n}$ by (1.3).)
Hence, there exists a branch $F_{1,x_{n},x_{n+1}}$ of $f^{-1}$ satisfying $F_{1}(x_{n})=x_{n+1}$, well-defined in $D$. By the inductive hypothesis, there exists an inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$, defined in $D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1})$, and such that
$\textrm{diam}\left(F_{n}(D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1}))\right)\leq\eta\cdot M^{n},$
$\textrm{diam}_{\rho}\left(F_{n}(D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1}))\right)\leq\eta.$
Set $F_{n+1}=F_{1}\circ F_{n}$. Then, $F_{n+1}$ is well-defined in $D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1})$, and sends $x_{0}$ to $x_{n+1}$.
Finally, we shall prove the bounds on the diameter. For the Euclidean bound, we shall use Koebe’s distortion estimate (Thm. 2.1). Indeed, note that $F_{n+1}$ is well-defined in $D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1})$ (which is strictly larger than $D(x_{0},r\prod_{m\geq{n+1}}(1-b_{m})^{-1})$, and the ratio between the two radii is $(1-b_{n})$). Then, we apply Koebe’s distortion estimate to the larger disk $D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1})$ in which $F_{n+1}$ is conformal, and with smaller disk $D(x_{0},r\prod_{m\geq n+1}(1-b_{m})^{-1})$. We get
[…]
$\leq 32\left|(f^{n+1})^{\prime}(x_{n+1})\right|^{-\frac{1}{4}}rP,$
where we used that $b_{n}=\frac{1}{2}\left|(f^{n+1})^{\prime}(x_{n+1})\right|^{-\frac{1}{4}}$, $\left|F_{n+1}^{\prime}(x_{0})\right|=\left|(f^{n+1})^{\prime}(x_{n+1})\right|^{-1}$ and $\prod_{m\geq n+1}(1-b_{m})^{-1}\leq P$. Finally, applying (2.1) and the estimate $\left|(f^{n})^{\prime}(x_{n})\right|^{-\frac{1}{4}}<M^{n}$, we get that
[…]
as desired.
We shall prove now that
[…]
The previous estimate on the Euclidean diameter, together with Lemma 3.12 (note that $v_{i}\notin D(x_{n+1},2r)$, where $r=\eta\cdot M^{n}$, by (1.4) together with the assumption $M^{n/2}<2\cdot\frac{\eta}{\varepsilon}$), yields
$F_{n+1}(D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1}))\subset D(x_{n+1},\eta\cdot\frac{M^{n}}{2})\subset D_{\rho}(F_{n+1}(x_{0}),8\cdot\eta\cdot M^{n}\cdot\rho(x_{n+1})).$
Since the backward orbit $\left\{x_{n}\right\}_{n}$ satisfies that, for $n\geq n_{1}$, $x_{n}\notin\bigcup\limits_{i=1}^{k}D(v_{i},\frac{4}{\varepsilon}\cdot M^{n/2})$, it follows that
$\rho(x_{n+1})=\frac{\varepsilon^{2}}{\left|x_{n+1}-v_{i}\right|^{2}}\leq\frac{\varepsilon^{2}}{\frac{16}{\varepsilon^{2}}\cdot M^{n}}=\frac{1}{16\cdot M^{n}}.$
Putting everything together, we have
$F_{n+1}(D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1}))\subset D_{\rho}(x_{n+1},\frac{\eta}{2}),$
as desired. This ends the proof of the claim. $\square$
As noted before, this last claim ends the proof of Theorem 5.1. $\square$
6 Entire functions and the first return map. Theorem B
In this section, we extend Theorem A to parabolic and recurrent Baker domains of entire maps. Note that the main challenge is that there does not exist an invariant probability which is absolutely continuous with respect to harmonic measure. However, the existence of an invariant $\sigma$-finite measure in $\partial U$ absolutely continuous with respect to $\omega_{U}$ will allow us to perform Pesin theory, by means of the first return map.
We shall start by constructing Rokhlin’s natural extension (note that this is indeed possible due to the existence of the $\sigma$-invariant measure). We do this by showing that Rokhlin’s natural extension is compatible with the use of first return maps if the transformation we consider is recurrent. This allows us to move from our problem of finding inverse branches in a space endowed with a $\sigma$-invariant measure to a probability space, where we can perform Pesin theory in a stardard way. We do this construction of the first return map and Rokhlin’s natural extension in Section 6.1, and finally we develop Pesin theory in Section 6.2.
Remark 6.1.
In this section, we restrict ourselves to entire functions, i.e. functions for which $\infty$ is (possibly) the only essential singularity of $f$, and it has no finite preimages (there are no poles). Hence, the infinite sequence $\left\{f^{n}(z)\right\}_{n}$ is well-defined for all $z\in\mathbb{C}$. This is indeed the crucial fact that we need for our proof to work. In fact, if $f$ were rational, the statement would still true: what we really need is being able to iterate infinitely many times $f$ without falling in an essential singularity (compare with Lemma 6.5.1).
6.1 The first return map and Rokhlin’s natural extension
We are working under the assumption that $U$ is either an attracting basin, a parabolic basin or a recurrent Baker domain. Note that, since we are assuming $f$ is entire, $U$ is always simply connected [Bak84].
We shall distinguish two cases. First, if $U$ is an attracting basin with attracting fixed point $p\in U$, then $\omega_{U}(p,\cdot)$ is an $f$-invariant probability in $\partial U$. The second possible scenario is when $U$ is a parabolic basin or a recurrent Baker domain. In this case, there does not exists an $f$-invariant measure which is absolutely continuous with respect to harmonic measure, but there exists a $\sigma$-finite measure which is invariant under $f$ and absolutely continuous with respect to $\omega_{U}$. Indeed, the measure
[…]
is invariant under the radial extension of the associated inner function $g$ (taken such that 1 is the Denjoy-Wolff point) and its push-forward
$\mu=(\varphi^{*})_{*}\lambda_{\mathbb{R}}$
is an infinite invariant measure supported in ^$\widehat{\partial}U$. Note that $\lambda_{\mathbb{R}}$ is a $\sigma$-finite measure.
In both cases, we shall denote this invariant probability by $\mu$, and, according to Section 2.5, one can consider the Rokhlin’s natural extension in both cases. Thus, $\widetilde{\partial U}$ is the space of backward orbits $\left\{x_{n}\right\}_{n}\subset\partial U$, with $f(x_{n+1})=x_{n}$ for $n\geq 0$, and $\widetilde{f}\colon\widetilde{\partial U}\to\widetilde{\partial U}$ is the automorphism which makes the following diagram commute.
[…]
On the other hand, one can get an equivalent construction of backward orbits by means of the first return map. Indeed, we can write $\partial U$, up to a set of zero measure, as
$\partial U=\bigcup\limits_{k\geq 0}X_{k},\hskip 14.22636pt\left\{X_{k}\right\}_{k}\subset\mathcal{B}(\partial U),\hskip 14.22636pt\mu(X_{k})\in(0,+\infty).$
For technical reasons, and without loss of generality we shall assume that the sets $\left\{X_{k}\right\}_{k}$ are pairwise disjoint, and each of them is at a positive distance of grand orbit finite points. Hence, up to a set of zero measure, $X$ can be writen as the union of pairwise disjoint measurable sets $\left\{X_{k}\right\}_{k}$, each of them at a positive distance of grand orbit finite points, and every point $x\in X_{k}$ coming back to $X_{k}$ infinitely often. This last requirement is possible due to recurrence.
Then, for all $k\geq 0$, consider the first return map to $X_{k}$, i.e.
$f_{X_{k}}\colon$ $X_{k}\longrightarrow X_{k}$
$x\longmapsto f^{T(x)}(x),$
where $T(x)$ denotes the first return time of $x$ to $X_{k}$. We consider the measure-theoretical dynamical system $(X_{k},\mu_{k},f_{X_{k}})$, where
$\mu_{k}(A)\coloneqq\frac{\mu(A\cap X_{k})}{\mu(X_{k})},$
for every measurable set $A\subset\partial U$. Note that $(X_{k},\mu_{k})$ is a probability space.
The following properties of the first return map $f_{X_{k}}$ will be needed.
Lemma 6.2.
(First return map) Let $f_{X_{k}}\colon X_{k}\to X_{k}$ be defined as above. Then, the following holds.
  1. (1.1)
    $\mu_{k}$ is invariant under $f_{X_{k}}$. In particular, $f_{X_{k}}$ is recurrent with respect to $\mu_{k}$.
  2. (1.2)
    $f_{X_{k}}$ is ergodic with respect to $\mu_{k}$.
  3. (1.3)
    If $\log\left|f^{\prime}\right|\in L^{1}(\mu)$ , then $\log\left|f^{\prime}_{X_{k}}(x)\right|\coloneqq\log\left|(f^{T(x)})^{\prime}(x)\right|\in L^{1}(\mu_{k})$ and
    $\int_{X_{k}}\log\left|f^{\prime}_{X_{k}}\right|d\mu_{k}=\dfrac{1}{\mu(X_{k})}\int_{\partial U}\log\left|f^{\prime}\right|d\mu.$
Proof.
The three claims are standard facts of measure-theoretical first return maps. More precisely, (1.1) and (1.2) follow from [URM22, Prop. 10.2.1] and [URM22, Prop. 10.2.7], respectively. Statement (1.3) comes from [URM22, Prop. 10.2.5], applied to $\varphi=\log\left|f^{\prime}\right|$ and $\varphi_{X_{k}}=\log\left|f^{\prime}_{X_{k}}\right|$square
Since $(X_{k},\mu_{k})$ is a Lebesgue probability space, and $\mu_{k}$ is $f_{k}$-invariant, we shall consider its Rohklin’s natural extension, denoted by $(\widetilde{X_{k}},\widetilde{f_{k}})$, and given by the projecting morphisms $\left\{\pi_{X_{k,n}}\right\}_{n}$. Thus, $\widetilde{X_{k}}$ is the space of backward orbits $\left\{x^{k}_{n}\right\}_{n}\subset\partial U$, with $f_{X_{k}}(x^{k}_{n+1})=x^{k}_{n}$ for $n\geq 0$, and $\widetilde{f_{X_{k}}}\colon\widetilde{X_{k}}\to\widetilde{X_{k}}$ is the automorphism which makes the following diagram commute.
[…]
Since the natural extension of a probability space inherits the ergodic properties of the original dynamical system, we have that $\widetilde{\mu_{k}}$ is an $\widetilde{f_{X_{k}}}$-invariant, ergodic and recurrent probability (Prop. 2.13).
We claim that both constructions of spaces of backward orbits are essentially the same, with the only difference that, when considering the first return map, orbits starting at the set $X_{k}$ are written ‘packed’ according to their visits to $X_{k}$.
Indeed, given a backward orbit $\left\{x^{k}_{n}\right\}_{n}\subset X_{k}$ for $f_{X_{k}}$, we can associate to it unambiguously a backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ for $f$ as follows. Let $x_{0}\coloneqq x^{k}_{0}$, and let $x_{T(x^{k}_{1})}\coloneqq x^{k}_{1}$. Since $f_{X_{k}}(x^{k}_{1})=f^{T(x^{k}_{1})}(x^{k}_{1})=x^{k}_{0}$, for $n=1,\dots,T(x^{k}_{1})-1$, let $x_{n}\coloneqq f^{T(x^{k}_{1})-n}(x^{k}_{1})$. The rest of the backward orbit is defined recursively. We say that the $f$-backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ is associated to the $f_{X_{k}}$-backward orbit $\left\{x^{k}_{n}\right\}_{n}\subset X_{k}$. In the same way, if a $f$-backward orbit visits $X_{k}$ infinitely often, we can associate a $f_{X_{k}}$-backward orbit to it.
As noted above, for every $f_{X_{k}}$-backward orbit we can associate a $f$-backward orbit. Moreover, the converse is true $\widetilde{\mu}$-almost everywhere.
Lemma 6.3.
(Distribution of $f_{X_{k}}$-backward orbits in $\partial U$) Let $(\partial U,\mu,f)$ and $(X_{k},\mu_{k},f_{X_{k}})$, and consider their natural extensions as before. Then, the following holds.
  1. (a)
    If $A$ is a measurable set in $\partial U$ with $\mu(A)>0$, then for $\widetilde{\mu_{k}}$-almost every $\left\{x^{k}_{n}\right\}_{n}\in\widetilde{X_{k}}$, with associated backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ under $f$, there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A$.
  2. (b)
    For $\widetilde{\mu}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ with $x_{0}\in X_{k}$, $x_{n}\in X_{k}$ infinitely often. Hence, to $\widetilde{\mu}$-almost every $f$-backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ with $x_{0}\in X_{k}$, we can associate a $f_{X_{k}}$-backward orbit.
  3. (c)
    If $A$ is a measurable set in $\partial U$ with $\mu(A)>0$, then for $\widetilde{\mu}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ under $f$, there exists a sequence $n_{k}\to\infty$ such that $x_{n_{k}}\in A$.
Proof.
Note that if $U$ is an attracting basin, the Rokhlin’s natural extension is made out of an invariant probability, and hence the content of this lemma is automatically implied by Proposition 2.13. Therefore, the lemma is of special interest when $U$ is a parabolic basin or a recurrent Baker domain.
  1. (a)
    Assume, on the contrary, that there exists a measurable set $A\subset\partial U$ of positive measure, and a subset $K\subset\widetilde{X_{k}}$ of positive $\widetilde{\mu_{k}}$-measure, such that their associated $f$-backward orbits do not intersect $A$. Since $\widetilde{f_{X_{k}}}$ is ergodic and recurrent with respect to $\widetilde{\mu_{k}}$, it follows that, for $\widetilde{\mu_{k}}$-almost every backward orbit in $X_{k}$, its associated $f$-backward orbits does not intersect $A$ (since any $\widetilde{f_{X_{k}}}$-backward orbit visits $K$ infinitely often, Prop. 2.13).
    We claim that this implies that
    $K_{\infty}\coloneqq\bigcup_{n\in\mathbb{Z}}f^{n}(X_{k});\hskip 14.22636pt\mu(K_{\infty}\cap A)=0.$
    Indeed, on the contrary, assume $\mu(K_{\infty}\cap A)>0$. Since we can write
    $K_{\infty}\cap A=\bigcup_{n\in\mathbb{Z}}f^{n}(X_{k})\cap A=\bigcup_{n\in\mathbb{Z}}(f^{n}(X_{k})\cap A),$
    it follows that, for some $n\geq 0$, $\mu(f^{n}(X_{k})\cap A)>0$. Since $f|_{\partial U}$ is ergodic and recurrent, then $\mu$-almost every $x\in f^{n}(X_{k})\cap A$ visits $X_{k}$ under forward iteration. This already implies that, since $f|_{\partial U}$ is non-singular, there exists a subset $K^{\prime}$ of $X_{k}$ of positive measure such that $f_{X_{k}}(K^{\prime})$ intersects $A$ in some intermediate step. Since any $\widetilde{f_{X_{k}}}$-backward orbit visits $K^{\prime}$ infinitely often, we reach a contradiction.
    Finally, note that $K_{\infty}$ is an $f$-invariant set such that $\mu(K_{\infty})\geq\mu(X_{k})>0$ and $\mu(\partial U\smallsetminus K_{\infty})\geq\mu(A)>0$. This contradicts the fact that the system $(\partial U,\mu,f)$ is ergodic.
  2. (b)
    For the first statement, we proceed by contradiction. Assume there exists $K\subset\widetilde{\partial U}$ of positive $\widetilde{\mu}$-measure, such that for every $\left\{x_{n}\right\}_{n}\in K$, $x_{0}\in X_{k}$ and $x_{n}\notin X_{k}$, for all but finitely many $n\geq 1$. Then,
    $K=\bigcup_{m\geq 0}K_{m}=\bigcup_{m\geq 0}\left\{\left\{x_{n}\right\}\in K\colon x_{n}\notin X_{k},\textrm{ for all }n\geq m\right\}.$
    Note that $\widetilde{f}^{-1}(K)\subset K$, and, since $\widetilde{f}$ is measure-preserving, the set
    $K_{\infty}=\bigcup_{n\geq 0}f^{-n}(K),$
    has measure $\widetilde{\mu}(K)$ (in particular, $K_{\infty}$ has positive $\widetilde{\mu}$-measure and $\pi_{U,0}(K_{\infty})$ has positive $\mu$-measure). Note that $\pi_{U,0}(K_{\infty})$ consists of points whose orbit never enters $X_{k}$.
    However, since $f|_{\partial U}$ is ergodic and recurrent, then $\mu$-almost every $x\in\pi_{U,0}(K_{\infty})$ visits $\pi_{U,0}(K_{\infty})\subset X_{k}$ under forward iteration infinitely often. This is a contradiction with the definition of $K_{\infty}$, and finishes the proof of the first statement.
    The second statement is a straight-forward consequence of the first one, taking into account the definition of $f_{X_{k}}$-backward orbit associated with a $f$-backward orbit.
  3. (c)
    If follows from combining the previous statements. Indeed, taking into account that $\left\{X_{k}\right\}_{k}$ is a measurable partition of $\partial U$, according to (b), $\widetilde{\mu}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ can be written as a $f_{X_{k}}$-backward orbit, for some $k\geq 0$. Then, the statement follows applying (a).
$\square$
6.2 Pesin theory for the first return map. Proof of Theorem B
We shall start by rewriting Theorem B in terms of the space of backward orbits given by Rokhlin’s natural extension, as done in Section 5 for attracting basins.
Theorem 6.4.
(Inverse branches are well-defined almost everywhere) Let $f\colon\mathbb{C}\to\mathbb{C}$ be an entire function, and let $U$ be an invariant Fatou component for $f$. Let $\omega_{U}$ be a harmonic measure on $\partial U$. Assume
  1. (a)
    $f|_{\partial U}$ is recurrent with respect to $\omega_{U}$;
  2. (b)
    $\log\left|f^{\prime}\right|\in L^{1}(\omega_{U})$, and $\int_{\partial U}\log\left|f^{\prime}\right|d\omega_{U}>0$;
  3. (c)
    critical values of $f$ are finite.
Then, for $\widetilde{\mu}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\in\widetilde{\partial U}$, there exists $r\coloneqq r(\left\{x_{n}\right\}_{n})>0$ such that
  1. (i)
    the inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},r_{0})$;
  2. (ii)
    for every $r\in(0,r_{0})$, there exists $m\in\mathbb{N}$ such that
    $F_{m}(D(x_{0},r_{0}))\subset D(x_{0},r),$
    and diam $\textrm{\em diam }F^{j}_{m}(D(x_{0},r))\to 0$, as $j\to\infty$.
It is clear that Theorem 6.4 implies Theorem B.
Going one step further, using the partition $\left\{X_{k}\right\}_{k}$ developed in the previous section, we shall write Theorem 6.4 in terms of the first return maps $f_{X_{k}}\colon X_{k}\to X_{k}$. This is the content of Proposition 6.5. Applying Lemma 6.3, it is clear that Proposition 6.5 implies Theorem 6.4, and hence Theorem B.
Proposition 6.5.
(Generic inverse branches are well-defined for the first return map) Under the assumptions of Theorem 6.4, consider the partition $\left\{X_{k}\right\}_{k}$ as in Section 6.1. Then, for each $k\geq 0$, for $\widetilde{\mu_{k}}$-almost every backward orbit $\left\{x^{k}_{n}\right\}_{n}\in\widetilde{\partial U}$, there exists $r\coloneqq r(\left\{x^{k}_{n}\right\}_{n})>0$ such that
  1. (i)
    the inverse branch $F^{k}_{n}$ sending $x^{k}_{0}$ to $x^{k}_{n}$ is well-defined in $D(x_{0},r_{0})$;
  2. (ii)
    for every $r\in(0,r_{0})$, there exists $m\in\mathbb{N}$ such that
    $F^{k}_{m}(D(x^{k}_{0},r_{0}))\subset D(x^{k}_{0},r).$
6.3 Proof of Proposition 6.5
We start by proving the following lemma. Let $X_{k}$ be any of the sets in the partition of $\partial U$ considered previously. Recall that it is at a positive distance of any grand orbit finite point. Hence, we fix $\varepsilon>0$ such that, for all $x\in X_{k}$, the closed disk $\overline{D(x,\varepsilon)}$ does not contain grand orbit finite points.
Lemma 6.5.1.
(Blow-up property of $X_{k}$) For all $x\in X_{k}$, there exists $r\coloneqq r(x)>0$ such that
$f^{T}(D(x,r))\supset D(f^{T}(x),\varepsilon).$
Recall that $T=T(x)$ is the first return time of $x$ to the set $X_{k}$, and depends on $x$.
Proof.
Since $x\in\mathcal{J}(f)$ and repelling periodic points are dense in $\mathcal{J}(f)$, there exists a repelling periodic point $p\in D(x,R)$, of period say $q\geq 0$, for some $R>0$. Then, there exists a neigbourhood $V$ of $p$, $V\subset D(x,R)$, such that $V\subset f^{q}(V)$. Then, $V\subset f^{q}(V)\subset f^{2q}(V)\subset\dots$. In particular, since $\overline{D(f^{T}(x),\varepsilon)}$ is a compact subset of the plane which does not contain grand orbit finite points, by the blow-up property of the Julia set, there exists $n_{0}\geq 0$ such that for all $n\geq n_{0}$,
$f^{nq}(V)\supset D(f^{T}(x),\varepsilon).$
In particular, writing $n=mT$, for $m$ large enough, we get
$f^{nq}(V)=f^{T}(f^{mq}(V))\supset D(f^{T}(x),\varepsilon).$
Since $x\in V\subset f^{mq}(V)$ and $f^{mq}(V)$ is a compact subset of the plane ($f$ is assumed to be entire), we can take $r\coloneqq r(x)>0$ so that $f^{mq}(V)\subset D(x,r)$, it holds $f^{T}(D(x,r))\supset D(f^{T}(x),\varepsilon)$, as desired. $\square$
Next, we shall establish who are the ‘singular values’ for the first return map $f_{X_{k}}$. The previous lemma rules out the existence of asymptotic values for the first return map: any preimage under $f_{X_{k}}$ of the disk of radius $\varepsilon$ is a compact subset of the plane. Hence, the only problems when defining inverse branches come from the critical values of $f$. Note that, if there are no critical values of $f$ in $D(f_{X_{k}}(x),\varepsilon)$ and the first return time of $x$ is 1, then the corresponding branch of $f_{X_{k}}$ is well-defined in $D(f_{X_{k}}(x),\varepsilon)$. Inductively, if there are no critical values of $f^{n}$ in $D(f_{X_{k}}(x),\varepsilon)$ and the first return time of $x$ is $n$, then the corresponding branch of $f_{X_{k}}$ is well-defined in $D(f_{X_{k}}(x),\varepsilon)$. Hence, we observe an interplay between the points in the orbit of critical values of $f$ and the first return times, as the limitation to define the inverse branches of $f_{X_{k}}$.
Next we aim to give estimates on the first return times and the size of disks centered at ‘singular values of $f_{X_{k}}$’. This is the content of Lemma 6.5.2.
We use the following notation: let $\left\{v_{1},\dots,v_{N}\right\}$ be the critical values of $f$ (we assumed there are finitely many), and denote them by $CV(f)$. $T(x)$ stands for the first return time to $X_{k}$ of $x\in X_{k}$.
$A_{n}\coloneqq\left\{x\in X_{k}\colon T(x)=n\right\}$
$B_{n}\coloneqq\left\{x\in X_{k}\colon T(x)\geq n\right\}$
Lemma 6.5.2.
(Estimates on critical values and first returns) In the previous setting, the following holds.
  1. (2.1)
    If $\mu$ is a probability measure, then for $\widetilde{\mu_{k}}$-almost every $\left\{x_{n}\right\}_{n}$, there exists $n_{0}\coloneqq n_{0}(\left\{x_{n}\right\}_{n})\geq 1$ such that $x_{n}\notin B_{n_{0}}$, for all $n\geq n_{0}$.
  2. (2.2)
    If $\mu$ is an infinite $\sigma$-finite measure,
    $\sum\limits_{n}\mu_{k}(B_{n^{2}})<\infty.$
  3. (2.3)
    $\sum_{n}\mu_{k}(D(CV(f^{n^{2}}),C\cdot\lambda^{n}))<\infty$, for any $\lambda\in(0,1)$.
Proof.
We shall prove the previous statements separately.
  1. (2.1)
    It is a direct consequence of Birkhoff Ergodic Theorem 2.10 applied to First Return Maps, see e.g. [URM22, Prop. 10.3.1] which asserts that
    $\lim\limits_{n}\dfrac{T(x_{n})}{T(x_{n+1})}=1\hskip 14.22636pt\textrm{for }\widetilde{\mu_{k}}{\textrm{-almost every}}\left\{x_{n}\right\}_{n}.$
    (Note that [URM22, Prop. 10.3.1] is stated for forward orbits, but, acccording to Theorem 2.10 it also applies to backward orbits in Rokhlin’s natural extension.)
  2. (2.2)
    By Kac’s Lemma [URM22, Lemma 10.2.6],
[…]
    This implies that $\mu(A_{n})$ cannot decrease faster than $1/n^{2}$ on average. We claim that this implies that $\mu(B_{n^{2}})<1/n^{2}$ on average, so $\sum\limits_{n}\mu_{k}(B_{n^{2}})<\infty$. Indeed, inductively, if $\mu(B_{n^{2}})<1/n^{2}$, then
    $\mu(B_{{n+1}^{2}})-\mu(B_{n^{2}})=\mu(A_{n^{2}})+\dots+\mu(A_{{n+1}^{2}-1})>\frac{2n+1}{n^{4}},$
    implying that $\mu(B_{{n+1}^{2}})<1/(n+1)^{2}$. This ends the proof of the second statement.
  3. (2.3)
    First note that, since $X_{k}$ has finite measure, the measures $\mu_{k}$ and $\omega_{U}$ are comparable. Note also that $f^{n^{2}}$ has $n^{2}\cdot N$ critical values (where $N$ stands for the number of critical values of $f$). Then, applying the standard estimate of the harmonic measure of disks (Lemma 2.25), we have
    $\sum_{n}\mu_{k}(D(CV(f^{n^{2}}),C\cdot\lambda^{n}))<\sum_{n}\mu(D(CV(f^{n^{2}}),C\cdot\lambda^{n}))$
    $=$ $L\cdot\sum_{n}\omega_{U}(D(CV(f^{n^{2}}),C\cdot\lambda^{n}))=L\cdot\sum_{n}C^{1/2}\cdot n\cdot\lambda^{n/2}<\infty,$
    for some constant $L>0$, as desired.
$\square$
From here, the proof ends as the one of Theorem 5.1: proving that orbits under $f_{X_{k}}$ do not come close to the ‘critical values of $f_{X_{k}}$’, and finally constructing inductively the required inverse branches of $f_{X_{k}}$, which turn out to be a composition of inverse branches for the original map $f$, as explained in Section 6.1. For convenience, we outline the steps of the proof, although not giving all the details as in Theorem 5.1.
Set
$\chi\coloneqq\int_{X_{k}}\log\left|f^{\prime}_{X_{k}}\right|d\mu_{k}\in(0,+\infty),$
and let $M\in(e^{\frac{1}{4}\cdot\chi},1)$.
Lemma 6.5.3.
(Almost every orbit does not come close to singular values) For $\widetilde{\mu_{k}}$-almost every $\left\{x^{k}_{n}\right\}_{n}\in\widetilde{X_{k}}$, it holds
  1. (3.1)
    $x^{k}_{0}\notin\bigcup\limits_{c\in CV(f)}\bigcup\limits_{n\geq 0}f^{n}(c)$,
  2. (3.2)
    $\lim\limits_{n}\dfrac{1}{n}\log\left|(f_{X_{k}}^{n})^{\prime}(x^{k}_{n})\right|=\chi$,
  3. (3.3)
    inverse branches of $f_{X_{k}}$ are well-defined in $D(x^{k}_{n},\varepsilon\cdot M^{n})$, except for a finite number of $n$’s.
Proof.
Since the finite intersection of sets of full measure has full measure, it is enough to show that each of the conditions is satisfied in a set of full measure.
For condition (3.1), note that $\bigcup\limits_{c\in CV(f)}\bigcup\limits_{n\geq 0}f^{n}(c)$ is countable, and hence has zero $\mu_{k}$-measure. Requirement (3.2) follows from Birkhoff Ergodic Theorem 2.10 applied to the map $\log\left|f_{X_{k}}^{\prime}\right|$ (note that $\mu_{k}$ is an ergodic probability).
Condition (3.3) follows from Lemma 6.5.2 together with the first Borel-Cantelli Lemma 2.6. Indeed, if $\mu$ was a probability measure, then $x_{n}\notin B_{n_{0}}$, for all $n\geq n_{0}$. By the argument above, this implies that the ‘singular values of $f_{X_{k}}$’ come from the critical values of $f^{n_{0}}$ (which are in turn critical values of $f^{n}$ for all $n\geq n_{0}$). Therefore, if $D_{n}=(D(CV(f^{n^{2}}),\varepsilon\cdot M^{n}))$, then
$\sum_{n\geq 1}\widetilde{\mu_{k}}(\pi_{U,n}^{-1}(D_{n}))=\sum_{n\geq 1}\mu_{k}(D_{n})<\infty,$
implying that $x^{k}_{n}\notin D_{n}$, for all $n\geq n_{0}$ and $\widetilde{\mu_{k}}$-almost every backward orbit. (maybe taking $n$ larger). This already implies that inverse branches of $f_{X_{k}}$ are well-defined in $D(x^{k}_{n},\varepsilon\cdot M^{n})$, except for a finite number of $n$’s, for $\widetilde{\mu_{k}}$-almost every backward orbit.
It is left the case when $\mu$ was an infinite $\sigma$-finite measure. As before, by the first Borel-Cantelli Lemma 2.6, if $D_{n}=(D(CV(f^{n^{2}}),\varepsilon\cdot M^{n}))$, then
$\sum_{n\geq 1}\widetilde{\mu_{k}}(\pi_{U,n}^{-1}(D_{n}))=\sum_{n\geq 1}\mu_{k}(D_{n})<\infty,$
implying that $x^{k}_{n}\notin D_{n}$, for all $n\geq n_{0}$, for some $n_{0}$ and $\widetilde{\mu_{k}}$-almost every backward orbit. But according to (2.2) in Lemma 6.5.2,
$\sum_{n\geq 1}\widetilde{\mu_{k}}(\pi_{U,n}^{-1}(B_{n^{2}}))=\sum\limits_{n}\mu_{k}(B_{n^{2}})<\infty,$
so $x^{k}_{n+1}\notin B_{n^{2}}$, for all $n\geq n_{0}$ (maybe taking $n_{0}$ larger), and $\widetilde{\mu_{k}}$-almost every backward orbit. Thus, for all $n\geq n_{0}$, the return time of $x^{k}_{n+1}$ is less than $n^{2}$, so we only have to take into account the critical values of $f^{n^{2}}$ in order to define the inverse branch from $x^{k}_{n}$ to $x^{k}_{n+1}$. Since $x^{k}_{n}\notin D_{n}$, the claim follows. $\square$
Fix $\left\{x^{k}_{n}\right\}_{n}$ satifying the conditions of Lemma 6.5.1. By (3.2) and (3.3), there exists $n_{0}\in\mathbb{N}$ such that, for $n\geq n_{0}$, nverse branches of $f_{X_{k}}$ are well-defined in $D(x^{k}_{n},\varepsilon\cdot M^{n})$, and
$\left|(f^{n}_{X_{k}})^{\prime}(x^{k}_{n})\right|^{-\frac{1}{4}}<M^{n}<1.$
We set the following notation.
$S\coloneqq\sum_{n\geq 1}\left|(f^{n}_{X_{k}})^{\prime}(x^{k}_{n})\right|^{-\frac{1}{4}}$
$b_{n}\coloneqq\frac{1}{2}\left|(f^{n+1}_{X_{k}})^{\prime}(x^{k}_{n+1})\right|^{-\frac{1}{4}}$
$P=\prod_{n\geq 1}(1-b_{n})^{-1}$
Choose $r\coloneqq r(\left\{x^{k}_{n}\right\}_{n})>0$ such that
  1. (4.1)
    $32rP<\varepsilon$,
  2. (4.2)
    the branch $F^{k}_{n_{0}}$ of $f_{X_{k}}^{-n_{0}}$ sending $x^{k}_{0}$ to $x^{k}_{n_{0}}$ is well-defined in $D(x_{0},rP)$,
  3. (4.3)
    $\textrm{diam}\left(F^{k}_{n_{0}}(D(x_{0},r\prod\limits_{m\geq n_{0}}(1-b_{m})^{-1}))\right)\leq\varepsilon\cdot M^{n_{0}}$.
Using the same procedure as in Theorem 5.1 (Claim 5.2.2), one can prove inductively the following claim.
Claim 6.5.4.
(Inductive construction of the inverse branches) For every $n\geq n_{0}$, there exists a branch $F^{k}_{n}$ of $f^{-n}_{X_{k}}$ sending $x^{k}_{0}$ to $x^{k}_{n}$, defined in $D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1})$, and such that
$\textrm{\em diam}\left(F^{k}_{n}(D(x_{0},r\prod_{m\geq n}(1-b_{m})^{-1}))\right)\leq\varepsilon\cdot M^{n}.$
Letting $n\to\infty$, we get that all inverse branches of $f_{X_{k}}$ sending $x^{k}_{0}$ to $x^{k}_{n}$ are well-defined in $D(x_{0},r)$, with $r>0$. Moreover, as $n\to\infty$,
$\textrm{diam}\left(F^{k}_{n}(D(x_{0},r))\right)\leq\textrm{diam}\left(F^{k}_{n}(D(x_{0},r\prod\limits_{m\geq n}(1-b_{m})^{-1}))\right)\leq\varepsilon\cdot M^{n}\to 0.$
This ends the proof of Proposition 6.5. $\square$
7 Application: dynamics of centered inner functions on the unit circle. Corollary C
In this section we apply the techniques developed previously to a particular type of self-maps of the unit disk $\mathbb{D}$, the so-called inner functions.
Definition 7.1.
(Inner function) A holomorphic self-map of the unit disk $g\colon\mathbb{D}\to\mathbb{D}$ is an inner function if, for $\lambda$-almost every point $\xi\in\partial\mathbb{D}$,
$g^{*}(\xi)\coloneqq\lim\limits_{t\to 1^{-}}g(t\xi)\in\partial\mathbb{D}.$
If, in addition, $g(0)=0$, we say that $g$ is a centered inner function.
It follows from the definition that an inner function induces a measure-theoretical dynamical system $g^{*}\colon\partial\mathbb{D}\to\partial\mathbb{D}$ defined $\lambda$-almost everywhere. Morever, for centered inner functions, $g^{*}|_{\partial\mathbb{D}}$ preserves the Lebesgue measure $\lambda$ in ${\partial\mathbb{D}}$, and $g^{*}|_{\partial\mathbb{D}}$ is ergodic. Hence, the radial extension of a centered inner functions is a good candidate to perform Pesin theory. This is the content of Corollary C, and we devote this section to prove it.
Therefore, we shall see Corollary C as an application of the work done in Theorem A, for a particular class of inner functions (centered inner functions with finite entropy and thin singular values). On the one hand, the condition of $\log\left|g^{\prime}\right|\in L^{1}(\lambda)$ is known in the field as finite entropy, and it is a natural assumption to work with, which arises in many examples (see e.g. [DM91, Thm. B] and references therein). On the other hand, the geometric condition of thin singular values in $\partial\mathbb{D}$ (Def. 3.8) can be formulate for inner functions as if there exists $\varepsilon>0$, $\mu\in(0,1)$, $d\in\mathbb{N}$ and $v_{1},\dots,v_{k}\in{\partial}\mathbb{D}$ such that
  1. (a)
    for every $n\in\mathbb{N}$, there are at most $n^{d}$ singular values in
    $\left\{z\in\mathbb{D}\colon\left|z\right|>1-\mu^{n}\right\}\smallsetminus\bigcup_{i=1}^{k}D(v_{i},\varepsilon)$
    which are $\mu^{n}$-separated;
  2. (b)
    for $i=1,\dots,k$, there exists a horodisk $H_{i}$ of center $v_{i}\in\partial\mathbb{D}$ and radius $\delta_{i}>0$ such that all singular values in $D(v_{i},\varepsilon)\cap\mathbb{D}$ lie in the horodisk $H_{i}$.
For the definition of horodisk, see e.g. [Aba23, Def. 2.1.1], see also Figure 7.1.
[…]
Figure 7.1: Schematic representation of the notion of thin singular values for inner functions. Indeed, singular values in the orange region satisfy that there are at most $n^{d}$ singular values in $\left\{z\in\mathbb{D}\colon\left|z\right|>1-\mu^{n}\right\}$ which are $\mu^{n}$-separated; and all singular values in $D(v_{i},\varepsilon)\cap\mathbb{D}$(grey) lie in the horodisk $H_{i}$ (green), $i=1,\dots,k$.
Analogously to the previous sections, we will prove proof a generalized version of Corollary C, rewritten in terms of Rokhlin’s natural extension (Thm. 7.14). Although this section is devoted precisely to centered inner functions, we include some details on the distortion of inverse branches taken at points in the unit circle for general inner functions (not necessarily centered), which will be useful in the following sections (in particular, to prove density of periodic points). We include also basic background on iteration of inner functions, not necessarily centered.
7.1 Iteration of inner functions
Iteration of holomorphic self-maps of the unit disk
The asymptotic behaviour of iterates of a holomorphic self-map of the unit disk is essentially described by the Denjoy-Wolff theorem.
Theorem 7.2.
(Denjoy-Wolff, [Mil06, Thm. 5.2]) Let $g\colon\mathbb{D}\to\mathbb{D}$ be holomorphic, which is not the identity nor an elliptic Möbius transformation. Then, there exists a point $p\in\overline{\mathbb{D}}$, the Denjoy-Wolff point of $g$, such that for all $z\in\mathbb{D}$, $g^{n}(z)\to p$.
We say that $g$ is elliptic if the Denjoy-Wolff point $p$ belongs to the unit disk; otherwise, we say that $g$ is non-elliptic. Note that any elliptic inner function is conjugate to a centered inner function. Hence, the results presented in this section apply to all inner functions whose Denjoy-Wolff point lies in the unit disk.
Observe that, in the elliptic case, $g$ is holomorphic in a neighbourhood of the Denjoy-Wolff point $p\in\mathbb{D}$, which is fixed and it is either attracting (if $\left|g^{\prime}(p)\right|\in(0,1)$) or superattracting (if $g^{\prime}(p)=0$). In the former case, $g$ is conjugate to $z\mapsto\left|g^{\prime}(p)\right|z$ in a neighourhood of $p$ (by Koenigs Theorem, see e.g. [Mil06, Chap. 8]). In the latter case, the dynamics are conjugate to those of $z\mapsto z^{d}$, where $d$ stands for the local degree of $g$ at $p$ (by Böttcher Theorem, see e.g. [Mil06, Chap. 9]). An analogous result for the non-elliptic case is given by a result of Cowen [Cow81], which leads to a classification of non-elliptic self-maps of $\mathbb{D}$ in terms of the dynamics near the Denjoy-Wolff point.
Theorem 7.3.
(Cowen’s classification of self-maps of $\mathbb{D}$, [Cow81]) Let $g$ be a holomorphic self-map of $\mathbb{D}$ with Denjoy-Wolff point $p\in\partial\mathbb{D}$. Then, there exists a simply connected domain $V\subset\mathbb{D}$, there exists a domain $\Omega$ equal to $\mathbb{C}$ or ={Im $\mathbb{H}=\left\{\textrm{\em Im }z>0\right\}$, a holomorphic map $\psi\colon\mathbb{D}\to\Omega$, and a Möbius transformation $T\colon\Omega\to\Omega$, such that:
  1. (a)
    $V$ is an absorbing domain for $g$ in $\mathbb{D}$, i.e. $g(V)\subset V$ and for every compact set $K\subset\mathbb{D}$, there exists $n\geq 0$ such that $g^{n}(K)\subset V$;
  2. (b)
    $\psi(V)$ is an absorbing domain for $T$ in $\Omega$;
  3. (c)
    $\psi\circ g=T\circ\psi$ in $\mathbb{D}$;
  4. (d)
    $\psi$ is univalent in $V$.
Moreover, $T$ and $\Omega$ depend only on the map $g$, not on the fundamental set $V$. In fact (up to a conjugacy of $T$ by a Möbius transformation preserving $\Omega$), one of the following cases holds:
  • $\Omega=\mathbb{C}$, $T=\textrm{\em id}_{\mathbb{C}}+1$ (doubly parabolic type),
  • $\Omega=\mathbb{H}$, $T=\lambda\textrm{\em id}_{\mathbb{H}}$, for some $\lambda>1$ (hyperbolic type),
  • $\Omega=\mathbb{H}$, $T=\textrm{\em id}_{\mathbb{H}}\pm 1$ (simply parabolic type).
(a) Elliptic
(b) Doubly parabolic
(c) Hyperbolic
(d) Simply parabolic
Figure 7.2: The different types of convergence to the Denjoy-Wolff point.
Iteration of inner functions in the unit circle
Recall that, in general, inner functions present a highly discontinuous behaviour in $\partial\mathbb{D}$. Indeed, a point $\xi\in\partial\mathbb{D}$ is called a singularity of $g$ if $g$ cannot be continued analitically to any neighbourhood of $\xi$. Denote the set of singularities of $g$ by $E(g)$. If an inner function has finite degree, then it is a finite Blaschke product. In this case, $g$ has no singularities, and it extends to the Riemann sphere as a rational map. On the other hand, infinite degree inner functions must have at least one singularity. The following lemma characterizes the singularities of an inner function.
Lemma 7.4.
(Characterization of singularities, [Gar07, Thm. II.6.6]) Let $g\colon\mathbb{D}\to\mathbb{D}$ be an inner function. Then, $\xi\in E(g)$ if and only if, for any crosscut neighbourhood $N_{\xi}$ of $\xi$,
$\overline{g(N_{\xi})}=\overline{\mathbb{D}}.$
Throughout the paper, given any inner function $g$, we consider it continued to ^$\widehat{\mathbb{C}}\smallsetminus\overline{\mathbb{D}}$ by the reflection principle, and to $\partial\mathbb{D}\smallsetminus E(g)$ by analytic continuation. In other words, $g$ is considered as its maximal meromorphic extension
$g\colon\widehat{\mathbb{C}}\smallsetminus E(g)\to\widehat{\mathbb{C}}.$
On the other hand, one can consider the measure-theoretical dynamical system induced its the radial extension
$g^{*}\colon\partial\mathbb{D}\to\partial\mathbb{D},$
which is well-defined $\lambda$-almost everywhere. Recall that, if $g$ is an inner function, so is $g^{k}$ [BD99, Lemma 4], so the equality
$(g^{n})^{*}(\xi)=(g^{*})^{n}(\xi)$
holds $\lambda$-almost everywhere. Moreover, the radial extension $g^{*}$ is measurable (Thm. 2.15), and hence analyzable from the point of view of ergodic theory. The following is a recollection of ergodic properties of $g^{*}$.
Theorem 7.5.
(Ergodic properties of $g^{*}$) Let $g\colon\mathbb{D}\to\mathbb{D}$ be an inner function, with Denjoy-Wolff point $p\in\overline{\mathbb{D}}$, and let $g^{*}\colon\partial\mathbb{D}\to\partial\mathbb{D}$ be its radial extension to the unit circle. The following holds.
  1. (a)
    $g^{*}$ is non-singular. In particular, for $\lambda$-almost every $\xi\in\partial\mathbb{D}$, its infinite orbit under $g^{*}$, $\left\{(g^{n})^{*}(\xi)\right\}_{n}$, is well-defined.
  2. (b)
    $g^{*}$ is ergodic if and only if $g$ is elliptic or doubly parabolic.
  3. (c)
    If $g^{*}$ is recurrent, then it is ergodic. In this case, for every $A\in\mathcal{B}(\mathbb{D})$ with $\lambda(A)>0$, we have that for $\lambda$-almost every $\xi\in\partial\mathbb{D}$, there exists a sequence $n_{k}\to\infty$ such that $(g^{n_{k}})^{*}(\xi)\in A$. In particular, $\lambda$-almost every $\xi\in\partial\mathbb{D}$, $\left\{(g^{n})^{*}(\xi)\right\}_{n}$ is dense in $\partial\mathbb{D}$.
  4. (d)
    If $g$ is an elliptic inner function, then $g^{*}$ is recurrent.
  5. (e)
    The radial extension of a doubly parabolic inner function is not recurrent in general. However, if $g$ is doubly parabolic and the Denjoy-Wolff point $p$ is not a singularity for $g$, then $g^{*}$ is recurrent. Moreover, if $g$ is doubly parabolic and there exists $z\in\mathbb{D}$ and $r>1$ such that
    $\textrm{\em dist}_{\mathbb{D}}(g^{n+1}(z),g^{n}(z))\leq\frac{1}{n}+O\left(\frac{1}{n^{r}}\right),$
    as $n\to\infty$, then $g^{*}$ is recurrent.
  6. (f)
    Let $k$ be a positive integer. Then, $g^{k}$ is an inner function. Moreover, $g^{*}$ is ergodic (resp. recurrent) if and only if $(g^{k})^{*}$ is ergodic (resp. recurrent).
The previous theorem compiles different results in [Aar97, DM91, BFJK19, BEF+23]. A detailed explanation can be found in [Jov24, Thm. 3.10]. The next theorem deals with the existence of invariant measures.
Theorem 7.6.
(Invariant measures for $g^{*}|_{\partial\mathbb{D}}$, [DM91, Thm. A, C]) Let $g\colon\mathbb{D}\to\mathbb{D}$ be an inner function.
  1. (i)
    If $g$ is elliptic, assume 00 is the Denjoy-Wolff point of $g$. Then, the Lebesgue measure $\lambda$ is invariant under $g^{*}$.
  2. (ii)
    If $g$ is doubly-parabolic, assume $1$ is the Denjoy-Wolff point of $g$. Then, the $\sigma$-finite measure
[…]
    is invariant under $g^{*}$.
Note that the measure $\lambda_{\mathbb{R}}$ is the push-forward of the Lebesgue measure on $\mathbb{R}$ (up to multiplication by a constant) under any Möbius map transforming the upper half plane to the unit disk, and sending $\infty$ to 1.
7.2 Distortion of inverse branches in the unit circle
As explained before, inner functions may behave in a very discontinuous way along the unit circle, due to the possible presence of singularities. However, far from the singularities, the map is holomorphic and exhibits a strong symmetry, since it maps the unit circle to itself locally conformally. This can be used to give precise estimates on the distortion of the radial segment in terms of Stolz angles under an inner function.
Generalized radial arcs and Stolz angles
We shall start by reviewing some basic concepts, which describe different ways of approaching a point $\xi\in\partial\mathbb{D}$.
Definition 7.7.
(Crosscut neighbourhoods) Let $\xi\in\partial\mathbb{D}$. A crosscut $C$ is an open Jordan arc $C\subset\mathbb{D}$ such that $\overline{C}=C\cup\left\{a,b\right\}$, with $a,b\in\partial\mathbb{D}$. If $a=b$, we say that $C$ is degenerate; otherwise it is non-degenerate.
A crosscut neighbourhood of $\xi\in\partial\mathbb{D}$ is an open set $N\subset\mathbb{D}$ such that $\xi\in\partial N$, and $C\coloneqq\partial N\cap\mathbb{D}$ is a non-degenerate crosscut. We usually write $N_{\xi}$ or $N_{C}$, to stress the dependence on the point $\xi$ or on the crosscut $C$.
Moreover, recall that given $\xi\in\partial\mathbb{D}$ we have defined the Stolz angle of opening $\alpha\in\left(0,\pi/2\right)$ and length $\rho>0$ as the set of the form
$\Delta_{\alpha,\rho}(\xi)=\left\{z\in\mathbb{D}\colon\left|\textrm{Arg }\xi-\textrm{Arg }(\xi-z)\right|<\alpha,\left|z\right|>1-\rho\right\}.$
$\xi$
(a) Radial segment
text
$\xi$
(b) (Non-degenerate) crosscut
$\xi$
(c) Crosscut neighbourhood
$\xi$
(d) Stolz angle
text
Figure 7.3: Different sets related to $\xi\in\partial\mathbb{D}$.
Some times it is more convenient to work in the upper half-plane $\mathbb{H}$ rather than in the unit disk $\mathbb{D}$. The previous concepts can be defined analogously for points in $\partial\mathbb{H}$. In particular, the specific formulas for both the radial segment and Stolz angles at a point $x\in\mathbb{R}$ are
$R_{\rho}^{\mathbb{H}}(x)\coloneqq\left\{z\in\mathbb{H}\colon{\textrm{Im }}w<\rho,{\textrm{Re }}w=x\right\};$
$\Delta^{\mathbb{H}}_{\alpha,\rho}(x)\coloneqq\left\{z\in\mathbb{H}\colon{\textrm{Im }}w<\rho,\dfrac{\left|\textrm{Re }w-x\right|}{{\textrm{Im }}w}<\tan\alpha\right\}.$
A more flexible notion of radial segment and Stolz angle will be needed for our purposes.
Definition 7.8.
(Generalized radial arc and Stolz angle) Let $p\in\overline{\mathbb{D}}$ and let $\xi\in\partial\mathbb{D}$, $\xi\neq p$. Let $\rho>0$ and $\alpha\in(0,\pi/2)$.
  • If $p\in\mathbb{D}$, consider the Möbius transformation $M\colon\mathbb{D}\to\mathbb{D}$, $M(z)=\dfrac{p-z}{1-\overline{p}z}$. Then, the (generalized) radial segment $R_{\rho}(\xi,p)$ of length $\rho$ at $\xi$ is defined as the preimage under $M$ of the radial segment $R_{\rho}(M(\xi))$. Analogously, the (generalized) Stolz angle $\Delta_{\alpha,\rho}(\xi,p)$ of angle $\alpha$ and length $\rho$ is the preimage under $M$ of the Stolz angle $\Delta_{\alpha,\rho}(M(\xi))$. That is,
    $R_{\rho}(\xi,p)\coloneqq M^{-1}(R_{\rho}(M(\xi))),$
    $\Delta_{\alpha,\rho}(\xi,p)\coloneqq M^{-1}(\Delta_{\alpha,\rho}(M(\xi))).$
  • If $p\in\partial\mathbb{D}$, consider the Möbius transformation $M\colon\mathbb{D}\to\mathbb{H}$, $M(z)=i\dfrac{p+z}{p-z}$. Then, the (generalized) radial segment and Stolz angle at $\xi$ are defined as the preimages of the corresponding radial segment and Stolz angle at $M(\xi)\in\mathbb{R}$. That is,
    $R_{\rho}(\xi,p)\coloneqq M^{-1}(R_{\rho}^{\mathbb{H}}(M(\xi)))$
    $\Delta_{\alpha,\rho}(\xi,p)\coloneqq M^{-1}(\Delta_{\alpha,\rho}^{\mathbb{H}}(M(\xi))).$
See Figures 7.4 and 7.5.
Observe that $R_{\rho}(\xi)=R_{\rho}(\xi,0)$, and $\Delta_{\alpha,\rho}(\xi)=\Delta_{\alpha,\rho}(\xi,0)$. Note also that $R_{\rho}(\xi,p)$ is a curve landing non-tangentially at $\xi\in\partial\mathbb{D}$, while $\Delta_{\alpha,\rho}(\xi,p)$ is an angular neighbourhood of $\xi$, since Möbius transformations are conformal, and hence angle-preserving.
[…]
Figure 7.4: Radial arc and angular neighbourhood with respect to $p\in\mathbb{D}$.
[…]
Figure 7.5: Radial arc and angular neighbourhood with respect to $p\in\partial\mathbb{D}$.
Finally, the Lehto-Virtanen Theorem relates the existence of radial limits with the existence of limits taken along these generalized radial segments and Stolz angles, for a given map $h\colon\mathbb{D}\to\widehat{\mathbb{C}}$. Recall that we say that $h$ has angular limit at $\xi$ if, for any Stolz angle $\Delta$ at $\xi$, the limit
$\lim\limits_{z\to\xi,z\in\Delta}h(z)$
exists.
Theorem 7.9.
(Lehto-Virtanen, [Pom92, Sect. 4.1]) Let $h\colon\mathbb{D}\to\widehat{\mathbb{C}}$ be a meromorphic map omitting three values in $\widehat{\mathbb{C}}$. Let $\gamma$ be a curve in $\mathbb{D}$ landing at $\xi\in\partial\mathbb{D}$. If $h(\gamma)$ lands at a point $v\in\mathbb{C}$, then $h$ has angular limit at $\xi$ equal to $v$. In particular, radial and angular limits are the same.
Remark 7.10.
(Limit on generalized radial arcs and Stolz angles) Note that, in particular, the Lehto-Virtanen Theorem justifies that, for meromorphic maps omitting three values, it is equivalent to take the limit along the radial segment, than along any generalized radial segment (Def. 7.8). Likewise, the angular limit can be computed along generalized Stolz angles.
Singular values for inner functions
Given an inner function $g\colon\mathbb{D}\to\mathbb{D}$, consider it as its maximal meromorphic extension
$g\colon\widehat{\mathbb{C}}\smallsetminus E(g)\to\widehat{\mathbb{C}}.$
As defined in Section 3.1, we consider singular values of $g$ as those in which not all the inverse branches are well-defined. Due to the reflection symmetry along $\partial\mathbb{D}$ exhibit by inner functions, it is enough to consider
$SV(g,\overline{\mathbb{D}})=\left\{v\in\overline{\mathbb{D}}\colon v\textrm{ is singular}\right\}.$
Moreover, singular values in $\partial\mathbb{D}$ correspond to accumulation points of singular values in $\mathbb{D}$. Hence, the set of singular values of an inner function is determined by those in $\mathbb{D}$.
Proposition 7.11.
(Singular values in $\partial\mathbb{D}$, [Jov24, Prop. 4.2]) Let $g\colon\mathbb{D}\to\mathbb{D}$ be an inner function, and let $\xi\in\partial\mathbb{D}$. The following are equivalent.
  1. (a)
    There exists a crosscut $C$, with crosscut neighbourhood $N_{C}$ and $\xi\in\partial N_{C}$ such that $SV(g)\cap N_{C}=\emptyset$.
  2. (b)
    $v$ is regular, i.e. there exists $\rho\coloneqq\rho(\xi)>0$ such that all inverse branches $G_{1}$ of $g$ are well-defined in $D(\xi,\rho)$.
Distortion estimates around the unit circle
As an application of Koebe Distortion Theorem 2.1, together with the fact that $g$ preserves the unit circle $\lambda$-almost everywhere, we obtain the following result on how the radial segment is distorted in terms of Stolz angles when applying inverse branches. Note that the radius $\rho_{1}$ in which the estimate works depends only on domain of definition of the inverse branch, given by $\rho_{0}$, but not on the point $\xi\in\partial\mathbb{D}$, nor on the inverse branch $G_{1}$ considered.
Proposition 7.12.
(Control of radial limits in terms of Stolz angles, [Jov24, Prop. 4.5]) Let $g\colon\mathbb{D}\to\mathbb{D}$ be an inner function, with Denjoy-Wolff point $p\in\overline{\mathbb{D}}$. Let $\xi\in\partial\mathbb{D}$, $\xi\neq p$. Assume there exists $\rho_{0}>0$ such that $D(\xi,\rho_{0})\cap SV(g)\neq 0$. Then, for all $0<\alpha<\frac{\pi}{2}$, there exists $\rho_{1}\coloneqq\rho_{1}(\alpha,\rho_{0})<\rho_{0}$ such that all branches $G_{1}$ of $g^{-1}$ are well-defined in $D(\xi,\rho_{1})$ and, for all $\rho<\rho_{1}$,
$G_{1}(R_{\rho}(\xi,p))\subset\Delta_{\alpha,\rho}(G_{1}(\xi),p),$
where $R_{\rho}(\cdot,p)$ and $\Delta_{\alpha,\rho}(\cdot,p)$ stand for the generalized radial segment and Stolz angle with respect to $p$ (Def. 7.8).
It follows quite straightforward from Proposition 7.12, that, if $g^{*}|_{\partial\mathbb{D}}$ is recurrent and there exists a crosscut neighbourhood disjoint from the postsingular set, then iterated inverse branches are well-defined $\lambda$-almost everywhere, and radial segments are distorted in the previous fashion. This is the content of Theorem 7.13, and one has to see Corollary C as an improvement of it, only making assumptions on the singular values.
Theorem 7.13.
(Distortion of inverse branches on the unit circle, [Jov24, Thm. 4.6]) Let $g\colon\mathbb{D}\to\mathbb{D}$ be an inner function, such that $g^{*}|_{\partial\mathbb{D}}$ is recurrent. Assume there exists $\zeta\in\partial\mathbb{D}$ and a crosscut neighbourhood $N_{\zeta}$ of $\zeta$ such that $P(g)\cap N_{\zeta}=\emptyset$. Then, for $\lambda$-almost every $\xi\in\partial\mathbb{D}$, there exists $\rho_{0}\coloneqq\rho_{0}(\xi)>0$ such that all branches $G_{n}$ of $g^{-n}$ are well-defined in $D(\xi,\rho_{0})$. In particular, the set $E(g)$ of singularities of $g$ has $\lambda$-measure zero.
In addition, for all $0<\alpha<\frac{\pi}{2}$, there exists $\rho_{1}<\rho_{0}$ such that, for all $n\geq 0$, all branches $G_{n}$ of $g^{-n}$ are well-defined in $D(\xi,\rho_{1})$ and, for all $\rho<\rho_{1}$,
$G_{n}(R_{\rho}(\xi),p)\subset\Delta_{\alpha,\rho}(G_{n}(\xi),p),$
where $R_{\rho}(\cdot,p)$ and $\Delta_{\alpha,\rho}(\cdot,p)$ stand for the radial segment and the Stolz angle with respect to $p$ (Def. 7.8).
7.3 Pesin theory for centered inner functions with thin singular values. Proof of Corollary C
We shall start by rewriting Corollary C in terms of the Rokhlin’s natural extension of $\partial\mathbb{D}$. Indeed, $(\partial\mathbb{D},\mathcal{B}(\partial\mathbb{D}),\lambda)$ is a Lebesgue space (it is isomorphic, in the measure-theoretical sense, to the unit interval), and hence Theorem 2.12 guarantees the existence of Rokhlin’s natural extension. Thus, $\widetilde{\partial\mathbb{D}}$ is the space of backward orbits $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, with $g^{*}(\xi_{n+1})=\xi_{n}$ for $n\geq 0$, and $\widetilde{g^{*}}\colon\widetilde{\partial\mathbb{D}}\to\widetilde{\partial\mathbb{D}}$ is the automorphism which makes the following diagram commute.
{...}${\widetilde{\partial\mathbb{D}}}$${\widetilde{\partial\mathbb{D}}}$${\widetilde{\partial\mathbb{D}}}${...}{${{\scriptstyle\left\{\xi_{n+2}\right\}_{n}}}${${{\scriptstyle\left\{\xi_{n+1}\right\}_{n}}}${${{\scriptstyle\left\{\xi_{n}\right\}_{n}}}${...}${\partial\mathbb{D}}$${\partial\mathbb{D}}$${\partial\mathbb{D}}${...}${{\scriptstyle\xi_{n+2}}}$${{\scriptstyle\xi_{n+1}}}$${{\scriptstyle\xi_{n}}}$$\scriptstyle{\widetilde{g}}$$\scriptstyle{\widetilde{g}}$$\scriptstyle{\widetilde{g}}$$\scriptstyle{\widetilde{g}}$$\scriptstyle{\pi_{\mathbb{D},n}}$$\scriptstyle{\pi_{\mathbb{D},n}}$$\scriptstyle{\pi_{\mathbb{D},n}}$$\scriptstyle{g}$$\scriptstyle{g}$$\scriptstyle{g}$$\scriptstyle{g}$
This way, we rephrase Corollary C as follows.
Theorem 7.14.
(Pesin theory for centered inner function) Let $g\colon\mathbb{D}\to\mathbb{D}$ be an inner function, such that $g(0)=0$, and $\log\left|g^{\prime}\right|\in L^{1}(\partial\mathbb{D})$. Fix $\alpha\in(0,\pi/2)$. Assume singular values are thin in $\partial\mathbb{D}$. Then, for $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, there exists $\rho>0$ such that the inverse branch $G_{n}$ of $g^{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho)$, and, for all $\rho_{1}\in(0,\rho)$,
$G_{n}(R_{\rho_{1}}(\xi_{0}))\subset\Delta_{\alpha,\rho_{1}}(\xi_{n}).$
Moreover, the set of singularities $E(g)$ has zero $\lambda$-measure.
Using that $g^{*}|_{\partial\mathbb{D}}$ is ergodic and recurrent with respect to $\lambda$ (Thm. 7.5), it follows that for $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ and every set $A\subset\partial\mathbb{D}$ of positive measure, there exists a sequence $n_{k}\to\infty$ such that $\xi_{n_{k}}\in A$ (Prop. 2.13). Hence, it is clear that Theorem 7.14 implies Corollary C.
Proof of Theorem 7.14.
Proceeding exactly as in Theorem A, we find that, for $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, there exists $\rho_{0}>0$ such that the inverse branch $G_{n}$ of $g^{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho_{0})$. Note that all inverse branches $\left\{G_{n}\right\}_{n}$ are well-defined in a disk of uniform radius, namely in $D(\xi_{0},\rho_{0})$. Hence, we can apply Proposition 7.11, to see that, for all $\alpha\in(0,\pi/2)$ there exists $\rho<\rho_{0}$ such that for all $\rho_{1}\in(0,\rho)$,
$G_{n}(R_{\rho_{1}}(\xi_{0}))\subset\Delta_{\rho_{1}}(\xi_{n}).$
It is left to see that singularities have zero Lebesgue measure. Indeed, given any measurable set of positive measure in $\partial\mathbb{D}$, it is visited by $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}$ infinitely often. Now assume on the contrary that the set of singularities $E(g)$ has positive measure. Then, we can take $\left\{\xi_{n}\right\}_{n}$ visting $E(g)$ infinitely often, and satisfying that the inverse branches $\left\{G_{n}\right\}_{n}$ realizing such backward orbit are well-defined in $D(\xi_{0},\rho)$. Consider
$K\coloneqq\bigcup\limits_{n\geq 1}G_{n}(D(\xi_{0},\rho)).$
Since $K$ is open and $E(g)$ is closed, it is clear that $\lambda(K\cap E(g))>0$.
We claim that no point in $K$ is a singularity for $g$. Indeed, for any $\xi\in K$, there exists $n\geq 1$ such that $\xi\in G_{n}(D(\xi_{0},\rho))$. Hence,
$g|_{G_{n}(D(\xi_{0},\rho))}\colon G_{n}(D(\xi_{0},\rho))\longrightarrow G_{n-1}(D(\xi_{0},\rho))$
is univalent, so by Lemma 7.4, $\xi$ cannot be a singularity for $g$. This is a contradiction with the fact that $\lambda(K\cap E(g))>0$, and ends the proof of Corollary C. $\square$
8 Application: inner function associated to a Fatou component
Let $f\in\mathbb{K}$, and let $U$ be an invariant Fatou component for $f$, which we assume to be simply connected. Consider $\varphi\colon\mathbb{D}\to U$ to be a Riemann map. Then, $f\colon U\to U$ is conjugate by $\varphi$ to a holomorphic map $g\colon\mathbb{D}\to\mathbb{D}$, such that the diagram
${U}$${U}$${\mathbb{D}}$${\mathbb{D}}$$\scriptstyle{f}$$\scriptstyle{g}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi}$
commutes. It is well-known that $g$ is an inner function (see e.g. [Jov24, Prop. 5.6]). We say that $g$ is an inner function associated to $(f,U)$.
It is clear that the dynamics of $f|_{U}$ is inherited from the dynamics of $g|_{\mathbb{D}}$. Clearly, inner functions associated with attracting basins have their Denjoy-Wolff point $p\in\mathbb{D}$ (and hence are elliptic), whereas for parabolic basins and Baker domains $p\in\partial\mathbb{D}$. More precisely, inner functions associated with parabolic basins are always of doubly parabolic type, while Baker domains can exhibit any of the three possible behaviours [FH06], thus establishing a classification among them.
Moreover, the dynamics of $f|_{\partial U}$ can be studied by means of the radial extension $g^{*}|_{\partial\mathbb{D}}$, even though neither $\varphi$ nor $g$ extends continuously to $\partial\mathbb{D}$ in general. Following this idea, we analyze the correspondence between the radial limits for the inner function $g$ and the action of $f$ in $\partial U$, proving that $f\circ\varphi^{*}=\varphi^{*}\circ g^{*}$ $\lambda$-almost everywhere (Sect. 8.1). Then, we collect the ergodic properties of $f|_{\partial U}$ with respect to harmonic measure (Sect. 8.2), which were introduced in Theorem 3.3. These properties are needed in Section 9. Finally, we study how the determine a similar correspondence $f\circ\varphi^{*}=\varphi^{*}\circ g^{*}$ for infinite backward orbits, $\widetilde{\lambda}$-almost everywhere (Sect. 8.3).
8.1 Dynamics in the boundary of Fatou components
The conjugacy $f\circ\varphi=\varphi\circ g$ extends almost everywhere to $\partial\mathbb{D}$ by means of the radial extensions $\varphi^{*}\colon\partial\mathbb{D}\to\partial U$ and $g^{*}\colon\partial\mathbb{D}\to\partial\mathbb{D}$. More precisely, consider the following subsets of $\partial\mathbb{D}$.
$\Theta_{E}\coloneqq\left\{\xi\in\partial\mathbb{D}\colon\varphi^{*}(\xi)\in E(f)\right\}$
$\Theta_{\Omega}\coloneqq\left\{\xi\in\partial\mathbb{D}\colon\varphi^{*}(\xi)\in\Omega(f)\right\}$
Since $E(f)$ is countable, $\lambda(\Theta_{E})=0$, so $\lambda(\Theta_{\Omega})=1$. Moreover, the conjugacy $f\circ\varphi=\varphi\circ g$ extends for the radial extensions in $\Omega_{\Theta}$, as shown in the following lemma.
Lemma 8.1.
(Radial limits commute, [Jov24, Lemma 5.5]) Let $\xi\in\Theta_{\Omega}$, then $g^{*}(\xi)$ and $\varphi^{*}(g^{*}(\xi))$ are well-defined, and
$f(\varphi^{*}(\xi))=\varphi^{*}(g^{*}(\xi)).$
In this section we show that one can go further and relate backward orbits for the radial extension of the inner function $g^{*}$ with backward orbits for the boundary map $f|_{\partial U}$. Moreover, we will show how the natural extensions of $(\partial\mathbb{D},\lambda,g^{*})$ and $(\partial U,\omega_{U},f)$ are connected.
To do so, first we have to establish, in the spirit of Lemma 8.1, a relation between backward orbits for $g^{*}$ and backward orbits for $f$. More precisely, we prove that backward orbits associated to a well-defined sequence of inverse branches indeed commute by the Riemann map, as long as the radial limit at the initial point exists.
Proposition 8.2.
(Backward orbits commute) Let $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ be a backward orbit for $g^{*}$ such that there exists $\rho>0$ such that the inverse branch $G_{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho)$. Assume $\varphi^{*}(\xi_{0})$ exists. Then, $\varphi^{*}(\xi_{n})$ exists for all $n\geq 1$.
Moreover, for $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, $\left\{\varphi^{*}(\xi_{n})\right\}_{n}$ is well-defined and
$f(\varphi^{*}(\xi_{n+1}))=\varphi^{*}(g^{*}(\xi_{n+1}))=\varphi^{*}(\xi_{n}),$
for all $n\geq 0$.
Proof.
We note that, using an inductive argument, it is enough to prove that, if $\varphi^{*}(\xi_{0})$ exists and $\xi_{1}\in\partial\mathbb{D}$ is such that $g^{*}(\xi_{1})=\xi_{0}$, then $\varphi^{*}(\xi_{1})$ is well-defined, and
$f(\varphi^{*}(\xi_{1}))=\varphi^{*}(g^{*}(\xi_{1}))=\varphi^{*}(\xi_{0}).$
Let $R_{\xi_{1}}$ be the radius at $\xi_{1}$. Then, $g(R_{\xi_{1}})$ is a curve landing non-tangentially at $\xi_{0}$ (by Prop. 7.11). By the Lehto-Virtanen Theorem 7.9, radial and angular limits coincide, and hence $\varphi(g(R_{\xi_{1}}))=f(\varphi(R_{\xi_{1}}))$ is a curve landing non-tangentially at $\varphi^{*}(\xi_{0})$. Therefore, since preimages of a point under a holomorphic map are discrete and the singularities of $f$ are countable, $\varphi(R_{\xi_{1}})$ cannot accumulate on a continuum. Hence, $\varphi(R_{\xi_{1}})$ lands at a point in ^$\widehat{\partial}U$, as desired. $\square$
8.2 Ergodic properties of the boundary map $f\colon\partial U\to\partial U$
The previous conjugacy $f\circ\varphi=\varphi\circ g$ allows us to derive the ergodic properties of $f|_{\partial U}$ (with respect to the harmonic measure $\omega_{U}$) from the ergodic properties of $g^{*}|_{\partial\mathbb{D}}$ (with respect to the Lebesgue measure). Ergodic properties of $g^{*}|_{\partial\mathbb{D}}$, as well as their extension to $f|_{\partial U}$, have been widely studied [Aar78, DM91, BFJK19, Jov24]. The following theorem summarizes these well-known results (see e.g. [Jov24, Thm. 5.7] and the references therein).
Theorem 8.3.
(Ergodic properties of the boundary map) Let $f\in\mathbb{K}$, and let $U$ be an invariant simply connected Fatou component for $f$. Let $g$ be an inner function associated to $(f,U)$. Then, the following are satisfied.
  1. (i)
    If $U$ is either an attracting basin, a parabolic basin, or a Siegel disk, then $g^{*}|_{\partial\mathbb{D}}$ is ergodic and recurrent with respect to the Lebesgue measure $\lambda$.
  2. (ii)
    If $U$ is a doubly parabolic Baker domain, $g^{*}|_{\partial\mathbb{D}}$ is ergodic with respect to $\lambda$. In addition, assume one of the following conditions is satisfied.
    1. (a)
      $f|_{U}$ has finite degree.
    2. (b)
      Relevant singular values $SV(f,U)$ are compactly contained in $U$.
    3. (c)
      The Denjoy-Wolff point of $g$ is not a singularity for $g$.
    4. (d)
      There exists $z\in U$ and $r>1$ such that
      $\textrm{\em dist}_{U}(f^{n+1}(z),f^{n}(z))\leq\frac{1}{n}+O\left(\frac{1}{n^{r}}\right),$
      as $n\to\infty$, where $\textrm{\em dist}_{U}$ denotes the hyperbolic distance in $U$.
    Then, $g^{*}|_{\partial\mathbb{D}}$ is recurrent with respect to $\lambda$.
  3. (iii)
    If $g^{*}|_{\partial\mathbb{D}}$ is ergodic (resp. recurrent) with respect to $\lambda$, so is $f|_{{\partial}U}$ with respect to $\omega_{U}$.
  4. (iv)
    Let $k$ be a positive integer. Then, the inner function associated to $(f,U)$ has the same ergodic properties than the inner function associated to $(f^{k},U)$.
The following result concern the existence of invariant measures for $f|_{\partial U}$, built as the push-forward measures of those invariants for the radial extension of the associated inner function (7.6).
Corollary 8.4.
(Invariant measures for $f|_{\partial U}$) Let $f\in\mathbb{K}$, and let $U$ be an invariant simply connected Fatou component for $f$.
  1. (i)
    If $U$ is an attracting basin or a Siegel disk with fixed point $p\in U$, the harmonic measure $\omega_{U}(p,\cdot)$ is invariant under $f$.
  2. (ii)
    If $U$ is a parabolic basin or a doubly-parabolic Baker domain, with convergence point $p\in\widehat{\partial}U$. Then, the push-forward of the measure
[…]
    under the Riemann map $\varphi\colon\mathbb{D}\to U$, $\varphi^{*}(1)=p$, i.e.
    $\mu\coloneqq(\varphi^{*})_{*}\lambda_{\mathbb{R}},$
    is invariant under $f$. The support of $\mu$ is ^$\widehat{\partial}U$.
Note that, in particular, the measure $\mu$ is $\sigma$-finite. Hence, both in the case of an attracting basin, or a parabolic basin or a recurrent Baker domain, there exists an invariant measure in $f|_{\partial U}$, which is either finite or $\sigma$-finite, so we can consider the Rokhlin’s natural extension of this system (see Sect. 8.3).
Proof of Corollary 8.4.
  1. (i)
    Observe that the the harmonic measure $\omega_{U}=\omega_{U}(p,\cdot)$ is the push-forward of the Lebesgue measure $\lambda$ by a Riemann map $\varphi\colon\mathbb{D}\to U$, with $\varphi(0)=p$, which is invariant under $g^{*}$. Hence, using Lemma 8.1,
    $\omega_{U}(f^{-1}(A))=\lambda((\varphi^{*})^{-1}(f^{-1}(A)))=\lambda((g^{*})^{-1}((\varphi^{*})^{-1}(A)))=\lambda((\varphi^{*})^{-1}(A))=\omega_{U}(A),$
    for any measurable set $A\subset\widehat{\partial}U$, as desired.
  2. (ii)
    For the invariance of the measure, apply the same argument as before. The fact that supp $\textrm{supp }\mu=\widehat{\partial}U$ comes from the absolute continuity between $\lambda$ and $\lambda_{\mathbb{R}}$, together with supp $\textrm{supp }\omega_{U}=\widehat{\partial}U$. Indeed, let $x\in\widehat{\partial}U$ and $r>0$, then $\omega_{U}(D(x,r))>0$, so $\lambda((\varphi^{*})^{-1}(D(z,r)))>0$. By the absolute continuity between $\lambda$ and $\lambda_{\mathbb{R}}$, $\lambda_{\mathbb{R}}((\varphi^{*})^{-1}(D(z,r)))>0$, so $\mu(D(x,r))>0$, as desired.
$\square$
8.3 Generic inverse branches commute
We are interested now in the interplay between the backward orbits for the associated inner function $g$, and the backward orbits for $f$ in the dynamical plane.
Throughout the section, we let $U$ be either an attracting or parabolic basin, or a doubly parabolic Baker domain for a function $f\in\mathbb{K}$, and let $g\colon\mathbb{D}\to\mathbb{D}$ be its associated inner function. Hence, according to Theorem 7.6 and Corollary 8.4, there exists invariant measures for $(\widetilde{\partial\mathbb{D}},\widetilde{\lambda},\widetilde{g^{*}})$ and for $(\partial U,\omega_{U},f)$. Therefore, according to Section 2.5, we can consider the natural extension $(\widetilde{\partial\mathbb{D}},\widetilde{\lambda},\widetilde{g^{*}})$ of $(\partial\mathbb{D},\lambda,g^{*})$, given by the projecting morphisms $\left\{\pi_{\mathbb{D},n}\right\}_{n}$, and the natural extension $(\widetilde{\partial U},\widetilde{\omega_{U}},\widetilde{f})$ of $(\partial U,\omega_{U},f)$, given by the projecting morphisms $\left\{\pi_{U,n}\right\}_{n}$. We are interested in relating both natural extensions.
In views of Proposition 8.2, it is clear that the transformation
$\widetilde{\varphi^{*}}$ $\colon\widetilde{\partial\mathbb{D}}\longrightarrow\widetilde{\partial U}$
$\left\{\xi_{n}\right\}_{n}\mapsto\left\{\varphi^{*}(\xi_{n})\right\}_{n}$
is well-defined, and the following diagram, which shows in a more detailed way the actions of the morphisms on the considered backward orbits, commutes almost everywhere.
[…]
Now we claim that $\widetilde{\varphi^{*}}$ is measure-preserving. Indeed, one may take a basis for the $\sigma$-algebra in $\widetilde{\partial U}$ made of sets of the form $\pi^{-1}_{U,n}(A)$, where $A\subset\partial U$ measurable, and $n\geq 0$. It is enough to prove that $\widetilde{\varphi^{*}}$ preserves the measure of these sets. Indeed, using that $\varphi^{*}\circ\pi_{\mathbb{D},n}=\pi_{{U},n}\circ\widetilde{\varphi^{*}}$ $\widetilde{\lambda}$-almost everywhere, we have
$\widetilde{\omega_{U}}(\pi^{-1}_{U,n}(A))=\omega_{U}(A)=\lambda(\varphi^{*}(A))=\widetilde{\lambda}(\pi^{-1}_{\mathbb{D},n}\circ(\varphi^{*})^{-1}(A))=\widetilde{\lambda}((\widetilde{\varphi^{*}})^{-1}\circ\pi^{-1}_{U,n}(A)),$
where $A\subset\partial U$ measurable, and $n\geq 0$, as desired. In other words, $\widetilde{\omega_{U}}$ is the push-forward of $\widetilde{\lambda}$ by $\widetilde{\varphi^{*}}$.
Hence, the following diagram
${(\partial\mathbb{D},\lambda,g^{*})}$${(\widetilde{\partial\mathbb{D}},\widetilde{\lambda},\widetilde{g^{*}})}$${(\partial U,\omega_{U},f)}$${(\widetilde{\partial U},\widetilde{\omega_{U}},\widetilde{f}).}$$\scriptstyle{\varphi^{*}}$$\scriptstyle{\widetilde{\varphi^{*}}}$$\scriptstyle{\left\{\pi_{\mathbb{D},n}\right\}_{n}}$$\scriptstyle{\left\{\pi_{U,n}\right\}_{n}}$
commutes almost everywhere.
Proposition 8.5.
(Generic inverse branches commute) Let $f\in\mathbb{K}$, and let $U$ be an invariant simply connected Fatou component for $f$. Let $\varphi\colon\mathbb{D}\to U$ be a Riemann map, and let $g\colon\mathbb{D}\to\mathbb{D}$ be the inner function associated to $(f,U)$ by $\varphi$. Assume the following conditions are satisfied.
  1. (a)
    For $\widetilde{\omega_{U}}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$, there exists $r>0$ such that the inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},r)$.
  2. (b)
    For $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ there exists $\rho>0$ such that the inverse branch $G_{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho)$.
Then, $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ there exists $\rho,r>0$ such that the inverse branch $G_{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho)$, the inverse branch $F_{n}$ sending $\varphi^{*}(\xi_{0})$ to $\varphi^{*}(\xi_{n})$ is well-defined in $D(\varphi^{*}(\xi_{0}),r)$, and $\varphi^{*}\circ G_{n}(\xi_{0})=F_{n}\circ\varphi^{*}(\xi_{0})$, for all $n\geq 0$.
We note that, if $\widetilde{\omega_{U}}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$ satisfies an additional property (such as the ones proved in Theorem A), then it is straightforward to see that, for $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, the backward orbit $\left\{x_{n}\coloneqq\varphi^{*}(\xi_{n})\right\}_{n}$ satisfies this additional property.
Proof of Proposition 8.5.
The proof follows directly from the previous construction. Indeed, one shall write the first assumption as: for $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, there exists $r>0$ such that the inverse branch $F_{n}$ sending $\varphi^{*}(\xi_{0})$ to $\varphi^{*}(\xi_{n})$ is well-defined in $D(\varphi^{*}(\xi_{0}),r)$. Since the intersection of sets of full measure has full measure, we have that inverse branches $G_{n}$ and $F_{n}$ are well-defined along the backward orbit of $\xi_{0}$ and $\varphi^{*}(\xi_{0})$. By Proposition 8.2, such inverse branches commute. $\square$
Remark.
It follows from the previous construction that in Theorems A and B one can find first the backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ and define the backward orbit in the dynamical plane as their image by $\varphi^{*}$. Moreover, one can choose a countable collection of sets $\left\{K_{k}\right\}_{k}\subset\partial\mathbb{D}$ and ask that there exists a sequence $n_{k}\to\infty$ with $\xi_{n_{k}}\in K_{k}$.
9 Application: periodic boundary points. Corollary D
One possible application of Pesin theory in holomorphic dynamics is to prove that for some invariant Fatou components, periodic points are dense in their boundary. This was done in the seminal paper of Przytycki and Zdunik [PZ94] for simply connected attracting basins of rational maps (note that in this paper it is proved that periodic points are dense in the boundary of every attracting or parabolic basin of a rational map, but the proof relies on a different technique). In the spirit of [JF23] and [Jov24], we aim to prove a similar result for transcendental maps.
The goal in this section is to prove Corollary D, which states that, under the hyptothesis of either Theorem A or Theorem B, plus an extra hyptotheses on the singular values in $U$, accessible periodic boundary points are dense.
In view of the theory developed in the previous sections based on working in the space of backward orbits given by Rokhlin’s natural extension, we shall formulate an alternative (and more natural) version of Corollary D, in terms of backward orbits.
Theorem 9.1.
(Periodic points are dense) Let $f\in\mathbb{K}$, and let $U$ be an invariant simply connected Fatou component for $f$. Let $\varphi\colon\mathbb{D}\to U$ be a Riemann map, and let $g\colon\mathbb{D}\to\mathbb{D}$ be the inner function associated to $(f,U)$ by $\varphi$. Assume the following conditions are satisfied.
  1. (a)
    For $\widetilde{\omega_{U}}$-almost every backward orbit $\left\{x_{n}\right\}_{n}\subset\partial U$, there exists $r>0$ such that the inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},r)$, for every subsequence $\left\{x_{n_{j}}\right\}_{j}$ with $x_{n_{j}}\in D(x_{0},r)$, diam $\textrm{\em diam }F_{n_{j}}(D(x_{0},r))\to 0$, as $j\to\infty$.
  2. (b)
    For $\widetilde{\lambda}$-almost every backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ there exists $\rho>0$ such that the inverse branch $G_{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho)$.
Then, accessible periodic points are dense in $\partial U$.
However, as promised in the introduction, we aim to give a proof of the density of periodic boundary points which does not use Rohklin’s natural extension. To do so, we state Theorem 9.1 in a slightly different (and stronger) way.
Theorem 9.2.
(Periodic points are dense) Let $f\in\mathbb{K}$, and let $U$ be an invariant simply connected Fatou component for $f$. Let $\varphi\colon\mathbb{D}\to U$ be a Riemann map, and let $g\colon\mathbb{D}\to\mathbb{D}$ be the inner function associated to $(f,U)$ by $\varphi$. Assume that for every countable sequence of measurable sets $\left\{K_{k}\right\}_{k}\subset\partial\mathbb{D}$ with $\lambda(K_{k})>0$ and $\lambda$-almost every $\xi\in\partial\mathbb{D}$, there exists a backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, such that
  1. (a)
    $\xi=\xi_{0}$ and there exists $\rho>0$ such that the inverse branch $G_{n}$ sending $\xi_{0}$ to $\xi_{n}$ is well-defined in $D(\xi_{0},\rho)$, and there exists $n_{k}\to\infty$ with $\xi_{n_{k}}\in K_{k}$;
  2. (b)
    for the backward orbit $\left\{x_{n}\coloneqq\varphi^{*}(\xi_{n})\right\}_{n}\subset\partial U$, there exists $r>0$ such that the inverse branch $F_{n}$ sending $x_{0}$ to $x_{n}$ is well-defined in $D(x_{0},r)$, for every subsequence $\left\{x_{n_{j}}\right\}_{j}$ with $x_{n_{j}}\in D(x_{0},r)$, diam $\textrm{\em diam }F_{n_{j}}(D(x_{0},r))\to 0$, as $j\to\infty$.
Then, accessible periodic points are dense in $\partial U$.
According to Proposition 8.5, it is clear that Theorem 9.2 implies 9.1. We show now how to deduce Corollary D from Theorem 9.1, and later we give the proof of Theorem 9.2.
Proof of Corollary D.
On the one hand, it is clear that, by the conclusion of Theorem A and Theorem B, the second requirement of Theorem 9.1 holds.
On the other hand, we have to see the assumption of the existence of a crosscut neighbourhood $N_{C}$ in $U$ with $N_{C}\cap P_{U}(f)=\emptyset$ implies (b). Indeed, $\varphi^{-1}(N_{C})$ is a crosscut neighbourhood in $\mathbb{D}$ which contains no postsingular value for the inner function. Applying Theorem 7.13, for $\lambda$-almost $\xi\in\partial\mathbb{D}$, there exists $\rho>0$ such that, for all $n\geq 0$, all inverse branches of $g^{n}$ are well-defined in $D(\xi,\rho)$. Denote this set of backward orbits by $\widetilde{A}$. We have to see that $\widetilde{A}$ has full $\widetilde{\lambda}$-measure in $\widetilde{\partial\mathbb{D}}$. Indeed, note that
$\widetilde{A}=\pi_{\mathbb{D},0}^{-1}(\pi_{\mathbb{D},0}(\widetilde{A})),$
that is, the set $\widetilde{A}$ is made of all backward orbit with initial point in $\pi_{\mathbb{D},0}(\widetilde{A})$. Since $\lambda(\pi_{\mathbb{D},0}(\widetilde{A}))=1$ and $\pi_{\mathbb{D},0}$ is measure-preserving, this already implies the requirement (b) in Theorem 9.1. $\square$
9.1 Proof of Theorem 9.2
Let $x\in\partial U$ and $R>0$, we have to see that $f$ has a repelling periodic point in $D(x,R)\cap\partial U$, which is accessible from $U$.
We split the proof in several intermediate lemmas. We start by proving the existence of a backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$ such that for both $\left\{\xi_{n}\right\}_{n}$ and $\left\{\varphi^{*}(\xi_{n})\right\}_{n}$ the corresponding inverse branches are well-defined (and conformal), and certain estimates on the contraction are achieved.
In the sequel, we fix $\alpha\in(0,\pi/2)$, and we take all Stolz angles of opening $\alpha$, although in the notation we omit the dependence.
Lemma 9.2.1.
There exists a backward orbit $\left\{\xi_{n}\right\}_{n}\subset\partial\mathbb{D}$, and constants $m\in\mathbb{N}$, $0<\rho_{m}\leq\rho$, and $r\in(0,R/2)$ such that:
  1. (1.1)
    $x_{0}\coloneqq\varphi^{*}(\xi_{0})$ and $x_{m}\coloneqq\varphi^{*}(\xi_{m})$ are well-defined, and $x_{0}\in D(x,R/2)$ and $x_{m}\in D(x_{0},r/3)$;
  2. (1.2)
    the inverse branch $F_{m}$ of $f^{m}$ sending $x_{0}$ to $x_{m}$ is well-defined in $D(x_{0},r)$, and diam $\textrm{\em diam }F_{m}(D(x_{0},r))<r/3$;
  3. (1.3)
    the inverse branch $G_{m}$ of $g^{m}$ sending $\xi_{0}$ to $\xi_{m}$ is well-defined in $D(\xi_{0},\rho_{m})$, and satisfies
    $G_{m}(R_{\rho_{m}}(\xi_{0}))\subset\Delta_{\rho_{m}}(\xi_{m});$
  4. (1.4)
    $\Delta_{\rho}(\xi_{0})\cap\Delta_{\rho}(\xi_{m})\neq\emptyset$, and, if $z\in\Delta_{\rho}(\xi_{0})\cup\Delta_{\rho}(\xi_{m})$, then $\varphi(z)\in D(x_{0},r)$.
Proof.
Let $A_{n}=D(x^{n},r_{n})$ be a countable basis for $D(x,R)$ with the Euclidean topology, where $x^{n}\in\partial U$ and $A_{n}\subset D(x,R)$.
In order to apply the hyptothesis of the theorem, we shall construct an appropriate countable sequence of measurable sets $\left\{K_{k}\right\}_{k}$ of $\partial\mathbb{D}$. We do it as follows.
For all $n\geq 0$, let
$K^{n}=\left\{\xi\in\partial\mathbb{D}\colon\varphi^{*}(\xi)\in D(x^{n},{r_{n}}/{2})\right\}.$
By Lemma 2.20, it is clear that $\lambda(K^{n})>0$. By the Lehto-Virtanen Theorem 7.9, the angular limit exists whenever the radial limit exists. Therefore, there exists $\rho_{n}>0$ small enough so that
$K^{n}_{\rho_{n}}=\left\{\xi\in K^{n}\colon\Delta_{\rho_{n}}(\xi)\subset D(x^{n},{r_{n}}/{2})\right\}$
has positive $\lambda$-measure. By Theorem 2.5, we can assume that every point in $K^{n}_{\rho_{n}}$ is a Lebesgue density point for $K^{n}_{\rho_{n}}$. Then, if we take $\xi^{n}\in K^{n}_{\rho_{n}}$, there exists a circular interval $I_{\xi^{n}}$ around $\xi^{n}$ such that for any $\zeta_{1},\zeta_{2}\in I_{\xi^{n}}$,
$\Delta_{\rho_{n}}(\zeta_{1})\cap\Delta_{\rho_{n}}(\zeta_{2})\neq\emptyset.$
Then, $K^{n}_{\rho_{n}}\cap I_{\xi^{n}}$ has positive $\lambda$-measure. Note that this property only depends on the length of the interval, as long as $\xi^{n}$ is a Lebesgue density point for $K^{n}_{\rho_{n}}$. Then, it is clear that there exist finitely many circular intervals $I_{1}^{n},\dots,I^{n}_{i_{n}}$ with this property.
Let
$K^{1,n}_{*,i}\coloneqq K^{n}_{\rho_{n}}\cap I^{n}_{i},\hskip 14.22636pti=1,\dots,i_{n},$
$K^{1,n}_{*}\coloneqq\left\{K^{1,n}_{*,1},\dots,K^{1,n}_{*,i_{n}}\right\}.$
Then, we define the set $K^{j,n}_{*}$, as before, but replacing $\rho_{n}$ by $\rho_{n}/2^{j}$.
Having introduced all this notation of the sets $\left\{K^{j,n}_{*}\right\}_{n,j}$, we arrange the sequence $\left\{K_{k}\right\}_{k}$ as follows. We construct this sequence of sets inductively, adding at each step finitely many sets. Indeed, let us start by putting the block $K^{1,1}_{*}\coloneqq\left\{K^{1,1}_{*,1},\dots,K^{1,1}_{*,i_{1}}\right\}$ as the first elements of the sequence. Then, for the $k$-th step of the induction, we consider $A_{k}$ and let $A_{k_{1}},\dots,A_{k_{n}}$ be all the sets of $A_{1},\dots,A_{n}$ such that $A_{n}\subset A_{k_{i}}$. Then, we add to the sequence the blocks
$K^{1,k_{1}}_{*},\dots K^{1,k_{n}}_{*},\dots,K^{k,k_{1}}_{*},\dots,K^{k,k_{n}}_{*}.$
Basically, the idea is that, when one set is in the sequence $\left\{K_{k}\right\}_{k}$ for the first time, then it appears infinitely often. Moreover, the set of points in $\left\{K_{k}\right\}_{k}$ has measure $\lambda((\varphi^{*})^{-1}(D(x,R)))$. Indeed, the set of points in $\partial\mathbb{D}$ for which the radial limit exists has full measure. Let $\zeta$ be one of such points. Then, $\varphi^{*}(\zeta)\in A_{j}$, for some $j$, and for $\rho>0$ small enough, $\Delta_{\rho}(\zeta)\subset A_{j}$. Then, there exists $n\geq 0$ such that $A_{n}\subset A_{j}$ and $\rho<\rho_{j}/2^{n}$, so $\zeta\in K^{n}_{*,k_{n}}$, as desired.
By the assumption of the theorem, for $\lambda$-almost every $\xi_{0}\in\partial\mathbb{D}$, there exists a backward orbit $\left\{\xi_{n}\right\}_{n}$ such that the hypothesis on the definition of the inverse branches for $\left\{\xi_{n}\right\}_{n}$ and $\left\{x_{n}\coloneqq\varphi^{*}(\xi_{n})\right\}_{n}$ are accomplished, and there exists $n_{k}\to\infty$ with $\xi_{n_{k}}\in K_{k}$.
Without loss of generality, we assume $\xi_{0}$ is chosen so that $x_{0}\in D(x,R/2)$. Let $r>0$ be such that the inverse branches realizing the backward orbit $\left\{x_{n}\right\}_{n}$ are well-defined in $D(x_{0},r)$. There is no loss of generality on assuming $r\in(0,R/2)$.
On the one hand, since $\left\{A_{n}\right\}_{n}$ is a basis for $D(x,R)$, there exists $n_{0}$ such that
$x_{0}\in A_{n_{0}}\subset D(x_{0},r/3),$
and $\xi_{0}\in K^{n_{0}}_{*}$, by the previous remark. In particular, for $\rho_{n_{0}}$,
$\Delta_{\rho_{n_{0}}}(\xi_{0})\subset A_{n_{0}}\subset D(x_{0},r/3).$
On the other hand, by the construction of the sets $\left\{K_{n}\right\}_{n}$, the backward orbit visits $D(x_{0},r)$ infinitely many times. Let $n_{1}$ be large enough so that, for all $n\geq n_{1}$, if $x_{n}\in D(x_{0},r)$, then diam $\textrm{diam }F_{n}(D(x_{0},r))<r/3$.
By the construction of the sets $\left\{K_{n}\right\}_{n}$, there exists $m\geq\max\left\{n_{0},n_{1}\right\}$ such that $\xi_{m}\in K^{n_{0}}_{*}$. Hence, we take $r>0$, $\rho=\rho_{n_{0}}$, and $\xi_{0}$ and $\xi_{m}$ as above, and define $\rho_{m}>0$ as the radius such that the inverse branch $G_{m}$ sending $\xi_{0}$ to $x_{m}$ is defined around $\xi_{0}$ (such a radius exists by our assumptions on the orbit $\left\{\xi_{n}\right\}_{n}$). We have to check that, with these choices, the requirements are accomplished.
First, by the choice of $\left\{\zeta_{n}\right\}_{n}$, $\varphi^{*}(\xi_{0})\eqqcolon x_{0}$ and $\varphi^{*}(\xi_{m})\eqqcolon x_{m}$ are well-defined. Moreover, by the choice of $r$, we have $x_{0}\in D(x,R/2)$. Since $\xi_{0},\xi_{m}\in K^{n_{0}}_{*}$, we have
$\Delta_{\rho}(\xi_{0})\cap\Delta_{\rho}(\xi_{n})\neq\emptyset,$
and $\Delta_{\rho}(\xi_{0}),\Delta_{\rho}(\xi_{m})\subset A_{n_{0}}\subset D(x_{0},r/3)$. In particular, $x_{m}\in D(x_{0},r/3)$, so (1.1) and (1.4) holds. To see (1.2), note that $r$ has been chosen so that the inverse branches corresponding to $\left\{x_{n}\right\}_{n}$ are well-defined in $D(x_{0},r)$, and $m$ is large enough to that diam $\textrm{diam }F_{m}(D(x_{0},r))<r/3$, as desired. Requirement (1.3) is directly satisfied by the choice of $\rho_{m}$. Therefore, we have proved the lemma. $\square$
[…]
Figure 9.1: Situation after Lemma 9.2.1.
Next we prove the existence of a repelling periodic point in $D(x_{0},r)$. Note that, since $D(x_{0},r)\subset D(x,R)$, the proof of the next lemma ends the proof of the theorem.
Lemma 9.2.2.
The map $F_{m}$ has an attracting fixed point in $D(x_{0},r)$ which is accessible from $U$. Hence, $f$ has a repelling $m$-periodic point in $D(x_{0},r)\cap\partial U$.
Proof.
First note that $F_{m}(D(x_{0},r))\subset D(x_{0},r)$. Indeed, by (1.1) and (1.2), we have that $x_{m}\in D(x_{0},r/3)$ and diam $\textrm{diam }F_{m}(D(x_{0},r))<r/3$, so
$F_{m}(D(x_{0},r))\subset D(x_{m},2r/3)\subset D(x_{0},r).$
Therefore, by the Denjoy-Wolff Theorem, there exists a fixed point $p\in D(x_{0},r)$, which attracts all points in $D(x_{0},r)$ under the iteration of $F_{m}$. Hence, it is repelling under $f^{m}$ and thus belongs to $\mathcal{J}(f)$.
It is left to see that $p$ is accessible from $U$. To do so, first note that, by (1.3), the inverse branch $G_{m}$ of $g^{-m}$ is well-defined in $D(\xi_{0},\rho_{m})$, and it holds that
$\varphi\circ G_{m}=F_{m}\circ\varphi$
in $\Delta_{\rho_{m}}(\xi_{0})$. Moreover, we have that
$G_{m}(R_{\rho_{m}}(\xi_{0}))\subset\Delta_{\rho_{m}}(\xi_{m})\subset\Delta_{\rho}(\xi_{m}).$
By (1.4), $\Delta_{\rho}(\xi_{0})\cup\Delta_{\rho}(\xi_{m})$ is connected. Therefore, if we take $z\in R_{\rho_{m}}(\xi_{0})$, then $G_{m}(z)\in\Delta_{\rho}(\xi_{m})$, and we can find a curve $\gamma\subset\Delta_{\rho}(\xi_{0})\cup\Delta_{\rho}(\xi_{m})$ joining $z$ and $G_{m}(z)$. By (1.4), $\varphi(\gamma)\subset D(x_{0},r)$, and joins $\varphi(z)$ with $F_{m}(\varphi(z))$. See Figure 9.2.
[…]
Figure 9.2: The construction of the curve $\gamma$ in $\mathbb{D}$, and its image $\varphi(\gamma)$ in the dynamical plane.
Define
$\Gamma\coloneqq\bigcup\limits_{k\geq 0}F^{k}_{m}(\gamma).$
Then, $\Gamma\subset\partial U$ lands at $p$, ending the proof of Lemma 9.2.2, and hence of Theorem 9.1. $\square$
10 Further questions
Although in this paper we made an initial and exhaustive study of Pesin theory and Lyapunov exponents for the boundaries of Fatou components of transcendental maps, we left several interesting questions open, which arise in connection with our new results. We discuss them here.
First, it would be interesting to know whether our theorems hold without the assumption of thin singular values (or with weaker assumptions). A concrete field of interest would be determining when the Lyapunov exponent for the boundary of a Fatou component is well-defined, and when positive, in a more precise way than the estimates given in Section 4.
More ambitiously, it would be interesting to prove whether the equality between the Lyapunov exponents of the boundary map of an attracting basin and the radial extension of the associated inner function, that is
$\chi_{\lambda}(g)=\chi_{\omega_{U}}(f)$
holds. This is well-known for rational maps [Prz85], but the proof relies strongly on having finite degree and finitely many critical values in $\partial U$. Note that, if one can prove the previous equality, several tools used for finite entropy inner functions could be used in the setting of attracting basins. In particular, one would automatically deduce that $\chi_{\omega_{U}}(f)>0$.
Another direction for further work may be to construct examples of Fatou components for meromorphic functions with singular values accumulating to infinity through several accesses. We are not aware of any of such examples. It seems plausible to do so using approximation theory. It seems also possible to construct examples of centered inner functions with finite entropy with infinitely many singular values satisfying the conditions of Corollary C, using the results in [Ivr19] for prescribing critical points.
Finally, it is an appealing open question to know whether there always exists a periodic point in the boundary of an invariant Fatou component for functions in class $\mathbb{K}$. For rational functions, this is well-known since Fatou [Fat20, p. 81], but the proof uses strongly the compactness of $\partial U$ (compare with the treatment of accesses made for transcendental Fatou components in [BFJK17]). Hence, Corollary D answers this question for a large class of Fatou components, but it would be interesting to have a more general result. In the same direction, one can ask if periodic points are always dense in the boundaries of attracting basins, parabolic basins or doubly parabolic Baker domains (for a more complete discussion, see e.g. [Jov24]).